### v1

# Li Tana,b , Shengrong Wanga,c

Abstract
In this article, we propose a tamed theta Euler-Maruyama (EM) scheme for superlinearly growing neutral McKean-Vlasov stochastic differential delay equations driven by fractional Brownian motion with Hurst exponent $H\in(1/2,1)$. The analytical properties including uniqueness and existence, boundedness of moment and propagation of chaos are investigated. Moreover, the convergence rate of the numerical scheme is established.
Keywords :Neutral McKean-Vlasov SDEs; superlinearly growing; fractional Brownian motion; strong convergence rate; tamed theta EM
1 introduction
Stochastic differential equations (SDEs) with coefficients that depend on the law of the solution were first studied by McKean [1], who was inspired by Kac’s programme in Kinetic Theory [2]. Subsequently, they have become a research hotspot. Bossy and Talay [3, 4] chose to approximate the McKean-Vlasov limit by replicating the behavior with a system of $n$ weakly interacting particles. They also studied the numerical approximation of the solutions to the McKean-Vlasov SDEs. The existence and uniqueness of McKean-Vlasov SDEs were discussed in [5, 6]. Lacker [7] investigated the well-posedness and the propagation of chaos for McKean-Vlasov SDEs. Since the explicit solutions of McKean-Vlasov SDEs are difficult to obtain, many types of McKean-Vlasov SDEs can be approximated by implicit Euler-Maruyama (EM) scheme [8, 16], adaptive EM scheme [9, 11], tamed EM scheme [10, 14], truncated EM scheme [15] and so on. They approximated the solutions of McKean-Vlasov SDEs using numerical methods whose coefficients are globally (or locally) Lipschitz continuous or exhibit super-linear growth.
Recently, there has been a tendency for research to focus on the McKean-Vlasov SDEs driven by fractional Brownian motion. The form is
(1.1) ${\rm d}X_{t}=b(X_{t},\mathcal{L}(X_{t})){\rm d}t+\sigma(\mathcal{L}(X_{t})){\rm d}B_{t}^{H},$
where the coefficients $b:\mathbb{R}^{d}\times\mathcal{P}_{q}(\mathbb{R}^{d})\to\mathbb{R}^{d},\sigma:\mathcal{P}_{q}(\mathbb{R}^{d})\to\mathbb{R}^{d}\otimes\mathbb{R}^{d}$ are Borel-measurable, here, $\mathcal{P}_{q}(\mathbb{R}^{d})$ is the set of probability measures on $\mathbb{R}^{d}$ with finite $q$-th moment. $\mathcal{L}(X_{t})$ is the law of $X$ at $t$ and $B^{H}=\{B_{t}^{H},t\in[0,T]\}$ is a fractional Brownian motion with Hurst exponent $H\in(0,1)$ defined on the probability space $(\Omega,\mathcal{F},\mathbb{P})$. Moreover, $B^{H}$ is a centered Gaussian process with the covariance function
$R_{H}(t,s)=\mathbb{E}(B_{t}^{H}B_{s}^{H})=\frac{1}{2}(t^{2H}+s^{2H}-|t-s|^{2H}),\ \ \forall t,s\in[0,T].$
When $H=1/2$, the fractional Brownian motion $\{B_{t}^{H}\}_{t\geq 0}$ corresponds to a standard Bromnian motion. The self-similarity and long memory properties of fractional Brownian motion make it applicable in various fields, such as atmospheric sciences, financial engineering and others. From then on, scholars conducted studies on (1.1) separately. He et al. [17] established the theory of propagation of chaos and studied the strong convergence of the EM method when the coefficients of (1.1) satisfy the globally Lipschitz condition.
However, few studies considered the case where the coefficients of (1.1) grow super-linearly. Shen [18] demonstrated that, under certain averaging conditions, the solutions of McKean-Vlasov SDEs driven by fractional Brownian motion can be approximated by the solutions of the associated averaged McKean-Vlasov SDEs driven by fractional Brownian motion, in the sense of the mean square convergence, where the coefficients exhibit super-linear growth. Wang et al. [19] and Gao et al. [20] focus on the EM scheme for delayed McKean-Vlasov SDEs driven by fractional Brownian motion, whose drift coefficient contains a polynomial delay term.
Actually, in order to describe the impact of the past, stochastic delay differential equations (SDDEs) and neutral stochastic differential delay equations (NSDDEs) play a significant role in stochastic calculus, describing a wide variety of natural and man-made systems (see, e.g., [21, 22, 23]). The classical EM method is no longer applicable, as the coefficients of SDEs do not exhibit linear growth. It is known that the tamed EM shceme can be used to deal with superlinearly growing coefficients. Deng et al. [24] proposed two types of explicit tamed EM schemes for NSDDEs with superlinearly growing drift and diffusion coefficients and proved strong convergence for each of them, respectively. Cui et al. [25] studied the tamed EM methods to approximate the solutions for a type of of McKean-Vlasov NSDDEs in which the drift coefficients exhibit superlinear growth, and obtained the rate of convergence. We can refer to [23, 27] for more detail about tamed EM scheme in NSDDEs. It is worth noting that Tan and Yuan [28] modified the tamed theta EM scheme and showed the convergence of the numerical solutions, which provides us with research ideas for conducting this paper.
Drawing inspiration from the tamed approach introduced in [28] and previous models of McKean-Vlasov SDEs driven by fractional Brownian motion, we consider a tamed theta EM scheme for MV-NSDDEs driven by fractional Brownian motion with coefficients of superlinear growth, as an extension of (1.1). The form is
(1.2) ${\rm d}(X_{t}-D(X_{t-\tau}))=b(X_{t},X_{t-\tau},\mathcal{L}(X_{t})){\rm d}t+\sigma(\mathcal{L}(X_{t})){\rm d}B_{t}^{H},$
where $\tau>0$ is the delay.
The structure of the rest of the paper is as follows. In Section 2, the mathematical preliminaries on the McKean-Vlasov NSDDEs with superlinear growth driven by fractional Brownian motion are presented. In Section 3, existence and uniqueness of the solution is given. In Section 4, the propagation of chaos for this model is also analyzed. In Section 5, the EM scheme is established and the strong convergence rate is given.
2 Preliminaries
[…]
$\mathcal{P}_{q}(\mathbb{R}^{d})=\bigg{\{}\mu\in\mathcal{P}(\mathbb{R}^{d}):\int_{\mathbb{R}^{d}}|x|^{q}\mu({\rm d}x)<\infty\bigg{\}},$
which is a Polish space under the $L^{q}$-Wasserstein distance [30, Definition 6.1]. Then, for $\mu\in\mathcal{P}_{q}(\mathbb{R}^{d})$, define
$\mathcal{W}_{q}(\mu)=\bigg{(}\int_{\mathbb{R}^{d}}|x|^{q}\mu({\rm d}x)\bigg{)}^{1/q}$
For $\mu,\nu\in\mathcal{P}_{q}(\mathbb{R}^{d})$, we define the Wasserstein distance on $\mathcal{P}_{q}(\mathbb{R}^{d})$ by
$\mathbb{W}_{q}(\mu,\nu)=\bigg{(}\inf_{\pi\in\Pi(\mu,\nu)}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}|x-y|^{q}\pi({\rm d}x,{\rm d}y)\bigg{)}^{1/q},$
where $\Pi(\mu,\nu)$ is the set of probability measures on $\mathbb{R}^{d}\times\mathbb{R}^{d}$ with marginals $\mu$ and $\nu$. In other words, $\pi\in\Pi(\mu,\nu)$ is a probability measure on $\mathbb{R}^{d}\times\mathbb{R}^{d}$ such that $\pi(\cdot\times\mathbb{R}^{d})=\mu(\cdot)$ and $\pi(\cdot\times\mathbb{R}^{d})=\nu(\cdot)$. For $11/2<H<$ 1, we put
(2.1) $\phi(s,t)=\phi_{H}(s,t)=H(2H-1)|s-t|^{2H-2},\ \ s,t\in\mathbb{R},$
and recall that for $s,t>0$,
$\int_{0}^{t}\int_{0}^{s}\phi(u,v){\rm d}u{\rm d}v=R_{H}(t,s).$
Let $\mathcal{S}(\mathbb{R})$ be the Schwartz space of rapidly decreasing smooth functions on $R$, and if $b\in\mathcal{S}(\mathbb{R})$, denote
$\parallel f\parallel_{H}^{2}:=\int_{\mathbb{R}}\int_{\mathbb{R}}f(s)f(t)\phi(s,t){\rm d}s{\rm d}t<\infty.$
If we equip $\mathcal{S}(\mathbb{R})$ with the inner product
$\langle f,g\rangle_{H}:=\int_{\mathbb{R}}\int_{\mathbb{R}}f(s)g(t)\phi(s,t){\rm d}s{\rm d}t,\ \ f,g\in\mathcal{S}(\mathbb{R}),$
then the completion of $\mathcal{S}(\mathbb{R})$, denoted by $L_{\phi}^{2}(\mathbb{R})$,
$L_{\phi}^{2}(\mathbb{R}_{+})=\bigg{\{}g|g:\mathbb{R}_{+}\to\mathbb{R},\ \ \ |g|_{\phi}^{2}:=\int_{0}^{\infty}\int_{0}^{\infty}g(s)g(t)\phi(s,t){\rm d}s{\rm d}t<\infty\bigg{\}}.$
Definition 2.1.
[32, Definition 3.5.1] Let $g\in L_{\phi}^{2}(\mathbb{R})$. The $\phi$-derivative of a stochastic variable $F\in L^{p}$ in the direction of $\Phi_{g}$ is defined by
$D_{\Phi_{g}}F(\omega)=\lim_{\delta\to 0}\frac{1}{\delta}\bigg{\{}F\bigg{(}\omega+\delta\int_{0}^{\cdot}(\Phi_{g})(u){\rm d}u\bigg{)}-F(\omega)\bigg{\}},$
if the limit exists in $L^{p}$. Furthermore, if there is a process $(D_{s}^{\phi}F,s\geq 0)$ such that
$D_{\Phi_{g}}F=\int_{0}^{\infty}D_{s}^{\phi}Fg_{s}{\rm d}s\ \ \ {\rm almost\ surely},$
for all $g\in L_{\phi}^{2}(\mathbb{R})$, then $F$ is said to be $\phi$-differentiable.
For SDEs driven by fractional Brownian motion
(2.2) ${\rm d}X_{t}=G(t,X_{t}){\rm d}t+F(t,X_{t}){\rm d}B_{t}^{H}.$
Let $C^{2,1}(\mathbb{R}\times\mathbb{R}_{+};\mathbb{R})$ be the family of all real-valued functions $V(x,t)$ defined on $\mathbb{R}\times\mathbb{R}_{+}$, which is twice continuously differentiable in $x$ and once in $t$. Referring to the Itô formula for fractional Brownian motion given by [32, 33],
${\rm d}V(t,X_{t})=$ $\frac{\partial V(t,X_{t})}{\partial t}{\rm d}t+\frac{\partial V(t,X_{t})}{\partial X}G(t,X_{t}){\rm d}t+\frac{\partial V(t,X_{t})}{\partial X}F(t,X_{t}){\rm d}B_{t}^{H}$
$+\frac{\partial^{2}V(t,X_{t})}{\partial X^{2}}F(t,X_{t})D_{t}^{\phi}X_{t}{\rm d}t,$
or formally,
${\rm d}V(t,X_{t})=$ $\frac{\partial V(t,X_{t})}{\partial t}{\rm d}t+\frac{\partial V(t,X_{t})}{\partial X}G_{t,X_{t}}{\rm d}t+\frac{\partial V(t,X_{t})}{\partial X}F(t,X_{t}){\rm d}B_{t}^{H}$
$+\frac{\partial^{2}V(t,X_{t})}{\partial X^{2}}F(t,X_{t})\int_{0}^{t}\phi(t,r)F(r,X_{r}){\rm d}r{\rm d}t.$
In addition, [34], [35] also discussed the fractional Itô formula. Now, we are going to give the fractional Itô formula as follows.
Remark 2.1.
(Fractional Itô formula) Assume that the stochastic process $X_{t}$ follows (2.2) with $11/2<H<$ 1. We have
$V(t,X_{t})$
$=$ $V(X_{0})+\int_{0}^{t}\bigg{[}\frac{\partial V(s,X_{s})}{\partial t}+\frac{\partial V(s,X_{s})}{\partial X}G(s,X_{s})$
$+H(2H-1)\frac{\partial^{2}V(s,X_{s})}{\partial X^{2}}\int_{0}^{s}(s-r)^{2H-2}F^{2}(r,X_{r}){\rm d}r\bigg{]}{\rm d}s+\int_{0}^{t}\frac{\partial V(s,X_{s})}{\partial X}F(s,X_{s}){\rm d}B_{s}^{H}$
$=$ $V(X_{0})+\int_{0}^{t}\bigg{[}\frac{\partial V(s,X_{s})}{\partial t}+\frac{\partial V(s,X_{s})}{\partial X}G(s,X_{s})+H\frac{\partial^{2}V(s,X_{s})}{\partial X^{2}}s^{2H-1}F^{2}(s,X_{s})\bigg{]}{\rm d}s$
$+\int_{0}^{t}\frac{\partial V(s,X_{s})}{\partial X}F(s,X_{s}){\rm d}B_{s}^{H}.$
Since the kernel function $\phi$ is defined only for $11/2<H<$ 1, otherwise, $\phi$ has no meaning. We assume that $11/2<H<$ 1 throughout this paper. In this paper, we consider the following $d$-dimensional McKean-Vlasov NSDDEs driven by fractional Brownian motion as follows:
(2.5) $\left\{\begin{array}[]{ll}{\rm d}(X_{t}-D(X_{t-\tau}))=b(X_{t},X_{t-\tau},\mathcal{L}_{X_{t}}){\rm d}t+\sigma(\mathcal{L}_{X_{t}}){\rm d}B_{t}^{H},\ \ t>0,\\ X_{t}=\xi=\{\xi(t):-\tau\leq t\leq 0\}\in L_{\mathcal{F}_{0}}^{p}([-\tau,0];\mathbb{R}^{d}),\ p\geq\theta\geq 1\end{array}\right.$
where $b:\mathbb{R}^{d}\times\mathbb{R}^{d}\times\mathcal{P}_{q}(\mathbb{R}^{d})\to\mathbb{R}^{d}$, $\sigma:\mathcal{P}_{q}(\mathbb{R}^{d})\to\mathbb{R}^{d}\otimes\mathbb{R}^{d}$, $\mathcal{L}_{X_{t}}$ denotes the law of $X_{t}$. Throughout the paper, we shall denote $C$ by a generic positive constant, whose value may change from line to line. For any $x,y,\overline{x},\overline{y}$, we impose the following assumptions.
Assumption 2.2.
For any $s,t\in[-\tau,0]$ and $p>0$, there is a positive constant $K_{0}$ such that
$\mathbb{E}\bigg{(}\sup_{-\tau\leq s,t\leq 0}|\xi(s)-\xi(t)|^{p}\bigg{)}\leq K_{0}|s-t|^{p}.$
Assumption 2.3.
$D(0)=0$, and there exist positive constants $\lambda<1$ such that
$|D(x)-D(\overline{x})|\leq\lambda|x-\overline{x}|.$
Assumption 2.4.
There exist positive constants $l\geq 1$ and $K_{2},K_{3}$ such that
$\big{\langle}x-D(y)-\overline{x}+D(\overline{y}),b(x,y,\mu)-b(\overline{x},\overline{y},\nu)\big{\rangle}\leq K_{2}(|x-\overline{x}|^{2}+|y-\overline{y}|^{2}+(\mathbb{W}_{q}(\mu,\nu))^{2}),$
and
$|b(x,y,\mu)-b(\overline{x},\overline{y},\nu)|\leq K_{3}[(1+|x|^{l}+|\overline{x}|^{l}+|y|^{l}+|\overline{y}|^{l})(|x-\overline{x}|+|y-\overline{y}|)+\mathbb{W}_{q}(\mu,\nu)],$
for $x,y,\overline{x},\overline{y}\in\mathbb{R}^{d}$, $\mu,\nu\in\mathcal{P}_{q}(\mathbb{R}^{d})$.
Assumption 2.5.
There exist positive constants $K_{4},K_{5}$ such that
$\parallel\sigma(\mu)-\sigma(\nu)\parallel\leq K_{4}\mathbb{W}_{q}(\mu,\nu),$
and,
$|b(0,0,\mu)|\vee\parallel\sigma(\mu)\parallel\leq K_{5}(1+\mathbb{W}_{q}(\mu,\delta_{0})),$
for $\mu,\nu\in\mathcal{P}_{q}(\mathbb{R}^{d})$. Then, for initial experience distribution $\delta_{0}$
$|b(0,0,\delta_{0})|\vee\parallel\sigma(\delta_{0})\parallel\leq K_{6}.$
Remark 2.6.
Let Assumption 2.4 hold, then, for any $x,y\in\mathbb{R}^{d}$, we have
$|b(x,y,\mu)|\leq$ $|b(x,y,\mu)-b(0,0,\mu)|+|b(0,0,\mu)|$
$\leq$ $K_{3}(1+|x|^{l}+|y|^{l})(|x|+|y|)+K_{5}(1+\mathbb{W}_{q}(\mu,\delta_{0}))$
$\leq$ $C(1+|x|^{l+1}+|y|^{l+1}+\mathbb{W}_{q}(\mu,\delta_{0})),$
where $C=K_{3}\vee K_{5}$. Moreover, by Young’s inequality, we get
$(x-D(y))b(x,y,\mu)=$ $(x-D(y)-0+D(0))(b(x,y,\mu)-b(0,0,\delta_{0})+b(0,0,\delta_{0}))$
$\leq$ $K_{2}(|x|^{2}+|y|^{2}+\mathbb{W}_{q}^{2}(\mu,\delta_{0}))+(|x|+D(y))|b(0,0,\delta_{0})|$
$\leq$ $C(1+|x|^{2}+|y|^{2}+\mathbb{W}_{q}^{2}(\mu,\delta_{0})),$
where $C=K_{2}\vee(K_{2}+1)\vee(K_{2}+\lambda)\vee(K_{2}+K_{5}^{2})$.
Lemma 2.7.
[31, Lemma 4.1] Let $p>1$, $\varepsilon>0$ and $x,y\in\mathbb{R}$. Then
$|x+y|^{p}\leq\bigg{(}1+\varepsilon^{\frac{1}{p-1}}\bigg{)}^{p-1}\bigg{(}|x|^{p}+\frac{|y|^{p}}{\varepsilon}\bigg{)}.$
Especially for $\varepsilon=((1-a)/a)^{p-1}$, where $p\geq 2$ and $10<a<$ 1, we get
$|x+y|^{p}\leq\frac{1}{a^{p-1}}|x|^{p}+\frac{1}{(1-a)^{p-1}}|y|^{p}.$
3 Existence and Uniqueness
Since we are studying equations with super-linearly growing coefficients, it is necessary to consider the Carathéodory approximation. Now we define the Carathéodory approximation of (2.5) as follows. For every integer $n\geq 2/\tau$, define $X_{t}^{n}$ on $[-\tau,T]$ by
$X_{t}^{n}=\xi(t),\ \ t\in[-\tau,0],$
and
(3.1) $X_{t}^{n}-D(X_{t-\tau}^{n})=\xi(0)-D(\xi(-\tau))+\int_{0}^{t}b\big{(}X_{s-1/n}^{n},X_{s-\tau}^{n},\mathcal{L}_{X_{s-1/n}^{n}}\big{)}{\rm d}s+\int_{0}^{t}\sigma\big{(}\mathcal{L}_{X_{s-1/n}^{n}}\big{)}{\rm d}B_{s}^{H},$
for $t\in[0,T]$. It is important to note that each $X_{t}^{n}$ can be determined explicitly by the stepwise iterated Itô integrals over the intervals $[0,1/n]$, $(1/n,2/n]$ etc. To get the final results, we give the following lemmas.
Lemma 3.1.
Let Assumptions 2.2-2.5 hold. Then for $\overline{p}>1/H$ and $n\geq 2/\tau$, we get
$\mathbb{E}\bigg{(}\sup_{0\leq t\leq T}|X_{t}^{n}|^{\overline{p}}\bigg{)}\leq C.$
Proof.
By Assumption 2.4 and Lemma 2.7, we get
$|X_{t}^{n}|^{\overline{p}}=$ $|D(X_{t-\tau}^{n})+X_{t}^{n}-D(X_{t-\tau}^{n})|^{\overline{p}}$
$\leq$ $\frac{1}{\lambda^{p-1}}|D(X_{t-\tau}^{n})|^{\overline{p}}+\frac{1}{(1-\lambda)^{{\overline{p}}-1}}|X_{t}^{n}-D(X_{t-\tau}^{n})|^{\overline{p}}$
$\leq$ $\lambda|X_{t-\tau}^{n}|^{\overline{p}}+\frac{1}{(1-\lambda)^{p-1}}|X_{t}^{n}-D(X_{t-\tau}^{n})|^{\overline{p}},$
for any $t\in[0,T]$. Then
$\sup_{0\leq r\leq t}|X_{r}^{n}|^{\overline{p}}\leq$ $\lambda\parallel\xi\parallel^{p}+\lambda\sup_{0\leq r\leq t}|X_{r}^{n}|^{\overline{p}}+\frac{1}{(1-\lambda)^{{\overline{p}}-1}}\sup_{0\leq r\leq t}|X_{r}^{n}-D(X_{r-\tau}^{n})|^{\overline{p}},$
we may compute that
(3.2) $\sup_{0\leq r\leq t}|X_{r}^{n}|^{\overline{p}}\leq\frac{\lambda}{1-\lambda}\parallel\xi\parallel^{\overline{p}}+\frac{1}{(1-\lambda)^{\overline{p}}}\sup_{0\leq r\leq t}|X_{r}^{n}-D(X_{r-\tau}^{n})|^{\overline{p}}$
By Assumption 2.3, we get
$\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|X_{s}^{n}|^{\overline{p}}\bigg{)}$
$\leq$ $3^{\overline{p}-1}\mathbb{E}|\xi(0)-D(\xi(-\tau))|^{\overline{p}}+3^{\overline{p}-1}\mathbb{E}\bigg{(}\sup_{s\in[0,t]}\bigg{\lvert}\int_{0}^{s}b\left(X_{r-1/n}^{n},X_{r-\tau}^{n},\mathcal{L}_{X_{r-1/n}^{n}}\right){\rm d}r\bigg{\rvert}^{\overline{p}}\bigg{)}$
$+3^{\overline{p}-1}\mathbb{E}\bigg{(}\sup_{s\in[0,t]}\bigg{\lvert}\int_{0}^{s}\sigma\left(\mathcal{L}_{X_{r-1/n}^{n}}\right){\rm d}B_{r}^{H}\bigg{\rvert}^{\overline{p}}\bigg{)}$
$\leq$ $C+C\mathbb{E}\int_{0}^{t}\left[1+|X_{r-1/n}^{n}|^{(l+1)\overline{p}}+|X_{r-\tau}^{n}|^{(l+1){\overline{p}}}+\left(\mathbb{W}_{\theta}(\mathcal{L}_{X_{r-1/n}^{n}},\delta_{0})\right)^{\overline{p}}\right]{\rm d}r$
$+C\mathbb{E}\bigg{(}\sup_{s\in[0,t]}\bigg{\lvert}\int_{0}^{s}\sigma(\mathcal{L}_{X_{r-1/n}^{n}}){\rm d}B_{r}^{H}\bigg{\rvert}^{\overline{p}}\bigg{)}.$
Using the techniques of [12, Theorem 3.1]. Denote $\int_{s}^{t}(t-\beta)^{-\beta}(r-s)^{\beta-1}{\rm d}r:=C(\beta)$. Since for $\overline{p}H>1$, we can take some $\beta$ such that $1-H<\beta<1-1/{H}$, then by the stochastic Fubini theorem, the Hölder inequality, and combining with Lemma 2.4, we get
(3.3) $\mathbb{E}\bigg{(}\sup_{s\in[0,t]}\bigg{\lvert}\int_{0}^{s}\sigma(\mathcal{L}_{X_{r-1/n}^{n}}){\rm d}B_{r}^{H}\bigg{\rvert}^{\overline{p}}\bigg{)}$
$\leq$ $\frac{C(\beta)^{-{\overline{p}}}}{({\overline{p}}-1-\beta{\overline{p}})^{{\overline{p}}-1}}t^{{\overline{p}}-1-\beta{\overline{p}}}\int_{0}^{t}\mathbb{E}\bigg{|}\int_{0}^{s}(s-r)^{\beta-1}\sigma(\mathcal{L}_{X_{r-1/n}^{n}}){\rm d}B_{r}^{H}\bigg{|}^{{\overline{p}}}{\rm d}s$
$\leq$ $Ct^{{\overline{p}}-1-\beta{\overline{p}}}\int_{0}^{t}\left(\int_{0}^{s}(s-r)^{\frac{\beta-1}{H}}\parallel\sigma(\mathcal{L}_{X_{r-1/n}^{n}})\parallel^{\frac{1}{H}}{\rm d}r\right)^{\overline{p}H}{\rm d}s$
$\leq$ $Ct^{\overline{p}H-1}\int_{0}^{t}\parallel\sigma(\mathcal{L}_{X_{r-1/n}^{n}})\parallel^{\overline{p}}{\rm d}r$
$\leq$ $Ct^{\overline{p}H-1}L^{\overline{p}}\int_{0}^{t}\left[1+\mathbb{W}_{\theta}(\mathcal{L}_{X_{r-1/n}^{n}},\delta_{0})\right]^{\overline{p}}{\rm d}r.$
Referring to the proof in [6, Proposition 3.4], for the $q$-Wasserstein metric, we get
$\left(\mathbb{W}_{q}(\mathcal{L}_{X_{r-1/n}^{n}},\delta_{0})\right)^{\overline{p}}=\left(\mathcal{W}_{q}(\mathcal{L}_{X_{r-1/n}^{n}})\right)^{\overline{p}}\leq\mathbb{E}|X_{r-1/n}^{n}|^{\overline{p}}.$
To sum up, we obtain
$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}|X_{t}^{n}-D(X_{t-\tau}^{n})|^{\overline{p}}\bigg{)}\leq$ $C+C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{0\leq u\leq s}|X_{u-\tau}^{n}|^{(l+1)\overline{p}}\bigg{)}{\rm d}s,$
together with (3.2), we get
$\mathbb{E}\bigg{(}\sup_{0\leq r\leq t}|X_{r}^{n}|^{\overline{p}}\bigg{)}\leq$ $C+C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{0\leq u\leq s}|X_{u-\tau}^{n}|^{(l+1)\overline{p}}\bigg{)}{\rm d}s.$
We set up a finite sequence $\{u_{1},u_{2},\cdots,u_{[T/\tau]+1}\}$ where $u_{i}=([T/\tau]+2-i)\overline{p}(l+1)^{[T/\tau]+1-i}$, and $[T/\tau]$ denotes the integer part of $T/\tau$. It is easy to see that $u_{i}\geq\overline{p}$, $u_{[T/\tau]+1}=\overline{p}$ and $(l+1)u_{i+1}<u_{i}$ for $i=1,2,\cdots,[T/\tau]$. For $s\in[0,\tau]$, we immediately get
$\mathbb{E}\bigg{(}\sup_{0\leq r\leq t}|X_{r}^{n}|^{\overline{p}}\bigg{)}\leq C,$
since $\xi\in L_{\mathcal{F}_{0}}^{\overline{p}}([-\tau,0];\mathbb{R}^{d})$. For $s\in[0,2\tau]$, by the Hölder inequality, we get
$\mathbb{E}\bigg{(}\underset{s\in[0,2\tau]}{\sup}|X_{s}^{n}|^{u_{2}}\bigg{)}\leq C+C\int_{0}^{2\tau}\left[\mathbb{E}\bigg{(}\sup_{s\in[0,r]}|X_{s-\tau}^{n}|^{u_{1}}\bigg{)}\right]^{\frac{(l+1)u_{2}}{u_{1}}}{\rm d}r\leq C.$
By induction, we can derive the assertion. ∎
Lemma 3.2.
Let Assumptions 2.2-2.5 hold. Then for $0\leq s\leq t\leq T$ with $t-s<1$, and $p>1/H$, $(l+1)p\leq\overline{p}$, $n\geq 2/\tau$, we have
$\mathbb{E}|X_{t}^{n}-X_{s}^{n}|^{p}\leq C[(t-s)^{p}+(t-s)^{pH}],$
where $C$ is a positive constant depending on $p,H,\xi,T,L$ but independent of $s$ and $t$.
Proof.
By (3.1), we may compute
$X_{t}^{n}-X_{s}^{n}=$ $D(X_{t-\tau}^{n})-D(X_{s-\tau}^{n})+\int_{s}^{t}b\left(X^{n}_{r-1/n},X^{n}_{r-\tau},\mathcal{L}_{X_{r-1/n}^{n}}\right){\rm d}r+\int_{s}^{t}\sigma(\mathcal{L}_{X_{r-1/n}^{n}}){\rm d}B_{r}^{H}.$
By the elementary inequality, the Hölder inequality, Assumptions 2.3, 2.4, 2.5 and using same techniques as (3.3), we get
$\mathbb{E}|X_{t}^{n}-X_{s}^{n}|^{p}$
$\leq$ $3^{p-1}\mathbb{E}|D(X_{t-\tau}^{n})-D(X_{s-\tau}^{n})|^{p}+3^{p-1}\mathbb{E}\bigg{\lvert}\int_{s}^{t}b(X_{r-1/n}^{n},X_{r-\tau}^{n},\mathcal{L}_{X_{r-1/n}^{n}}){\rm d}r\bigg{\rvert}^{p}$
$+3^{p-1}\mathbb{E}\bigg{\lvert}\int_{s}^{t}\sigma(\mathcal{L}_{X_{r-1/n}^{n}}){\rm d}B_{r}^{H}\bigg{\rvert}^{p}$
$\leq$ $3^{p-1}\bigg{[}\lambda^{p}\mathbb{E}|X_{t-\tau}^{n}-X_{s-\tau}^{n}|^{p}+(t-s)^{p-1}$
×$\times\mathbb{E}\int_{s}^{t}C^{p}\bigg{(}1+|X_{r-1/n}^{n}|^{l+1}+|X_{r-\tau}^{n}|^{l+1}+\mathbb{W}_{q}(\mathcal{L}_{X_{r-1/n}^{n}},\delta_{0})\bigg{)}^{p}\bigg{]}{\rm d}r$
$+C(t-s)^{pH-1}\mathbb{E}\int_{s}^{t}\bigg{(}1+\mathbb{W}_{q}(\mathcal{L}_{X_{r-1/n}^{n}},\delta_{0})\bigg{)}^{p}\bigg{]}.$
For $0\leq s\leq t\leq\tau$, by Assumption 2.2, we get
$\mathbb{E}|X_{t}^{n}-X_{s}^{n}|^{p}$
$\leq$ $3^{p-1}\lambda^{p}\mathbb{E}|\xi(t-\tau)-\xi(s-\tau)|^{p}+C(t-s)^{p-1}\int_{s}^{t}\mathbb{E}\bigg{(}\sup_{u\in[-\tau,r]}|X_{u}^{n}|^{(l+1)p}\bigg{)}{\rm d}r$
$+C(t-s)^{pH-1}\int_{s}^{t}\mathbb{E}\bigg{(}\sup_{u\in[-\tau,r]}|X_{u}^{n}|^{p}\bigg{)}{\rm d}r$
$\leq$ $3^{p-1}\lambda^{p}K_{0}|t-s|^{p}+C(t-s)^{p-1}\int_{s}^{t}\mathbb{E}\bigg{(}\sup_{u\in[-\tau,r]}|X_{u}^{n}|^{(l+1)p}\bigg{)}{\rm d}r$
$+C(t-s)^{pH-1}\int_{s}^{t}\mathbb{E}\bigg{(}\sup_{u\in[-\tau,r]}|X_{u}^{n}|^{p}\bigg{)}{\rm d}r.$
Since $(l+1)p\leq\overline{p}$, we can applying Lemma 3.1 to obtain the assertion. For $\tau\leq s\leq t\leq T$, noting that $01-3^{p-1}\lambda^{p}>$ 0, we get the desired assertion by applying Lemma 3.1 again. ∎
Theorem 3.3.
Let Assumption 2.4 hold. Then for $p>1/H$ and $(l+1)p\leq\overline{p}$, (2.5) admits a global unique solution.
Proof.
We split the proof into two steps.
Step 1. Existence. For $m>n\geq 2/\tau$ and $t\in[0,T]$, by (2.5), we may compute
$\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|X^{m}_{s}-X_{s}^{n}|^{p}\bigg{)}$
$\leq$ $3^{p-1}\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|D(X_{s-\tau}^{m})-D(X_{s-\tau}^{n})|^{p}\bigg{)}$
$+3^{p-1}\mathbb{E}\bigg{(}\sup_{s\in[0,t]}\bigg{\lvert}\int_{0}^{s}\bigg{[}b\big{(}X_{r-1/m}^{m},X_{r-\tau}^{m},\mathcal{L}_{X_{r-1/m}^{m}}\big{)}-b\big{(}X_{r-1/n}^{n},X_{r-\tau}^{n},\mathcal{L}_{X_{r-1/n}^{n}}\big{)}\bigg{]}{\rm d}r\bigg{\rvert}^{p}\bigg{)}$
$+3^{p-1}\mathbb{E}\bigg{(}\sup_{s\in[0,t]}\bigg{\lvert}\int_{0}^{s}\bigg{[}\sigma\big{(}\mathcal{L}_{X_{r-1/m}^{m}}\big{)}-\sigma\big{(}\mathcal{L}_{X_{r-1/n}^{n}}\big{)}\bigg{]}{\rm d}B_{r}^{H}\bigg{\rvert}^{p}\bigg{)}$
=:absent:=:= : $W_{1}+W_{2}+W_{3},$
By Assumption 2.3, we get
$\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|D(X_{s-\tau}^{m})-D(X_{s-\tau}^{n})|^{p}\bigg{)}\leq\lambda^{p}\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|X_{s-\tau}^{m}-X_{s-\tau}^{n}|^{p}\bigg{)}.$
By the Hölder inequality, Assumption 2.4, we get
$W_{2}=$ $3^{p-1}\mathbb{E}\bigg{(}\sup_{s\in[0,t]}\bigg{\lvert}\int_{0}^{s}\bigg{[}b\big{(}X_{r-1/m}^{m},X_{r-\tau}^{m},\mathcal{L}_{X_{r-1/m}^{m}}\big{)}-b\big{(}X_{r-1/n}^{n},X_{r-\tau}^{n},\mathcal{L}_{X_{r-1/n}^{n}}\big{)}\bigg{]}{\rm d}r\bigg{\rvert}^{p}\bigg{)}$
$\leq$ $(3t)^{p-1}\mathbb{E}\int_{0}^{t}\bigg{\lvert}b\big{(}X_{r-1/m}^{m},X_{r-\tau}^{m},\mathcal{L}_{X_{r-1/m}^{m}}\big{)}-b\big{(}X_{r-1/n}^{n},X_{r-\tau}^{n},\mathcal{L}_{X_{r-1/n}^{n}}\big{)}\bigg{\rvert}^{p}{\rm d}r$
$\leq$ $(3t)^{p-1}K_{3}\mathbb{E}\int_{0}^{t}\bigg{[}\bigg{(}1+|X_{r-1/m}^{m}|^{l}+|X_{r-1/n}^{n}|^{l}+|X_{r-\tau}^{m}|^{l}+|X_{r-\tau}^{n}|^{l}\bigg{)}$
×$\times\bigg{(}|X_{r-1/m}^{m}-X_{r-1/n}^{n}|+|X_{r-\tau}^{m}-X_{s-\tau}^{n}|\bigg{)}+\mathbb{W}_{q}\big{(}\mathcal{L}_{X_{r-1/m}^{m}},\mathcal{L}_{X_{r-1/n}^{n}}\big{)}\bigg{]}^{p}{\rm d}r$
$\leq$ $C\mathbb{E}\int_{0}^{t}\bigg{(}1+|X_{r-1/m}^{m}|^{l}+|X_{r-1/n}^{n}|^{l}+|X_{r-\tau}^{m}|^{l}+|X_{r-\tau}^{n}|^{l}\bigg{)}^{p}$
×$\times\bigg{(}|X_{r-1/m}^{m}-X_{r-1/n}^{n}|+|X_{r-\tau}^{m}-X_{s-\tau}^{n}|\bigg{)}^{p}{\rm d}r$
$+C\mathbb{E}\int_{0}^{t}|X_{r-1/m}^{m}-X_{r-1/n}^{n}|^{p}{\rm d}r.$
Since for $l>1$, we have $2pl\leq(l+1)p\leq\overline{p}$, then by the Hölder inequality, Lemmas 3.1, 3.2, we get
$W_{2}\leq$ $C\int_{0}^{t}\bigg{[}\mathbb{E}\bigg{(}1+|X_{r-1/m}^{m}|^{l}+|X_{r-1/n}^{n}|^{l}+|X_{r-\tau}^{m}|^{l}+|X_{r-\tau}^{n}|^{l}\bigg{)}^{2p}\bigg{]}^{1/2}$
×$\times\bigg{[}\mathbb{E}\bigg{(}|X_{r-1/m}^{m}-X_{r-1/n}^{n}|+|X_{r-\tau}^{m}-X_{s-\tau}^{n}|\bigg{)}^{2p}\bigg{]}^{1/2}{\rm d}r+C\mathbb{E}\int_{0}^{t}|X_{r-1/m}^{m}-X_{r-1/n}^{n}|^{p}{\rm d}r$
$\leq$ $C\int_{0}^{t}\bigg{[}\mathbb{E}\bigg{(}|X_{r-1/m}^{m}-X_{r-1/n}^{n}|+|X_{r-\tau}^{m}-X_{r-\tau}^{n}|\bigg{)}^{2p}\bigg{]}^{1/2}{\rm d}r$
$+C\bigg{[}\bigg{(}\frac{1}{n}-\frac{1}{m}\bigg{)}^{p}+\bigg{(}\frac{1}{n}-\frac{1}{m}\bigg{)}^{pH}\bigg{]}.$
By Assumption 2.5, Lemmas 3.1, 3.2 and the same techniques as (3.3), we get
$W_{3}=$ $3^{p-1}\mathbb{E}\bigg{[}\sup_{s\in[0,t]}\bigg{\lvert}\int_{0}^{s}\bigg{(}\sigma(\mathcal{L}_{X_{r-1/m}^{m}})-\sigma(\mathcal{L}_{X_{r-1/n}^{n}})\bigg{)}{\rm d}B_{r}^{H}\bigg{\rvert}^{p}\bigg{]}$
$\leq$ $C\mathbb{E}\int_{0}^{t}|\sigma(\mathcal{L}_{X_{r-1/m}^{m}})-\sigma(\mathcal{L}_{X_{r-1/n}^{n}})|^{p}{\rm d}r$
$\leq$ $C\mathbb{E}\int_{0}^{t}K_{4}^{p}(\mathbb{W}_{q}(\mathcal{L}_{X_{r-1/m}^{m}}-\mathcal{L}_{X_{r-1/n}^{n}}))^{p}{\rm d}r$
$\leq$ $C\int_{0}^{t}\mathbb{E}|X_{r-1/m}^{m}-X_{r-1/n}^{n}|^{p}{\rm d}r$
$\leq$ $C\int_{0}^{t}\mathbb{E}|X_{r-1/m}^{m}-X_{r-1/n}^{m}|^{p}{\rm d}r+C\int_{0}^{t}\mathbb{E}|X_{r-1/n}^{m}-X_{r-1/n}^{n}|^{p}{\rm d}r$
$\leq$ $C\bigg{[}\bigg{(}\frac{1}{n}-\frac{1}{m}\bigg{)}^{p}+\bigg{(}\frac{1}{n}-\frac{1}{m}\bigg{)}^{pH}\bigg{]}+C\int_{0}^{t}\mathbb{E}|X_{r-1/n}^{m}-X_{r-1/n}^{n}|^{p}{\rm d}r.$
Combining $W_{1},W_{2},W_{3}$, we get
$\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|X^{m}_{s}-X_{s}^{n}|^{p}\bigg{)}$
$\leq$ $\lambda^{p}\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|X_{s-\tau}^{m}-X_{s-\tau}^{n}|^{p}\bigg{)}+C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{u\in[0,r]}|X_{u}^{m}-X_{u}^{n}|^{p}\bigg{)}{\rm d}r$
$+C\bigg{[}\bigg{(}\frac{1}{n}-\frac{1}{m}\bigg{)}^{p}+\bigg{(}\frac{1}{n}-\frac{1}{m}\bigg{)}^{pH}\bigg{]}+C\int_{0}^{t}\bigg{[}\mathbb{E}\bigg{(}\sup_{u\in[0,r]}|X_{u-\tau}^{m}-X_{u-\tau}^{n}|\bigg{)}^{2p}\bigg{]}^{1/2}{\rm d}r.$
Noting that $\lambda<1$, by the Gronwall inequality, we get
$\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|X^{m}_{s}-X_{s}^{n}|^{p}\bigg{)}$
$\leq$ $C\bigg{[}\bigg{(}\frac{1}{n}-\frac{1}{m}\bigg{)}^{p}+\bigg{(}\frac{1}{n}-\frac{1}{m}\bigg{)}^{pH}\bigg{]}+C\int_{0}^{t}\bigg{[}\mathbb{E}\bigg{(}\sup_{u\in[0,r]}|X_{u-\tau}^{m}-X_{u-\tau}^{n}|\bigg{)}^{2p}\bigg{]}^{1/2}{\rm d}r.$
Define a sequence $\{u_{1},u_{2},\cdots,u_{[T/\tau]+1}\}$ where $u_{i}=([T/\tau]+2-i)p2^{[T/\tau]+1-i}$, and $[T/\tau]$ denotes the integer part of $T/\tau$. It is easy to see that $u_{i}\geq p$, $u_{[T/\tau]+1}=p$ and $2u_{i+1}<u_{i}$ for $i=1,2,\cdots,[T/\tau]$. For $t\in[0,\tau]$, we get
(3.4) $\mathbb{E}\bigg{(}\sup_{s\in[0,\tau]}|X^{m}_{s}-X_{s}^{n}|^{u_{1}}\bigg{)}\to 0,\ {\rm for}\ m,n\to\infty.$
For $t\in[0,2\tau]$, by the Hölder inequality, we get
$\mathbb{E}\bigg{(}\underset{s\in[0,2\tau]}{\sup}|X^{m}_{s}-X_{s}^{n}|^{u_{2}}\bigg{)}$
[…]
by (3.4),
$\mathbb{E}\bigg{(}\underset{s\in[0,2\tau]}{\sup}|X^{m}_{s}-X_{s}^{n}|^{u_{2}}\bigg{)}\to 0,\ {\rm for}\ m,n\to\infty.$
By induction, we can see that for $t\in[0,T]$,
(3.5) $\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|X^{m}_{s}-X_{s}^{n}|^{p}\bigg{)}\to 0,\ {\rm for}\ m,n\to\infty.$
Consequently, $(X^{n})_{n\geq 2/\tau}$ is a Cauchy sequence for $p>1/H$. Denote the limit of $(X^{n})$ by $X$ and letting $m\to\infty$ in (3.5),
(3.6) $\lim_{n\to\infty}\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|X_{s}-X_{s}^{n}|^{p}\bigg{)}=0.$
Then, we should show that $X_{t}$ is a solution to (2.5). For $t\in[0,T]$, we get
$\mathbb{E}|X_{t}-X^{n}_{t-1/n}|^{p}=$ $\mathbb{E}|X_{t}-X^{n}_{t}+X^{n}_{t}-X^{n}_{t-1/n}|^{p}$
$\leq$ $2^{p-1}\mathbb{E}|X_{t}-X^{n}_{t}|^{p}+2^{p-1}\mathbb{E}|X^{n}_{t}-X^{n}_{t-1/n}|^{p}.$
By Lemma 3.2 and (3.6), we get
$\mathbb{E}|X_{t}-X^{n}_{t-1/n}|^{p}\to 0,\ \ {\rm as}\ \ n\to\infty.$
Taking $n\to\infty$ in (3.1), we get
$X_{t}-D(X_{t-\tau})=\xi(0)+D(\xi(-\tau))+\int_{0}^{t}b(X_{s},X_{s-\tau},\mathcal{L}_{X_{s}}){\rm d}s+\int_{0}^{t}\sigma(\mathcal{L}_{X_{s}}){\rm d}B_{s}^{H},\ \ t\in[0,T].$
Step 2. Uniqueness. Let $X_{t}$ and $Y_{t}$ be two solutions of (2.5) with $X_{t}=Y_{t}=\xi,\ \ t\in[-\tau,0]$. By Lemma 2.7 again, we get
$|X_{t}-Y_{t}|^{p}\leq\lambda|X_{t-\tau}-Y_{t-\tau}|^{p}+\frac{1}{(1-\lambda)^{p-1}}|X_{t}-D(X_{t-\tau})+Y_{t}-D(Y_{t-\tau})|^{p},$
which implies,
$\sup_{0\leq r\leq t}|X_{r}-Y_{r}|^{p}$
$\leq$ $\sup_{-\tau\leq r\leq 0}|X_{r}-Y_{r}|^{p}+\lambda\sup_{0\leq r\leq t}|X_{r}-Y_{r}|^{p}+\frac{1}{(1-\lambda)^{p-1}}\sup_{0\leq r\leq t}|X_{r}-D(X_{r-\tau})+Y_{r}-D(Y_{r-\tau})|^{p}.$
Since $X_{t}=Y_{t}=\xi,\ \ t\in[-\tau,0]$ and $\lambda\in(0,1)$,
$\sup_{0\leq r\leq t}|X_{r}-Y_{r}|^{p}\leq\frac{1}{(1-\lambda)^{p}}\sup_{0\leq r\leq t}|X_{r}-D(X_{r-\tau})+Y_{r}-D(Y_{r-\tau})|^{p}=\frac{1}{(1-\lambda)^{p}}\sup_{0\leq r\leq t}|\Pi(r)|^{p},$
where $\Pi(r)=X_{r}-D(X_{r-\tau})+Y_{r}-D(Y_{r-\tau})$. For $t\in[0,T]$, by the Itô formula and the elementary inequality, we get
$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}|\Pi(s)|^{p}\bigg{)}\leq$ $p\mathbb{E}\int_{0}^{t}|\Pi(r)|^{p-2}\langle\Pi(r),b(X_{r},X_{r-\tau},\mathcal{L}_{X_{r}})-b(Y_{r},Y_{r-\tau},\mathcal{L}_{Y_{r}})\rangle{\rm d}r$
$+p(p-1)H\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}\int_{0}^{s}|\Pi(r)|^{p-2}|\sigma(\mathcal{L}_{X_{r}})-\sigma(\mathcal{L}_{Y_{r}})|^{2}{\rm d}r\bigg{)}$
$+p\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}\int_{0}^{s}|\Pi(r)|^{p-1}|\sigma(\mathcal{L}_{X_{r}})-\sigma(\mathcal{L}_{Y_{r}})|{\rm d}B_{r}^{H}\bigg{)}.$
$=$ $J_{1}+J_{2}+J_{3}.$
By Assumption 2.4, the Young’s inequality, we get
$J_{1}=$ $C\mathbb{E}\int_{0}^{t}|X_{r}-D(X_{r-\tau})+Y_{r}-D(Y_{r-\tau})|^{p}{\rm d}r$
$+C\mathbb{E}\int_{0}^{t}\bigg{[}|X_{r}-Y_{r}|^{2}+|X_{r-\tau}-Y_{r-\tau}|^{2}+(\mathbb{W}_{q}(\mathcal{L}_{X_{r}},\mathcal{L}_{Y_{r}}))^{2}\bigg{]}^{p/2}{\rm d}r$
$\leq$ $C\mathbb{E}\int_{0}^{t}\bigg{(}\sup_{0\leq u\leq r}|X_{u}-Y_{u}|^{p}\bigg{)}{\rm d}r+C\mathbb{E}\int_{0}^{t}\bigg{(}\sup_{0\leq u\leq r}|X_{u-\tau}-Y_{u-\tau}|^{p}\bigg{)}{\rm d}r.$
By Assumption 2.5 and the Young’s inequality, we get
$J_{2}\leq$ $C\mathbb{E}\int_{0}^{t}|X_{r}-D(X_{r-\tau})+Y_{r}-D(Y_{r-\tau})|^{p}{\rm d}r+C\mathbb{E}\int_{0}^{t}\bigg{[}(\mathbb{W}_{q}(\mathcal{L}_{X_{r}},\mathcal{L}_{Y_{r}}))^{2}\bigg{]}^{p/2}{\rm d}r$
$\leq$ $C\mathbb{E}\int_{0}^{t}\bigg{(}\sup_{0\leq u\leq r}|X_{u}-Y_{u}|^{p}\bigg{)}{\rm d}r+C\mathbb{E}\int_{0}^{t}\bigg{(}\sup_{0\leq u\leq r}|X_{u-\tau}-Y_{u-\tau}|^{p}\bigg{)}{\rm d}r.$
In the way in (3.3), by Assumption 2.5, the Young’s inequality, the Hölder inequality, we get
$J_{3}\leq$ $C\mathbb{E}\int_{0}^{t}\bigg{(}\sup_{0\leq u\leq r}|X_{u}-Y_{u}|^{p}\bigg{)}{\rm d}r+C\mathbb{E}\int_{0}^{t}\bigg{(}\sup_{0\leq u\leq r}|X_{u-\tau}-Y_{u-\tau}|^{p}\bigg{)}{\rm d}r.$
Combining $J_{1}$-$J_{3}$ and applying the Gronwall inequality, we get
$\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|X_{s}-Y_{s}|^{p}\bigg{)}\leq C\mathbb{E}\int_{0}^{t}\bigg{(}\sup_{0\leq u\leq r}|X_{u-\tau}-Y_{u-\tau}|^{p}\bigg{)}{\rm d}r.$
By induction, we derive that $X_{t}=Y_{t},t\in[0,T]$, $\mathbb{P}$-a.s. The proof is complete. ∎
4 Propagation of Chaos
Since the coefficients of (2.5) depend on the distributions of state variables, we use stochastic interacting particle systems to approximate it. Firstly, considering a non-interacting particle system is given by
(4.1) ${\rm d}(X_{t}^{i}-D(X_{t-\tau}^{i}))=b(X_{t}^{i},X_{t-\tau}^{i},\mathcal{L}_{X_{t}^{i}}){\rm d}t+\sigma(\mathcal{L}_{X_{t}^{i}}){\rm d}B_{t}^{H,i},\ \ X_{0}^{i}=Y_{0}^{i}=\xi,\ \ i=1,2,\cdots,N,$
here, $|\xi_{t}^{i}-\xi_{s}^{i}|^{p}\leq|t-s|^{p}$. To deal with $\mathcal{L}_{X_{t}^{i}}$, we exploit the following interacting particle system of the form
(4.2) ${\rm d}(X_{t}^{i,N}-D(X_{t-\tau}^{i,N}))=b(X_{t}^{i,N},X_{t-\tau}^{i,N},\mu_{t}^{X,N}){\rm d}t+\sigma(\mu_{t}^{X,N}){\rm d}B_{t}^{H,i},\ \ i=1,2,\cdots,N,$
where the empirical measures is defined by
$\mu^{X,N}_{t}(\cdot)=\frac{1}{N}\sum_{j=1}^{N}\delta_{X_{t}^{j,N}}(\cdot),$
here, $\delta_{x}$ denotes the Dirac measure at point $x$. In order to prove the desired result, we give the following Lemma.
Lemma 4.1.
[17, Lemma 3.2] For two empirical measures $\mu_{t}^{N}=\frac{1}{N}\sum\limits_{j=1}\limits^{N}\delta_{Z^{1,j,N}_{t}}$ and $\nu_{t}^{N}=\frac{1}{N}\sum\limits_{j=1}\limits^{N}\delta_{Z_{t}^{2,j,N}}$ , then
$\mathbb{E}\left(\mathbb{W}_{q}(\mu_{t}^{N},\nu_{t}^{N})\right)\leq\mathbb{E}\bigg{(}\frac{1}{N}\sum_{j=1}^{N}\bigg{\lvert}Z_{t}^{1,j,N}-Z_{t}^{2,j,N}\bigg{\rvert}^{q}\bigg{)}^{1/q}.$
Next, we consider the theory of the propagation of chaos.
Lemma 4.2.
Let Assumptions 2.2-2.5 hold. For $\overline{p}>1/H$, we have
$\mathbb{E}\bigg{(}\sup_{t\in[-\tau,T]}|X_{t}^{i}|^{\overline{p}}\bigg{)}+\mathbb{E}\bigg{(}\sup_{t\in[-\tau,T]}|X_{t}^{i,N}|^{\overline{p}}\bigg{)}\leq C,\ \ i=1,2,\cdots,N,$
where $C$ is a positive constant depending on ${\overline{p}},T,L,H,\xi$.
Proof.
We will proceed in the same way in Lemma 3.1. By Assumption 2.4 and Lemma 2.7, we get
$|X_{t}^{i,N}|^{\overline{p}}=|D(X_{t-\tau}^{i,N})+X_{t}^{n}-D(X_{t-\tau}^{i,N})|^{\overline{p}}\leq\lambda|X_{t-\tau}^{i,N}|^{\overline{p}}+\frac{1}{(1-\lambda)^{{\overline{p}}-1}}|X_{t}^{i,N}-D(X_{t-\tau}^{i,N})|^{\overline{p}},$
for any $t\in[0,T]$. We immediately get
$\sup_{0\leq r\leq t}|X_{r}^{i,N}|^{\overline{p}}\leq\frac{\lambda}{1-\lambda}\parallel\xi\parallel^{\overline{p}}+\frac{1}{(1-\lambda)^{{\overline{p}}}}\sup_{0\leq r\leq t}|X_{r}^{i,N}-D(X_{r-\tau}^{i,N})|^{\overline{p}}$
By the Itô formula, we get
$|X_{t}^{i,N}-D(X_{t-\tau}^{i,N})|^{\overline{p}}$
$\leq$ $|\xi(0)-D(\xi(-\tau))|^{\overline{p}}$
$+p\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{{\overline{p}}-2}\big{\langle}X_{s}^{i,N}-D(X_{s-\tau}),b\big{(}X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N}\big{)}\big{\rangle}{\rm d}s$
$+{\overline{p}}\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{p-1}\sigma\big{(}\mu_{s}^{X,N}\big{)}{\rm d}B_{s}^{H}$
$+{\overline{p}}({\overline{p}}-1)\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{{\overline{p}}-2}\sigma\big{(}\mu_{s}^{X,N}\big{)}D_{s}^{\phi}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|{\rm d}s$
=:absent:=:= : $|\xi(0)-D(\xi(-\tau))|^{\overline{p}}+I_{1}(t)+I_{2}(t)+I_{3}(t).$
Then,
$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}|X_{t}^{i,N}-D(X_{t-\tau}^{i,N})|^{\overline{p}}\bigg{)}\leq\mathbb{E}|\xi(0)-D(\xi(-\tau))|^{\overline{p}}+\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}I_{1}+I_{2}+I_{3}\bigg{)}$
By Assumptions 2.3, 2.4, and the Young’s inequality, we get
$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}I_{1}(s)\bigg{)}$
$=$ $\mathbb{E}\int_{0}^{t}{\overline{p}}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{{\overline{p}}-2}\big{\langle}X_{s}^{i,N}-D(X_{s-\tau}^{i,N}),b\big{(}X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N}\big{)}\big{\rangle}{\rm d}s$
$\leq$ $C{\overline{p}}\mathbb{E}\int_{0}^{t}[|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{p-2}\times(1+|X_{s}^{i,N}|^{2}+|X_{s-\tau}^{i,N}|^{2}+(\mathbb{W}_{q}(\mu_{s}^{X,N},\delta_{0})))^{2}]{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}[|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{\overline{p}}+(1+|X_{s}^{i,N}|^{2}+|X_{s-\tau}^{i,N}|^{2}+(\mathbb{W}_{q}(\mu_{s}^{X,N},\delta_{0}))^{2})^{{\overline{p}}/2}]{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t\wedge\nu_{R}}[(1+|X_{s}^{i,N}|^{\overline{p}}+|X_{s-\tau}^{i,N}|^{\overline{p}}+\big{(}\mathbb{W}_{q}(\mu_{s}^{X,N},\delta_{0})\big{)}^{{\overline{p}}}]{\rm d}s.$
By the Hölder inequality, Young’s inequality and Assumption 2.5 and in the way in (3.3), we get
$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}I_{2}(s)\bigg{)}\leq$ $C\mathbb{E}\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{{\overline{p}}-1}|\sigma(\mu_{s}^{X,N})|{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{\overline{p}}{\rm d}s+C\mathbb{E}\int_{0}^{t}((\mathbb{W}_{q}(\mu_{s}^{X,N},\delta_{0}))^{\overline{q}}+1){\rm d}s.$
By Assumption 2.5, Remark 2.1, the Young’s inequality, we get
$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}I_{3}(s)\bigg{)}$
$\leq$ $C\mathbb{E}\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{\overline{p}}{\rm d}s+C\mathbb{E}\int_{0}^{t}\bigg{[}|\sigma\big{(}\mu_{s}^{X,N}\big{)}|^{2}\int_{0}^{s}H(2H-1)|r-s|^{2H-2}{\rm d}r\bigg{]}^{{\overline{p}}/2}{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{\overline{p}}{\rm d}s+2CHT^{2H-1}\mathbb{E}\int_{0}^{t}\lvert\sigma\big{(}\mu_{s}^{X,N}\big{)}\rvert^{\overline{p}}{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}[(|X_{s}^{n}|^{\overline{p}}+|X_{s-\tau}^{n}|^{\overline{p}}+\big{(}\mathbb{W}_{q}(\mu_{s}^{X,N},\delta_{0})\big{)}^{\overline{p}}]{\rm d}s$
$=$ $C\mathbb{E}\int_{0}^{t}[(|X_{s}^{n}|^{\overline{p}}+|X_{s-\tau}^{n}|^{\overline{p}}+C\mathbb{E}\int_{0}^{t}\bigg{(}\frac{1}{N}\sum_{j=1}^{N}|X_{r}^{j,N}|^{q}\bigg{)}^{{\overline{p}}/q}{\rm d}r,$
where for the last part, we uses the fact that
$\mathbb{W}_{\theta}(\mu_{s}^{X,N},\delta_{0})=\mathcal{W}_{q}(\mu_{s}^{X,N})=\bigg{(}\frac{1}{N}\sum_{j=1}^{N}|X_{s}^{j,N}|^{q}\bigg{)}^{1/q}.$
Since all $j$ are identically distributed, by the Minkowski inequality and Lemma 4.1, we get
$\mathbb{E}\bigg{(}\frac{1}{N}\sum_{j=1}^{N}|X_{s}^{j,N}|^{q}\bigg{)}^{{\overline{p}}/q}\leq\bigg{(}\frac{1}{N}\sum_{j=1}^{N}\parallel|X_{s}^{j,N}|^{q}\parallel_{L^{1/q}}\bigg{)}^{\overline{p}/q}=\bigg{(}\frac{1}{N}\sum_{j=1}^{N}(\mathbb{E}|X_{s}^{j,N}|)^{q}\bigg{)}^{\overline{p}/q}=\mathbb{E}|X_{s}^{i,N}|^{\overline{p}}.$
To sum up, we obtain
$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}|X_{t}^{i,N}-D(X_{t-\tau}^{i,N})|^{\overline{p}}\bigg{)}\leq$ $C+C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{0\leq u\leq s}|X_{u}^{i,N}|^{\overline{p}}\bigg{)}{\rm d}s+C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{0\leq u\leq s}|X_{u-\tau}^{i,N}|^{\overline{p}}\bigg{)}{\rm d}s,$
Thus, we get
$\mathbb{E}\bigg{(}\sup_{0\leq r\leq t}|X_{r}^{i,N}|^{\overline{p}}\bigg{)}\leq$ $C+C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{0\leq u\leq s}|X_{u}^{i,N}|^{\overline{p}}\bigg{)}{\rm d}s+C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{0\leq u\leq s}|X_{u-\tau}^{i,N}|^{\overline{p}}\bigg{)}{\rm d}s.$
For $t\in[0,\tau]$, we immediately get
$\mathbb{E}\bigg{(}\sup_{0\leq r\leq t}|X_{r}^{i,N}|^{\overline{p}}\bigg{)}\leq C+C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{0\leq u\leq s\wedge\nu_{R}}|X_{u}^{i,N}|^{\overline{p}}\bigg{)}{\rm d}s,$
since $\xi\in L_{\mathcal{F}_{0}}^{\overline{p}}([-\tau,0];\mathbb{R}^{d})$. We get the desired result by Gronwall inequality. By induction, we can derive the assertion. ∎
Lemma 4.3.
Let Assumption 2.4 hold. For $\overline{p}>1/H$ and $\overline{p}>2p$, we have
$\mathbb{E}\bigg{(}\sup_{t\in[0,T]}|X_{t}^{i}-X_{t}^{i,N}|^{p}\bigg{)}\leq C\times\left\{\begin{array}[]{ll}N^{-\frac{1}{2}},\ \ p>\frac{d}{2},\\ N^{-\frac{1}{2}}\log(1+N),\ p=\frac{d}{2},\\ N^{-\frac{p}{d}},\ p\in[2,\frac{d}{2}),\end{array}\right.$
where $C$ is a positive constant depending on $p,T,H,L,q$ but independent of $N$.
Proof.
For $i=1,2,\cdots,N$, by (4.2) and (4.1), set
$\Theta_{t}^{i}=X_{t}^{i}-D(X_{t-\tau}^{i})-X_{t}^{i,N}+D(X_{t-\tau}^{i,N}).$
By Itô formula, we get
$|\Theta_{t}^{i}|^{p}-|\Theta_{0}^{i}|^{p}$
$\leq$ $p\int_{0}^{t}|\Theta_{s}^{i}|^{p-1}\bigg{[}b(X_{s}^{i},X_{s-\tau}^{i},\mathcal{L}_{X_{s}^{i}})-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})\bigg{]}{\rm d}s$
$+p\int_{0}^{t}|\Theta_{t}^{i}|^{p-1}\bigg{[}\sigma(\mathcal{L}_{X_{s}^{i}})-\sigma(\mu_{s}^{X,N})\bigg{]}{\rm d}B_{s}^{H,i}$
$+p(p-1)\int_{0}^{t}|\Theta_{s}^{i}|^{p-2}\bigg{(}\sigma(\mathcal{L}_{X_{s}^{i}})-\sigma(\mu_{s}^{X,N})\bigg{)}\bigg{[}\int_{0}^{s}\phi(s,r)(\sigma(\mathcal{L}_{X_{r}^{i}})-\sigma(\mu_{r}^{X,N})){\rm d}r\bigg{]}{\rm d}s$
=:absent:=:= : $Q_{1}(t)+Q_{2}(t)+Q_{3}(t).$
For any $t\in[0,T]$, by Assumption 2.4, the Young’s inequality, we get
$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}Q_{1}(s)\bigg{)}$
$=$ $p\int_{0}^{t}|\Theta_{s}^{i}|^{p-1}\bigg{[}b(X_{s}^{i},X_{s-\tau}^{i},\mathcal{L}_{X_{s}^{i}})-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})\bigg{]}{\rm d}s$
$\leq$ $p\mathbb{E}\int_{0}^{t}|\Theta_{s}^{i}|^{p-2}\langle\Theta_{s}^{i},b(X_{s}^{i},X_{s-\tau}^{i},\mathcal{L}_{X_{s}^{i}})-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})\rangle{\rm d}s$
$\leq$ $C\int_{0}^{t}|\Theta_{s}^{i}|^{p}{\rm d}s+C\int_{0}^{t}\mathbb{E}[|X_{s}^{i}-X_{s}^{i,N}|^{p}+|X_{s-\tau}^{i}-X_{s-\tau}^{i,N}|^{p}+(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N}))^{p}]{\rm d}s.$
In the same way in (3.3), by Assumption 2.5, the Young’s inequality and the Hölder inequality, we get
$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}Q_{2}(s)\bigg{)}\leq$ $C\mathbb{E}\int_{0}^{t}|\Theta_{s}^{i}|^{\overline{p}-1}|\sigma(\mathcal{L}_{X_{s}^{i}})-\sigma(\mu_{s}^{X,N})|{\rm d}s$
$\leq$ $C\int_{0}^{t}|\Theta_{s}^{i}|^{p}{\rm d}s+C\int_{0}^{t}\mathbb{E}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N}))^{p}{\rm d}s.$
By Assumption 2.5, Remark 2.1, the Young’s inequality, we get
$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}Q_{3}(s)\bigg{)}\leq$ $C\mathbb{E}\int_{0}^{t}|\Theta_{s}^{i}|^{p}{\rm d}s+2CHT^{2H-1}\mathbb{E}\int_{0}^{t}\lvert\sigma\big{(}\mathcal{L}_{X_{u}^{i}}\big{)}-\sigma\big{(}\mu_{u}^{X,N}\big{)}\rvert^{p}{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}[(|\Theta_{s}^{i}|^{p}+\big{(}\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N})\big{)}^{p}]{\rm d}s.$
Noting that
$\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N})\leq\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X})+\mathbb{W}_{q}(\mu_{s}^{X},\mu_{s}^{X,N}),$
For $q<p$, by Lemma 4.1, we get
$\mathbb{E}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N}))^{p}=$ $\mathbb{E}[(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N}))^{q}]^{p/{q}}$
$\leq$ $\mathbb{E}\bigg{(}2^{q-1}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X}))^{q}+2^{q-1}(\mathbb{W}_{q}(\mu_{s}^{X},\mu_{s}^{X,N}))^{q}\bigg{)}^{p/q}$
$\leq$ $C\mathbb{E}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X}))^{p}+C\mathbb{E}(\mathbb{W}_{q}(\mu_{s}^{X},\mu_{s}^{X,N}))^{p}$
$\leq$ $C\mathbb{E}\bigg{(}\frac{1}{N}\sum_{j=1}^{N}|X_{s}^{j}-X_{s}^{j,N}|^{q}\bigg{)}^{p/q}+C\mathbb{E}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X}))^{p}$
$\leq$ $C\mathbb{E}|X_{s}^{i}-X_{s}^{i,N}|^{p}+C\mathbb{E}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X}))^{p}.$
Applying the Gronwall inequality, we get
$\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|\Theta_{s}^{i}|^{p}\bigg{)}$
$\leq$ $C\int_{0}^{t}\mathbb{E}|X_{s}^{i}-X_{s}^{i,N}|^{p}{\rm d}s+C\int_{0}^{t}\mathbb{E}|X_{s-\tau}^{i}-X_{s-\tau}^{i,N}|^{p}{\rm d}s+C\int_{0}^{t}\mathbb{E}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N}))^{p}{\rm d}s$
$\leq$ $\mathbb{E}\int_{0}^{t}\bigg{(}\sup_{s\in[0,t]}|X_{s}^{i}-X_{s}^{i,N}|^{p}\bigg{)}{\rm d}s+C\int_{0}^{t}\mathbb{E}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N}))^{p}{\rm d}s$
$\leq$ $\mathbb{E}\int_{0}^{t}\bigg{(}\sup_{s\in[0,t]}|X_{s}^{i}-X_{s}^{i,N}|^{p}\bigg{)}{\rm d}s+C\int_{0}^{t}\mathbb{E}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X}))^{p}{\rm d}s.$
According to (2.7), we obtain
$|X_{s}^{i}-X_{s}^{i,N}|^{p}\leq\lambda|X_{t-\tau}^{i}-X_{t-\tau}^{i,N}|^{p}+\frac{1}{(1-\lambda)^{p-1}}|\Theta_{s}^{i}|^{p}.$
Noting that $10<\lambda<$ 1, by the Gronwall inequality, we get
$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}|X_{s}^{i}-X_{s}^{i,N}|^{p}\bigg{)}\leq$ $C\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}|\Theta_{s}^{i}|^{p}\bigg{)}\leq C\int_{0}^{t}\mathbb{E}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X}))^{p}{\rm d}s.$
Since $(\mathbb{W}_{p}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X}))^{p}$ is controlled by the Wasserstein distance estimate([13], Theorem 1), for $\overline{p}>2p$, we get
$\mathbb{E}(\mathbb{W}_{p}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X}))^{p}\leq$ $C(\mathbb{E}|X_{s}^{i}|^{\overline{p}})^{p/{\overline{p}}}\times\left\{\begin{array}[]{ll}N^{-\frac{1}{2}},\ \ p>\frac{d}{2},\\ N^{-\frac{1}{2}}\log(1+N),\ p=\frac{d}{2},\\ N^{-\frac{p}{d}},\ p\in[2,\frac{d}{2}).\end{array}\right.$
Then, by Lemma 4.2, we complete the proof. ∎
5 The tamed theta scheme
In this section, we consider the tamed theta scheme to approximate solution of (4.2). Since the drift coefficient exhibits super-linear growth, we define
(5.1) $b_{\Delta}(x,y,\mu)=\frac{b(x,y,\mu)}{1+\Delta^{\alpha}|b(x,y,\mu)|},$
for $x,y\in\mathbb{R}^{d}$, $\mu\in\mathcal{P}_{q}(\mathbb{R}^{d})$ and $\alpha\in(0,1/2]$. Assume that for any time $T>0$, there exist positive constants $M,m\in\mathbb{N}$ such that $\Delta=\frac{\tau}{m}=\frac{T}{M}$, where $\Delta\in(0,1)$ is the step-size. Consequently, for the time grid $t_{k}=t_{0}+k\Delta$, it holds that $t_{k+1}-\tau=t_{k+1-m}$. For $k=-m,\cdots,0,$ we set $Y_{t_{k}}^{i,N}=\xi^{i}(t_{k})$ and $t_{0}=0$. For $k=0,1,\cdots,M$, we now define the discrete EM scheme for (4.2) as follows,
(5.2) $Y_{t_{k+1}}^{i,N}-D(Y_{t_{k+1-m}}^{i,N})=$ $Y_{t_{k}}^{i,N}-D(Y_{t_{k-m}}^{i,N})+\theta b_{\Delta}(Y_{t_{k+1}}^{i,N},Y_{t_{k+1-m}}^{i,N},\mu_{t_{k+1}}^{Y,N})\Delta$
$+(1-\theta)b_{\Delta}(Y_{t_{k}}^{i,N},Y_{t_{k-m}}^{i,N},\mu_{t_{k}}^{Y,N})\Delta+\sigma(\mu_{t_{k}}^{Y,N})\Delta B_{t_{k}}^{H,i},$
where the empirical measures $\mu_{t_{k}}^{Y,N}(\cdot):=\frac{1}{N}\sum\limits_{j=1}^{N}\delta_{Y_{t_{k}}^{j,N}}(\cdot)$ and $\Delta B_{t_{k}}^{H,i}=B_{t_{k+1}}^{H,i}-B_{t_{k}}^{H,i}$. Let
(5.5) $\overline{Y}_{\Delta}^{i,N}(t)=\left\{\begin{array}[]{ll}\xi^{i}(t),\ \ -\tau\leq t\leq 0,\\ \sum\limits_{k=0}^{[T/\Delta]}Y_{t_{k}}^{i,N}{\rm I}_{[k\Delta,(k+1)\Delta)}(t),\ \ t\geq 0,\end{array}\right.$
and
$\overline{Y}_{\Delta+}^{i,N}(t)=\left\{\begin{array}[]{ll}\xi^{i}(t),\ \ -\tau\leq t\leq 0,\\ \sum\limits_{k=0}^{[T/\Delta]}Y_{t_{k+1}}^{i,N}{\rm I}_{[k\Delta,(k+1)\Delta)}(t),\ \ t\geq 0,\end{array}\right.$
and the continuous EM solution writes as
(5.6) $Y_{\Delta}^{i,N}(t)=$ $D(\overline{Y}_{\Delta}^{i,N}(t))+\xi^{i}(0)-D(\xi^{i}(-\tau))+\theta\int_{0}^{t}b_{\Delta}(\overline{Y}_{\Delta+}^{i,N}(s),\overline{Y}_{\Delta+}^{i,N}(s-\tau),\overline{\mu}^{Y_{+},N}(s)){\rm d}s$
$+(1-\theta)\int_{0}^{t}b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s)){\rm d}s+\int_{0}^{t}\sigma(\overline{\mu}^{Y,N}(s)){\rm d}B_{s}^{H,i},\ \ t>0,$
where $\overline{\mu}_{s}^{Y,N}:=\frac{1}{N}\sum\limits_{j=1}^{N}\delta_{\overline{Y}_{t}^{j,N}}(\cdot)$, and $Y_{\Delta}^{i,N}(t)=\xi^{i}(t)$ when $t_{k}\in[-\tau,0]$. However, this $Y_{\Delta}^{i,N}(t)$ is not $\mathcal{F}_{t}$-adapted, it does not meet the fundamental requirement in the Itô stochastic analysis. To avoid calculus, we use the discrete split-step theta scheme (see [28]) described as follows: For $k=-m,\cdots,-1$, set $z_{t_{k}}^{i,N}=\xi^{i}(k\Delta)$. For $k=0,1,\cdots,M-1$, we rewrite (5.2) as
(5.9) $\left\{\begin{array}[]{ll}Y_{t_{k}}^{i,N}=D(Y_{t_{k-m}}^{i,N})+z_{t_{k}}^{i,N}-D(z_{t_{k-m}})+\theta b_{\Delta}(Y_{t_{k}}^{i,N},Y_{t_{k-m}}^{i,N},\mu_{t_{k}}^{Y,N})\Delta,\\ z_{t_{k+1}}^{i,N}=D(z_{t_{k+1-m}}^{i,N})+z_{t_{k}}^{i,N}-D(z_{t_{k-m}}^{i,N})+b_{\Delta}(Y_{t_{k}}^{i,N},Y_{t_{k-m}}^{i,N},\mu_{t_{k}}^{Y,N})\Delta+\sigma(\mu_{t_{k}}^{Y,N})\Delta B_{t_{k}}^{H,i},\end{array}\right.$
This scheme can also be rewritten as
$z_{t_{k+1}}^{i,N}-D(z_{t_{k+1-m}}^{i,N})=z_{t_{0}}^{i,N}-D(z_{t_{-m}}^{i,N})+\sum_{j=0}^{k}b_{\Delta}(Y_{t_{j}}^{i,N},Y_{t_{j-m}}^{i,N},\mu_{t_{j}}^{Y,N})\Delta+\sum_{j=0}^{k}\sigma(\mu_{t_{j}}^{Y,N})\Delta B_{t_{j}}^{H,i}$
$=$ $\xi^{i}(0)-D(\xi^{i}(-\tau))-\theta b_{\Delta}(\xi^{i}(0),\xi^{i}(-\tau),\mu_{t_{0}}^{i,N})\Delta+\sum_{j=0}^{k}b_{\Delta}(Y_{t_{j}}^{i,N},Y_{t_{j-m}}^{i,N},\mu_{t_{j}}^{Y,N})\Delta$
$+\sum_{j=0}^{k}\sigma(\mu_{t_{j}}^{Y,N})\Delta B_{t_{j}}^{H,i}.$
Similarly, we define the corresponding continuous-time split-step tamed theta solution $Z_{\Delta}^{i,N}(t)$ as follows: For any $t\in[-\tau,0]$, $Z_{\Delta}^{i,N}(t)=\xi^{i}(t)$, $Z_{\Delta}^{i,N}(0)=\xi^{i}(0)-\theta b_{\Delta}(\xi^{i}(0)$, $\xi^{i}(-\tau),\mu_{0}^{Y,N})\Delta$. For any $t\in[0,T]$,
(5.10) ${\rm d}[Z_{\Delta}^{i,N}(t)-D(Z_{\Delta}^{i,N}(t-\tau))]=b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(t),\overline{Y}_{\Delta}^{i,N}(t-\tau),\overline{\mu}^{Y,N}(t)){\rm d}t+\sigma(\overline{\mu}^{Y,N}(t)){\rm d}B_{t}^{H,i},$
where $\overline{Y}_{\Delta}^{i,N}(t)$ is defined by (5.5). Combining (5.6), (5.9) and (5.10), we have
$Y_{\Delta}^{i,N}(t)-D(Y_{\Delta}^{i,N}(t-\tau))-\theta b_{\Delta}(Y_{\Delta}^{i,N}(t),Y_{\Delta}^{i,N}(t-\tau),\mu^{Y,N}(t))\Delta=Z_{\Delta}^{i,N}(t)-D(Z_{\Delta}^{i,N}(t-\tau)).$
Let $\widetilde{Y}_{\Delta}^{i,N}(t)=Y_{\Delta}^{i,N}(t)-D(Y_{\Delta}^{i,N}(t-\tau))-\theta b_{\Delta}(Y_{\Delta}^{i,N}(t),Y_{\Delta}^{i,N}(t-\tau),\mu^{Y,N}(t))\Delta$. We can rewrite (5.10) as
(5.11) $\widetilde{Y}_{\Delta}^{i,N}(t)=\widetilde{Y}_{\Delta}^{i,N}(0)+\int_{0}^{t}b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s)){\rm d}s+\int_{0}^{t}\sigma(\overline{\mu}^{Y,N}(s)){\rm d}B_{s}^{H,i}.$
It is clear that $\widetilde{Y}_{\Delta}^{i,N}$ coincide with $\overline{Y}_{\Delta}^{i,N}(t)-D(\overline{Y}_{\Delta}^{i,N}(t-\tau))-\theta b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(t),\overline{Y}_{\Delta}^{i,N}(t-\tau),\overline{\mu}^{Y,N}(t))$ at grid points $t=k\Delta$, $k=0,1,\cdots,M-1$, this also means that the continuous-time tamed theta solution $Y_{\Delta}^{i,N}(t)$ and the discrete-time tamed theta solution $\overline{Y}_{\Delta}^{i,N}(t)$ take the same values at grid points $t=k\Delta$, $k=0,1,\cdots,M-1$.
Lemma 5.1.
There exists a positive constant $K_{7}\geq 1$ such that
$|b_{\Delta}(x,y,\mu)|\leq\min\{K_{7}\Delta^{-\alpha}(1+|x|+|y|+\mathbb{W}_{q}(\mu,\delta_{0})),|b(x,y,\mu)|\}.$
Proof.
For $\forall x,y>0$, by (5.1), it is easy to see that $|b_{\Delta}(x,y,\mu)|\leq|b(x,y,\mu)|$. On the other hand, we immediately get
$|b_{\Delta}(x,y,\mu)|=\Delta^{-\alpha}\frac{|b(x,y,\mu)|}{\Delta^{-\alpha}+|b(x,y,\mu)|}\leq\Delta^{-\alpha}\leq K_{7}\Delta^{-\alpha}(1+|x|+|y|+\mathbb{W}_{q}(\mu,\delta_{0})).$
Lemma 5.2.
Let Assumption 2.4 hold. One observes
$\langle x-D(y)-\overline{x}+D(\overline{y}),b_{\Delta}(x,y,\mu)-b_{\Delta}(\overline{x},\overline{y},\nu)\rangle\leq K_{2}(|x-\overline{x}|^{2}+|y-\overline{y}|^{2}+(\mathbb{W}_{q}(\mu,\nu))^{2}),$
and
$|b(x,y,\mu)-b_{\Delta}(x,y,\mu)|^{p}\leq K_{8}\Delta^{\alpha p}(1+|x|^{2(l+1)p}+|y|^{2(l+1)p}+(\mathbb{W}_{q}(\mu,\delta_{0}))^{2p}),$
where $K_{8}$ is a positive constant.
Proof.
We divide it into two case. Case 1, for $b(x,y,\mu)\cdot b(\overline{x},\overline{y},\nu)<0$, we get
$\langle x-D(y)-\overline{x}+D(\overline{y}),b_{\Delta}(x,y,\mu)-b_{\Delta}(\overline{x},\overline{y},\nu)\rangle$
$=$ $\bigg{\langle}x-D(y)-\overline{x}+D(\overline{y}),\frac{b(x,y,\mu)}{1+\Delta^{\alpha}|b(x,y,\mu)|}-\frac{b(\overline{x},\overline{y},\nu)}{1+\Delta^{\alpha}|b(\overline{x},\overline{y},\nu)|}\bigg{\rangle}$
$\leq$ $K_{2}(|x-\overline{x}|^{2}+|y-\overline{y}|^{2}+(\mathbb{W}_{q}(\mu,\nu))^{2}).$
Case 2, for $b(x,y,\mu)\cdot b(\overline{x},\overline{y},\nu)>0$, we get
$\langle x-D(y)-\overline{x}+D(\overline{y}),b_{\Delta}(x,y,\mu)-b_{\Delta}(\overline{x},\overline{y},\nu)\rangle$
$=$ $\bigg{\langle}x-D(y)-\overline{x}+D(\overline{y}),\frac{b(x,y,\mu)-b(\overline{x},\overline{y},\nu)}{(1+\Delta^{\alpha}|b(x,y,\mu)|)(1+\Delta^{\alpha}|b(\overline{x},\overline{y},\nu))}\bigg{\rangle}$
$+\bigg{\langle}x-D(y)-\overline{x}+D(\overline{y}),\frac{\Delta^{\alpha}[b(x,y,\mu)\cdot b(\overline{x},\overline{y},\nu)-b(x,y,\mu)\cdot b(\overline{x},\overline{y},\nu)]}{(1+\Delta^{\alpha}|b(x,y,\mu)|)(1+\Delta^{\alpha}|b(\overline{x},\overline{y},\nu))}\bigg{\rangle}$
$\leq$ $K_{2}(|x-\overline{x}|^{2}+|y-\overline{y}|^{2}+(\mathbb{W}_{q}(\mu,\nu))^{2}).$
Furthermore,
$|b(x,y,\mu)-b_{\Delta}(x,y,\mu)|^{p}\leq\Delta^{\alpha p}|b(x,y,\mu)|^{2p},$
the second part is obvious. ∎
Remark 5.3.
Given Lemmas 5.1, 5.2 hold. We have
$\langle x-D(y),b_{\Delta}(x,y,\mu)\rangle\leq C(1+|x|^{2}+|y|^{2}+(\mathbb{W}_{q}(\mu,\delta_{0}))^{2}).$
5.1 Moment bounds
In order to get the desired results, we give the following lemma.
Lemma 5.4.
Let Assumption 2.2-2.5 and Lemmas 5.1, 5.2 hold. For $\overline{p}>1/H$, we have
$\sup_{0\leq t\leq T}\mathbb{E}|Y_{\Delta}^{i,N}(t)|^{\overline{p}}\leq C,$
where the positive constant $C$ is independent of $\Delta$.
Proof.
For $a>0$, let $\lfloor a\rfloor$ be the integer part of $a$. Applying the Itô formula to $[1+|\widetilde{Y}_{\Delta}^{i,N}(t)|^{2}]^{\overline{p}/2}$, we get
$\mathbb{E}[1+|\widetilde{Y}_{\Delta}^{i,N}(t)|^{2}]^{\overline{p}/2}$
$\leq$ $\mathbb{E}[1+|\widetilde{Y}_{\Delta}^{i,N}(0)|^{2}]^{\overline{p}/2}+p\mathbb{E}\int_{0}^{t}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\rangle$
$+\overline{p}(\overline{p}-2)\mathbb{E}\int_{0}^{t}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}\sigma(\overline{\mu}^{Y,N}(s))D_{s}^{\phi}[1+|\widetilde{Y}_{\Delta}(s)|^{2}]{\rm d}s$
$\leq$ $\mathbb{E}[1+|\widetilde{Y}_{\Delta}^{i,N}(0)|^{2}]^{\overline{p}/2}+\overline{p}(\overline{p}-2)\mathbb{E}\int_{0}^{t}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}|\sigma(\overline{\mu}^{Y,N}(s))|^{2}\int_{0}^{s}\phi(s,r){\rm d}r{\rm d}s$
$+\overline{p}\mathbb{E}\int_{0}^{t}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}\langle\overline{Y}_{\Delta}^{i,N}(s)-D(\overline{Y}_{\Delta}^{i,N}(s-\tau)),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\rangle$
$+\overline{p}\mathbb{E}\int_{0}^{t}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(s)-\widehat{Y}_{\Delta}^{i,N}(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\rangle{\rm d}s$
=:absent:=:= : $\mathbb{E}[1+|\widetilde{Y}_{\Delta}^{i,N}(0)|^{2}]^{\overline{p}/2}+G_{1}(t)+G_{2}(t)+G_{3}(t).$
where $\widehat{Y}_{\Delta}^{i,N}(s)=\overline{Y}_{\Delta}^{i,N}(s)-D(\overline{Y}_{\Delta}^{i,N}(s-\tau))-\theta b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\Delta$. By the properties of $\phi$ and the Young’s inequality, Lemma 5.1, we get
$G_{1}(t)$
$\leq$ $CHT^{2H-1}\mathbb{E}\int_{0}^{t}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}|\sigma(\overline{\mu}^{Y,N}(s))|^{2}{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}\big{(}1+[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}}{2}}+|\overline{Y}_{\Delta}^{i,N}(s)|^{\overline{p}}\big{)}{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}\bigg{(}1+|Y_{\Delta}^{i,N}(s)|^{\overline{p}}+|Y_{\Delta}^{i,N}(s-\tau)|^{\overline{p}}+|\theta b_{\Delta}(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))\Delta|^{\overline{p}}+|\overline{Y}_{\Delta}^{i,N}(s)|^{\overline{p}}\bigg{)}{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}\bigg{(}1+|Y_{\Delta}^{i,N}(s)|^{\overline{p}}+|Y_{\Delta}^{i,N}(s-\tau)|^{\overline{p}}+|\overline{Y}_{\Delta}^{i,N}(s)|^{\overline{p}}\bigg{)}{\rm d}s$
$+C\Delta^{(1-\alpha)\overline{p}}\mathbb{E}\int_{0}^{t}\bigg{(}1+|Y_{\Delta}^{i,N}(s)|^{\overline{p}}+|Y_{\Delta}^{i,N}(s-\tau)|^{\overline{p}}\bigg{)}$
$\leq$ $C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)|^{\overline{p}}{\rm d}s.$
By Remark 5.3 and the Young’s inequality, we get
$G_{2}(t)\leq$ $C\mathbb{E}\int_{0}^{t}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}\bigg{(}1+|\overline{Y}_{\Delta}^{i,N}(s)|^{2}+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|^{2}+(\mathbb{W}_{q}(\mu^{\overline{Y},N}(s),\delta_{0}))^{2}\bigg{)}{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}\bigg{(}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}}{2}}+1+|\overline{Y}_{\Delta}^{i,N}(s)|^{\overline{p}}+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|^{\overline{p}}\bigg{)}{\rm d}s$
$\leq$ $C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)|^{\overline{p}}{\rm d}s.$
We may compute that
$G_{3}(t)=$ $\overline{p}\mathbb{E}\int_{0}^{t}[1+|\widehat{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(s)-\widehat{Y}_{\Delta}^{i,N}(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\rangle{\rm d}s$
$+p\mathbb{E}\int_{0}^{t}\bigg{[}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}-[1+|\widehat{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}\bigg{]}$
×$\times\langle\widetilde{Y}_{\Delta}^{i,N}(s)-\widehat{Y}_{\Delta}^{i,N}(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\rangle{\rm d}s$
=:absent:=:= : $\overline{p}G_{31}(t)+\overline{p}G_{32}(t),$
where $\widetilde{Y}_{\Delta}^{i,N}(t)-\widehat{Y}_{\Delta}^{i,N}(s)=\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u)){\rm d}u+\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}\sigma(\overline{\mu}^{Y,N}(u)){\rm d}B_{u}^{H,i}$. By Lemma 5.1 and the Young’s inequality, we get
$G_{31}(t)$
$=$ $\mathbb{E}\int_{0}^{t}[1+|\widehat{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}\bigg{\langle}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(u)){\rm d}u,$
$b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\bigg{\rangle}{\rm d}s$
$+\mathbb{E}\int_{0}^{t}[1+|\widehat{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}\bigg{\langle}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}\sigma(\overline{\mu}^{Y,N}(u)){\rm d}B_{u}^{H,i},b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\bigg{\rangle}{\rm d}s$
$\leq$ $\mathbb{E}\int_{0}^{t}[1+|\widehat{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(u))|{\rm d}u$
×$\times|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))|{\rm d}s$
$\leq$ $\Delta\mathbb{E}\int_{0}^{t}[1+|\widehat{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))|^{2}{\rm d}s$
$\leq$ $C\Delta^{1-2\alpha}\mathbb{E}\int_{0}^{t}\bigg{[}(1+|\overline{Y}_{\Delta}^{i,N}(u)|^{\overline{p}}+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|^{\overline{p}})+\Delta^{(1-\alpha)\overline{p}}(1+|\overline{Y}_{\Delta}^{i,N}(u)|^{\overline{p}}+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|^{\overline{p}})\bigg{]}{\rm d}s$
$\leq$ $C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)|^{\overline{p}}{\rm d}s.$
Applying Itô formula to $[1+|\widetilde{Y}_{\Delta}(s)|^{2}]^{\frac{\overline{p}-2}{2}}$, we get
$[1+|\widetilde{Y}_{\Delta}(s)|^{2}]^{\frac{\overline{p}-2}{2}}$
$\leq$ $[1+|\widetilde{Y}_{\Delta}^{i,N}(0)|^{2}]^{\frac{\overline{p}-2}{2}}$
$+(p-2)\int_{0}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u))\rangle{\rm d}u$
$+(\overline{p}-2)(\overline{p}-4)\int_{0}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\sigma(\overline{\mu}^{Y,N}(u))D_{r}^{\phi}[1+|\widetilde{Y}_{\Delta}(u)|^{2}]^{\frac{\overline{p}-2}{2}}{\rm d}u$
$+(\overline{p}-2)\int_{0}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),\sigma(\overline{\mu}^{Y,N}(u))\rangle{\rm d}B_{u}^{H,i}.$
Then,
$[1+|\widehat{Y}_{\Delta}(s)|^{2}]^{\frac{\overline{p}-2}{2}}$
$\leq$ $[1+|\widehat{Y}_{\Delta}^{i,N}(0)|^{2}]^{\frac{\overline{p}-2}{2}}$
$+(\overline{p}-2)\int_{0}^{\lfloor\frac{s}{\Delta}\rfloor\Delta}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u))\rangle{\rm d}u$
$+(\overline{p}-2)(\overline{p}-4)\int_{0}^{\lfloor\frac{s}{\Delta}\rfloor\Delta}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\sigma(\overline{\mu}^{Y,N}(u))D_{u}^{\phi}[1+|\widetilde{Y}_{\Delta}(u)|^{2}]^{\frac{\overline{p}-2}{2}}{\rm d}u$
$+(\overline{p}-2)\int_{0}^{\lfloor\frac{s}{\Delta}\rfloor\Delta}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),\sigma(\overline{\mu}^{Y,N}(u))\rangle{\rm d}B_{u}^{H,i}.$
One immediately gets
$G_{32}(t)\leq$ $(\overline{p}-2)\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u))\rangle{\rm d}u$
×$\times\langle\widetilde{Y}_{\Delta}^{i,N}(s)-\widehat{Y}_{\Delta}^{i,N}(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\rangle{\rm d}s$
$+(\overline{p}-2)(\overline{p}-4)\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\sigma(\overline{\mu}^{Y,N}(u))D_{u}^{\phi}[1+|\widetilde{Y}_{\Delta}(u)|^{2}]^{\frac{\overline{p}-2}{2}}{\rm d}u$
×$\times\langle\widetilde{Y}_{\Delta}^{i,N}(s)-\widehat{Y}_{\Delta}^{i,N}(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\rangle{\rm d}s$
$+(\overline{p}-2)\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),\sigma(\overline{\mu}^{Y,N}(u))\rangle{\rm d}B_{u}^{H,i}$
×$\times\langle\widetilde{Y}_{\Delta}^{i,N}(s)-\widehat{Y}_{\Delta}^{i,N}(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\rangle{\rm d}s$
=:absent:=:= : $(\overline{p}-2)G_{321}(t)+(\overline{p}-2)(\overline{p}-4)G_{322}(t)+(\overline{p}-2)G_{323}(t).$
By Lemma 5.1, the Young’s inequality, we get
$G_{321}(t)\leq$ $\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u))\rangle{\rm d}u$
×$\times\bigg{\langle}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u)){\rm d}u,b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\bigg{\rangle}{\rm d}s$
$+\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u))\rangle{\rm d}u$
×$\times\bigg{\langle}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}\sigma(\overline{\mu}^{Y,N}(u)){\rm d}B_{u}^{H,i},b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\bigg{\rangle}{\rm d}s$
$\leq$ $\Delta\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-3}{2}}|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))|^{3}{\rm d}u{\rm d}s$
$+C\mathbb{E}\int_{0}^{t}\bigg{[}\bigg{(}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-3}{2}}|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u),\overline{\mu}^{Y,N}(u))|{\rm d}u$
×$\times|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))|\bigg{)}^{\frac{\overline{p}}{\overline{p}-1}}+\bigg{\lvert}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}\sigma(\overline{\mu}^{Y,N}(u)){\rm d}B_{u}^{H,i}\bigg{\rvert}^{\overline{p}}\bigg{]}{\rm d}s$
$\leq$ $C\Delta^{2-3\alpha}\mathbb{E}\int_{0}^{t}(1+|Y_{\Delta}^{i,N}(s)|^{\overline{p}}+|Y_{\Delta}^{i,N}(s-\tau)|^{\overline{p}}+|\overline{Y}_{\Delta}^{i,N}(s)|^{p}+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|^{\overline{p}}){\rm d}s$
$+C\Delta^{2-3\alpha}\Delta^{(1-\alpha)\overline{p}}\mathbb{E}\int_{0}^{t}(1+|Y_{\Delta}^{i,N}(s)|^{\overline{p}}+|Y_{\Delta}^{i,N}(s-\tau)|^{p}+|\overline{Y}_{\Delta}^{i,N}(s)|^{\overline{p}}$
$+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|^{\overline{p}}){\rm d}s$
$+C\mathbb{E}\int_{0}^{t}\bigg{[}\bigg{(}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-3}{2}}|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))|^{2}{\rm d}u|\bigg{)}^{\frac{\overline{p}}{\overline{p}-1}}$
$+\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}\lvert\sigma(\overline{\mu}^{Y,N}(u))\rvert^{\overline{p}}{\rm d}u{\rm d}s$
$\leq$ $C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(s)|^{\overline{p}}{\rm d}s.$
Similarly, we get
$G_{322}(t)\leq C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(s)|^{\overline{p}}{\rm d}s.$
By Assumption 2.4, Lemma 5.1 and [32, Lemma 3.1.3], noting that $\mathbb{E}\int_{0}^{t}F_{s}{\rm d}B_{s}^{H,i}=0$, we get
$G_{323}(t)=$ $\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),\sigma(\overline{\mu}^{Y,N}(u))\rangle{\rm d}B_{u}^{H,i}$
×$\times\bigg{\langle}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u)){\rm d}u,b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\bigg{\rangle}{\rm d}s$
$+\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),\sigma(\overline{\mu}^{Y,N}(u))\rangle{\rm d}B_{u}^{H,i}$
×$\times\bigg{\langle}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}\sigma(\overline{\mu}^{Y,N}(u)){\rm d}B_{u}^{H,i},b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\bigg{\rangle}{\rm d}s$
$=$ $\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),\sigma(\overline{\mu}^{Y,N}(u))\rangle{\rm d}B_{u}^{H,i}$
×$\times\bigg{\langle}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}\sigma(\overline{\mu}^{Y,N}(u)){\rm d}B_{u}^{H,i},b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\bigg{\rangle}{\rm d}s$
$\leq$ $2HT^{2H-2}\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-3}{2}}|\sigma(\overline{\mu}^{Y,N}(u))|^{2}{\rm d}u$
×$\times|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))|{\rm d}s$
$\leq$ $C\Delta^{1-\alpha}\mathbb{E}\int_{0}^{t}(|Y_{\Delta}^{i,N}(s)|^{\overline{p}}+|Y_{\Delta}^{i,N}(s-\tau)|^{\overline{p}}+|\overline{Y}_{\Delta}^{i,N}(s))|^{\overline{p}}){\rm d}s$
$+C\Delta^{1-\alpha}\Delta^{(1-\alpha)\overline{p}}\mathbb{E}\int_{0}^{t}|\overline{Y}_{\Delta}^{i,N}(s))|^{\overline{p}}{\rm d}s$
$\leq$ $C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(s)|^{\overline{p}}{\rm d}s.$
To sum up, we derive
$G_{3}(t)\leq C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(s)|^{\overline{p}}{\rm d}s.$
Combining $G_{1}-G_{3}$, we get
(5.12) $\sup_{0\leq u\leq t}\mathbb{E}|\widetilde{Y}_{\Delta}^{i,N}(u)|^{\overline{p}}\leq\sup_{0\leq u\leq t}\mathbb{E}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\overline{p}/2}\leq C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)|^{\overline{p}}{\rm d}s.$
Noting that $|x-y|^{\overline{p}}\geq 2^{1-\overline{p}}|x|^{\overline{p}}-|y|^{\overline{p}}$, by Assumption 2.3, Lemma 5.1, we get
$|\widetilde{Y}_{\Delta}^{i,N}(u)|^{\overline{p}}\geq$ $2^{1-\overline{p}}|Y_{\Delta}^{i,N}(t)-D(Y_{\Delta}^{i,N}(t-\tau))|^{\overline{p}}-|\theta b_{\Delta}(Y_{\Delta}^{i,N}(t),Y_{\Delta}^{i,N}(t-\tau),\mu^{Y,N}(t))\Delta|^{\overline{p}}$
$\geq$ $2^{1-\overline{p}}[2^{1-\overline{p}}|Y_{\Delta}^{i,N}(t)|^{\overline{p}}-|D(Y_{\Delta}^{i,N}(t-\tau))|^{\overline{p}}]-|\theta b_{\Delta}(Y_{\Delta}^{i,N}(t),Y_{\Delta}^{i,N}(t-\tau),\mu^{Y,N}(t))\Delta|^{\overline{p}}$
$\geq$ $(2^{2-2\overline{p}}-4^{\overline{p}-1}\theta^{\overline{p}}K_{7}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}})|Y_{\Delta}^{i,N}(t)|^{\overline{p}}-(2^{1-\overline{p}}\lambda^{\overline{p}}+4^{\overline{p}-1}\theta^{p}K_{7}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}})|Y_{\Delta}^{i,N}(t-\tau)|^{\overline{p}}$
$-4^{\overline{p}-1}\theta^{\overline{p}}K_{7}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}}(\mathbb{W}_{q}(\mu^{Y,N}(t)),\delta_{0})^{\overline{p}}-4^{\overline{p}-1}\theta^{\overline{p}}K_{7}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}}.$
Thus, we obtain
$\sup_{0\leq u\leq t}\mathbb{E}|Y_{\Delta}^{i,N}(t)|^{\overline{p}}\leq(2^{2-2\overline{p}}-3\times 4^{\overline{p}-1}\theta^{\overline{p}}K_{7}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}}-2^{1-\overline{p}}\lambda^{\overline{p}})^{-1}$
×$\times\bigg{[}\sup_{0\leq u\leq t}\mathbb{E}|\widetilde{Y}_{\Delta}^{i,N}(u)|^{\overline{p}}+4^{\overline{p}-1}\theta^{\overline{p}}K_{7}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}}+(2^{1-\overline{p}}\lambda^{\overline{p}}+4^{\overline{p}-1}\theta^{\overline{p}}K_{7}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}})\mathbb{E}\parallel\xi\parallel^{\overline{p}}\bigg{]}.$
Then, we derive from (5.12),
$\sup_{0\leq u\leq t}\mathbb{E}|Y_{\Delta}^{i,N}(t)|^{\overline{p}}\leq C+C\int_{0}^{t}\sup_{0\leq u\leq t}\mathbb{E}|Y_{\Delta}^{i,N}(t)|^{\overline{p}}.$
Applying the Gronwall inequality, we complete the proof. ∎
5.2 Strong convergence rate
Lemma 5.5.
Let Assumptions 2.2-2.5 hold. For $p>1/H$, then
$\mathbb{E}\bigg{(}\sup_{0\leq k\leq M-1}\sup_{t_{k}\leq t\leq t_{k+1}}|Y_{\Delta}^{i,N}(t)-Y_{\Delta}^{i,N}(t_{k})|^{p}\bigg{)}\leq C\Delta^{[(1-\alpha)\wedge H]p},$
where $C$ is a positive constant independent of $\Delta$.
Proof.
For $t\in[t_{k},t_{k+1}]$, by the elementary inequality, we may compute
$\mathbb{E}\bigg{(}\sup_{t_{k}\leq t\leq t_{k+1}}|\widetilde{Y}_{\Delta}^{i,N}(t)-\widetilde{Y}_{\Delta}^{i,N}(t_{k})|^{p}\bigg{)}$
$=$ $\mathbb{E}\bigg{(}\sup_{t_{k}\leq t\leq t_{k+1}}\bigg{\lvert}\int_{t_{k}}^{t}b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s)){\rm d}s+\int_{t_{k}}^{t}\sigma(\overline{\mu}^{Y,N}(s)){\rm d}B_{s}^{H,i}\bigg{\rvert}^{p}\bigg{)}$
$\leq$ $2^{p-1}\mathbb{E}\bigg{(}\sup_{t_{k}\leq t\leq t_{k+1}}\bigg{\lvert}\int_{t_{k}}^{t}b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s)){\rm d}s\bigg{\rvert}^{p}\bigg{)}$
$+2^{p-1}\mathbb{E}\bigg{(}\sup_{t_{k}\leq t\leq t_{k+1}}\bigg{\lvert}\int_{t_{k}}^{t}\sigma(\overline{\mu}^{Y,N}(s)){\rm d}B_{s}^{H,i}\bigg{\rvert}^{p}\bigg{)}$
=:absent:=:= : $J_{1}+J_{2}.$
By the Hölder inequality and Lemmas 5.1, 5.4, we get
$J_{1}\leq$ $2^{p-1}\Delta^{p-1}\mathbb{E}\int_{t_{k}}^{t_{k+1}}|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))|^{p}{\rm d}s$
$\leq$ $2^{p-1}\Delta^{p-1}\mathbb{E}\int_{t_{k}}^{t_{k+1}}K_{6}\Delta^{-\alpha p}(1+|\overline{Y}_{\Delta}^{i,N}(s)|+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|+\mathbb{W}_{q}(\overline{\mu}^{Y,N}(s),\delta_{0}))^{p}{\rm d}s$
$\leq$ $C\Delta^{(1-\alpha)p}+C\Delta^{(1-\alpha)p}\mathbb{E}\int_{t_{k}}^{t_{k+1}}(|\overline{Y}_{\Delta}^{i,N}(s)|^{p}+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|^{p}){\rm d}s$
$\leq$ $C\Delta^{(1-\alpha)p}.$
Using the same technique as in (3.3), along with the Hölder inequality, the Fubini theorem and Lemmas 2.5, 5.4, we get
$J_{2}\leq C\Delta^{pH-1}\int_{t_{k}}^{t_{k+1}}\mathbb{E}(1+\mathbb{W}_{\theta}(\overline{\mu}_{s}^{Y,N},\delta_{0}))^{p}{\rm d}s\leq C\Delta^{pH}.$
Recalling that $\alpha\in(0,1/2]$, we obtian
$\mathbb{E}\bigg{(}\sup_{t_{k}\leq t\leq t_{k+1}}|\widetilde{Y}_{\Delta}^{i,N}(t)-\widetilde{Y}_{\Delta}^{i,N}(t_{k})|^{p}\bigg{)}\leq C\Delta^{[(1-\alpha)\wedge H]p}.$
Denoting by $\widetilde{D}(t,t_{k}):=D(Y_{\Delta}^{i,N}(t-\tau))-D(Y_{\Delta}^{i,N}(t_{k}-\tau))$, and $\widetilde{b}(t,t_{k}):=b_{\Delta}(Y_{\Delta}^{i,N}(t),Y_{\Delta}^{i,N}(t-\tau),\mu^{Y,N}(t))-b_{\Delta}(Y_{\Delta}^{i,N}(t_{k}),Y_{\Delta}^{i,N}(t_{k}-\tau),\mu^{Y,N}(t_{k}))$, together with Assumption 2.3, we get
(5.13) $|\widetilde{Y}_{\Delta}^{i,N}(t)-\widetilde{Y}_{\Delta}^{i,N}(t_{k})|^{p}$
$\geq$ $2^{1-p}|Y_{\Delta}^{i,N}(t)-Y_{\Delta}^{i,N}(t_{k})-\widetilde{D}(t,t_{k})|^{p}-\theta^{p}\Delta^{p}|\widetilde{b}(t,t_{k})|^{p}$
$\geq$ $2^{2-2p}|Y_{\Delta}^{i,N}(t)-Y_{\Delta}^{i,N}(t_{k})|^{p}-2^{1-p}\lambda^{p}|Y_{\Delta}^{i,N}(t-\tau)-Y_{\Delta}^{i,N}(t_{k}-\tau)|^{p}-\theta^{p}\Delta^{p}|\widetilde{b}(t,t_{k})|^{p}.$
For $0\leq t\leq t_{1}=\Delta$, then $t-\tau<t_{1}-\tau<0$, we derive from (5.13) and Lemma 5.4,
$\mathbb{E}\bigg{(}\sup_{0\leq t\leq t_{1}}|Y_{\Delta}^{i,N}(t)-Y_{\Delta}^{i,N}(t_{0})|^{p}\bigg{)}\leq C\mathbb{E}\bigg{(}\sup_{0\leq t\leq t_{1}}|\widetilde{Y}_{\Delta}^{i,N}(t)-\widetilde{Y}_{\Delta}^{i,N}(t_{0})|^{p}\bigg{)}+C\Delta^{(1-\alpha)p}\leq C\Delta^{[(1-\alpha)\wedge H]p}.$
Similarly, for $t_{1}\leq t<t_{2}$, we obtain
$\mathbb{E}\bigg{(}\sup_{t_{1}\leq t\leq t_{2}}|Y_{\Delta}^{i,N}(t)-Y_{\Delta}^{i,N}(t_{1})|^{p}\bigg{)}$
$\leq$ $C\mathbb{E}\bigg{(}\sup_{t_{1}\leq t\leq t_{2}}|\widetilde{Y}_{\Delta}^{i,N}(t)-\widetilde{Y}_{\Delta}^{i,N}(t_{1})|^{p}\bigg{)}+C\mathbb{E}\bigg{(}\sup_{0\leq t\leq(t_{2-m}\vee 0)}|\widetilde{Y}_{\Delta}^{i,N}(t)-\widetilde{Y}_{\Delta}^{i,N}(t_{1-m})|^{p}\bigg{)}+C\Delta^{(1-\alpha)p}$
$\leq$ $C\Delta^{[(1-\alpha)\wedge H]p}.$
By induction,
$\mathbb{E}\bigg{(}\sup_{t_{k}\leq t\leq t_{k+1}}|Y_{\Delta}^{i,N}(t)-Y_{\Delta}^{i,N}(t_{1})|^{p}\bigg{)}\leq$ $C\mathbb{E}\bigg{(}\sup_{t_{k}\leq t\leq t_{k+1}}|\widetilde{Y}_{\Delta}^{i,N}(t)-\widetilde{Y}_{\Delta}^{i,N}(t_{k})|^{p}\bigg{)}+C\Delta^{(1-\alpha)p}$
$\leq$ $C\Delta^{[(1-\alpha)\wedge H]p}.$
we get the desired assertion. ∎
Lemma 5.6.
Let Assumptions 2.2-2.5 and Lemmas 5.1, 5.2 hold, then it holds that for $p>1/H$ and $\overline{p}>2(l+1)p$,
$\mathbb{E}\bigg{(}\sup_{0\leq t\leq T}|X_{t}^{i,N}-Y_{\Delta}^{i,N}(t)|^{p}\bigg{)}\leq C\Delta^{\alpha p},$
where $C$ is a positive constant independent of $\Delta$.
Proof.
For $t\in[0,T]$, denote $I(t)=\widetilde{Y}_{\Delta}^{i,N}(t)-X_{t}^{i,N}+D(X_{t-\tau}^{i,N})$, then
$I(t)=$ $I(0)+\int_{0}^{t}(b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})){\rm d}s$
$+\int_{0}^{t}(\sigma(\overline{\mu}^{Y,N}(s))-\sigma(\mu_{s}^{X,N})){\rm d}B_{s}^{H,i},$
where $I_{0}=-\theta b_{\Delta}(\xi(0),\xi(-\tau),\delta_{0})\Delta$. The Itô formula yields,
$|I(t)|^{p}\leq$ $|I(0)|^{p}$
$+p\int_{0}^{t}|I(s)|^{p-2}\langle I(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})\rangle{\rm d}s$
$+HT^{2H-1}p(p-1)\int_{0}^{t}|I(s)|^{p-2}|\sigma(\overline{\mu}^{Y,N}(s))-\sigma(\mu_{s}^{X,N})|^{2}{\rm d}s$
$+p\int_{0}^{t}|I(s)|^{p-2}\langle I(s),\sigma(\overline{\mu}^{Y,N}(s))-\sigma(\mu_{s}^{X,N})\rangle{\rm d}B_{s}^{H,i}$
$\leq$ $|I(0)|^{p}+\sum_{i=1}^{6}H_{i}(t),$
where
$H_{1}(t)=$ $p\int_{0}^{t}|I(s)|^{p-2}\langle I(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))$
$-b(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))\rangle{\rm d}s$
$H_{2}(t)=$ $p\int_{0}^{t}|I(s)|^{p-2}\langle I(s),b(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))$
$-b(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))\rangle{\rm d}s$
$H_{3}(t)=$ $p\int_{0}^{t}|I(s)|^{p-2}\langle I(s),b(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))$
$-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})\rangle{\rm d}s,$
$H_{4}(t)=$ $HT^{2H-1}p(p-1)\int_{0}^{t}|I(s)|^{p-2}|\sigma(\overline{\mu}^{Y,N}(s))-\sigma(\mu^{Y,N}(s))|^{2}{\rm d}s$
$H_{5}(t)=$ $HT^{2H-1}p(p-1)\int_{0}^{t}|I(s)|^{p-2}|\sigma(\mu^{Y,N}(s))-\sigma(\mu_{s}^{X,N})|^{2}{\rm d}s$
$H_{6}(t)=$ $p\int_{0}^{t}|I(s)|^{p-2}\langle I(s),\sigma(\overline{\mu}^{Y,N}(s))-\sigma(\mu_{s}^{X,N})\rangle{\rm d}B_{s}^{H,i}.$
By Assumption 2.3, Lemmas 5.1, 5.2, 5.4, and the Young’s inequality, the Hölder inequality, we get
$\mathbb{E}\bigg{(}\sup_{0\leq u\leq t}H_{1}(u)\bigg{)}$
$\leq$ $C\mathbb{E}\int_{0}^{t}|I(s)|^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))$
$-b(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))|^{p}{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}\bigg{(}|Y_{\Delta}^{i,N}(s)-X_{s}^{i,N}|^{p}+|Y_{\Delta}^{i,N}(s-\tau)-X_{s-\tau}^{i,N}|^{p}$
$+\theta^{p}\Delta^{p}|b_{\Delta}(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))|^{p}\bigg{)}{\rm d}s$
$+C\Delta^{\alpha p}\mathbb{E}\int_{0}^{t}(1+|\overline{Y}_{\Delta}^{i,N}(s)|^{2(l+1)p}+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|^{2(l+1)p}+(\mathbb{W}_{q}(\overline{\mu}^{Y,N},\delta_{0}))^{2p}){\rm d}s$
$\leq$ $C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(s)-X_{s}^{i,N}|^{p}+C\Delta^{(1-\alpha)p}+C\Delta^{\alpha p},$
where for the last term we used the fact that $\overline{p}>2(l+1)p$. By the elementary inequality, the Young’s inequality, the Hölder inequality, Assumptions 2.3, 2.4 and Lemmas 5.1, 5.4, 5.5, we get
$\mathbb{E}\bigg{(}\sup_{0\leq u\leq t}H_{2}(u)\bigg{)}$
$\leq$ $C\mathbb{E}\int_{0}^{t}|I(s)|^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}|b(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))-b(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))|^{p}{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}\bigg{(}|Y_{\Delta}^{i,N}(s)-X_{s}^{i,N}|^{p}+|Y_{\Delta}^{i,N}(s-\tau)-X_{s-\tau}^{i,N}|^{p}+\theta^{p}\Delta^{p}|b_{\Delta}(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))|\bigg{)}{\rm d}s$
$+C\mathbb{E}\int_{0}^{t}\bigg{[}(1+|\overline{Y}_{\Delta}^{i,N}(s)|^{l}+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|^{l}+|Y_{\Delta}^{i,N}(s)|^{l}+|Y_{\Delta}^{i,N}(s-\tau)|^{l})^{p}$
×$\times(|\overline{Y}_{\Delta}^{i,N}(s)-Y_{\Delta}^{i,N}(s)|+|\overline{Y}_{\Delta}^{i,N}(s-\tau)-Y_{\Delta}^{i,N}(s-\tau)|+(\mathbb{W}_{q}(\overline{\mu}^{Y,N}(s),\mu^{Y,N})(s)))^{p}\bigg{]}{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}\bigg{(}|Y_{\Delta}^{i,N}(s)-X_{s}^{i,N}|^{p}+|Y_{\Delta}^{i,N}(s-\tau)-X_{s-\tau}^{i,N}|^{p}\bigg{)}{\rm d}s+C\Delta^{(1-\alpha)p}$
$+C\int_{0}^{t}[\mathbb{E}(1+|\overline{Y}_{\Delta}^{i,N}(s)|^{l}+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|^{l}+|Y_{\Delta}^{i,N}(s)|^{l}+|Y_{\Delta}^{i,N}(s-\tau)|^{l})^{2p}]^{1/2}$
×$\times[\mathbb{E}(|\overline{Y}_{\Delta}^{i,N}(s)-Y_{\Delta}^{i,N}(s)|+|\overline{Y}_{\Delta}^{i,N}(s-\tau)-Y_{\Delta}^{i,N}(s-\tau)|)^{2p}]^{1/2}{\rm d}s$
$+C\int_{0}^{t}\mathbb{E}(|\overline{Y}_{\Delta}^{i,N}(s)-Y_{\Delta}^{i,N}(s)|)^{p}{\rm d}s$
$\leq$ $C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}{\rm d}s+C\Delta^{(1-\alpha)p}+C\Delta^{[(1-\alpha)\wedge H]p}.$
By Assumption 2.4 and Lemmas 5.1, 5.4, we get
$\mathbb{E}\bigg{(}\sup_{0\leq u\leq t}H_{3}(u)\bigg{)}$
$\leq$ $C\mathbb{E}\int_{0}^{t}|I(s)|^{p-2}[|Y_{\Delta}^{i,N}(s)-X_{s}^{i,N}|^{2}+|Y_{\Delta}^{i,N}(s-\tau)-X_{s-\tau}^{i,N}|^{2}+(\mathbb{W}_{q}(\mu^{Y,N}(s),\mu^{X,N}(s)))^{2}]{\rm d}s$
$+C\mathbb{E}\int_{0}^{t}|I(s)|^{p-2}|\theta b_{\Delta}(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))\Delta||b(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))$
$-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})|{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}[|Y_{\Delta}^{i,N}(s)-X_{s}^{i,N}|^{p}+\theta^{p}\Delta^{p}|b_{\Delta}(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))|^{p}]{\rm d}s$
$+C\Delta^{1-\alpha}\mathbb{E}\int_{0}^{t}|I(s)|^{p-2}[(1+|Y_{\Delta}^{i,N}(s)|^{l}+|Y_{\Delta}^{i,N}(s-\tau)|^{l}+|X_{s}^{i,N}|^{l}+|X_{s-\tau}^{i,N}|^{l})$
×$\times(|Y_{\Delta}^{i,N}(s)-X_{s}^{i,N}|+|Y_{\Delta}^{i,N}(s-\tau)-X_{s-\tau}^{i,N}|)+\mathbb{W}_{q}(\mu^{Y,N}(s),\mu_{s}^{X,N})]{\rm d}s$
$\leq$ $C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}{\rm d}s+C\Delta^{(1-\alpha)(1+p)}.$
By the Young’s inequality, Assumption 2.5, we get
$\mathbb{E}\bigg{(}\sup_{0\leq u\leq t}H_{4}(u)\bigg{)}+\mathbb{E}\bigg{(}\sup_{0\leq u\leq t}H_{5}(u)\bigg{)}$
$\leq$ $C\mathbb{E}\int_{0}^{t}|I(s)|^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}(\mathbb{W}_{q}(\overline{\mu}^{Y,N}(s)),\mu^{Y,N}(s))^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}(\mathbb{W}_{q}(\mu^{Y,N}(s)),\mu_{s}^{X,N})^{p}{\rm d}s$
$\leq$ $C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}{\rm d}s+C\Delta^{(1-\alpha)p}+C\Delta^{[(1-\alpha)\wedge H]p}.$
In the way in (3.3), by the Young’s inequality, we get
$\mathbb{E}\bigg{(}\sup_{0\leq u\leq t}|H_{6}(u)|^{p}\bigg{)}$
$\leq$ $C\mathbb{E}\int_{0}^{t}|I(s)|^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}(\mathbb{W}_{q}(\overline{\mu}^{Y,N}(s)),\mu^{Y,N}(s))^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}(\mathbb{W}_{q}(\mu^{Y,N}(s)),\mu_{s}^{X,N})^{p}{\rm d}s$
$\leq$ $C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}{\rm d}s+C\Delta^{(1-\alpha)p}+C\Delta^{[(1-\alpha)\wedge H]p}.$
Combining $H_{1}-H_{6}$, we get
$\mathbb{E}\bigg{(}\sup_{0\leq u\leq t}|I(u)|^{p}\bigg{)}\leq C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{0\leq u\leq s}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}\bigg{)}{\rm d}s+C\Delta^{\alpha p}.$
Recalling the definition of $I(t)$ and by Assumption 2.3, we may compute
$|I(t)|^{p}\geq$ $2^{1-p}|Y_{\Delta}^{i,N}(t)-X_{t}^{i,N}-D(Y_{\Delta}^{i,N}(t-\tau))+D(X_{t-\tau}^{i,N})|^{p}$
$-|\theta b_{\Delta}(Y_{\Delta}^{i,N}(t),Y_{\Delta}^{i,N}(t-\tau),\mu^{Y,N}(t))\Delta|^{p}$
$\geq$ $2^{1-p}[2^{1-p}|Y_{\Delta}^{i,N}(t)-X_{t}^{i,N}|^{p}-|D(Y_{\Delta}^{i,N}(t-\tau))-D(X_{t-\tau}^{i,N})|^{p}]$
$-|\theta b_{\Delta}(Y_{\Delta}^{i,N}(t),Y_{\Delta}^{i,N}(t-\tau),\mu^{Y,N}(t))\Delta|^{p}$
$\geq$ $2^{2-2p}|Y_{\Delta}^{i,N}(t)-X_{t}^{i,N}|^{p}-2^{1-p}\lambda^{p}|Y_{\Delta}^{i,N}(t-\tau)-X_{t-\tau}^{i,N}|^{p}$
$-|\theta b_{\Delta}(Y_{\Delta}^{i,N}(t),Y_{\Delta}^{i,N}(t-\tau),\mu^{Y,N}(t))\Delta|^{p}.$
By Lemma 5.1, 5.4, we derive
$\mathbb{E}\bigg{(}\sup_{0\leq u\leq t}|X_{u}^{i,N}-Y_{\Delta}^{i,N}(u)|^{p}\bigg{)}\leq$ $\mathbb{E}\bigg{(}\sup_{0\leq u\leq t}|I(u)|^{p}\bigg{)}+C\Delta^{(1-\alpha)p}$
$\leq$ $C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{0\leq u\leq s}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}\bigg{)}{\rm d}s+C\Delta^{\alpha p}.$
Applying the Gronwall inequality, we can derive the assertion. ∎
Theorem 5.7.
Let Assumptions 2.2-2.5 hold. For $p>1/H$ and $\overline{p}\geq 2(l+1)p$, it holds that
$\mathbb{E}\bigg{(}\sup_{t\in[0,T]}|X_{t}^{i}-Y_{t}^{i.N}|^{p}\bigg{)}\leq C\times\left\{\begin{array}[]{ll}N^{-\frac{1}{2}}+\Delta^{\alpha p},\ \ p>\frac{d}{2},\\ N^{-\frac{1}{2}}\log(1+N)+\Delta^{\alpha p},\ p=\frac{d}{2},\\ N^{-\frac{p}{d}}+\Delta^{\alpha p},\ p\in[2,\frac{d}{2}),\end{array}\right.$
where $C$ is a constant independent of $N$ and $\Delta$.
Proof.
By Lemmas 4.3 and 5.6, the result is obvious. ∎

### v2

@@
-Li Tana,b , Shengrong Wanga,c
+Strong convergence of tamed theta scheme for superlinearly growing McKean-Vlasov NSDDEs driven by fractional Brownian motionsthanks: Supported by Jiangxi Provincial Natural Science Foundation(No., 20252BAC250002, GJJ2404201)
 Abstract
-In this article, we propose a tamed theta Euler-Maruyama (EM) scheme for superlinearly growing neutral McKean-Vlasov stochastic differential delay equations driven by fractional Brownian motion with Hurst exponent $H\in(1/2,1)$.
-The analytical properties including uniqueness and existence, boundedness of moment and propagation of chaos are investigated.
+In this article, we study the McKean-Vlasov neutral stochastic differential delay equations driven by fractional Brownian motion with super-linearly growing coefficients, where the Hurst exponent $H\in(1/2,1)$.
+The existence and uniqueness of the exact solution were shown by the Picard iteration.
+Besides, we propose a tamed theta Euler-Maruyama scheme for this equation, analyzed the moment boundness and propagation of chaos etc.
 Moreover, the convergence rate of the numerical scheme is established.
-Keywords :Neutral McKean-Vlasov SDEs; superlinearly growing; fractional Brownian motion; strong convergence rate; tamed theta EM
+Keywords : McKean-Vlasov NSDDEs; super-linearly growing coefficients; fractional Brownian motion; strong convergence rate; tamed theta EM scheme
 1 introduction
-Stochastic differential equations (SDEs) with coefficients that depend on the law of the solution were first studied by McKean [1], who was inspired by Kac’s programme in Kinetic Theory [2].
-Subsequently, they have become a research hotspot.
-Bossy and Talay [3, 4] chose to approximate the McKean-Vlasov limit by replicating the behavior with a system of $n$ weakly interacting particles.
-They also studied the numerical approximation of the solutions to the McKean-Vlasov SDEs.
-The existence and uniqueness of McKean-Vlasov SDEs were discussed in [5, 6].
-Lacker [7] investigated the well-posedness and the propagation of chaos for McKean-Vlasov SDEs.
-Since the explicit solutions of McKean-Vlasov SDEs are difficult to obtain, many types of McKean-Vlasov SDEs can be approximated by implicit Euler-Maruyama (EM) scheme [8, 16], adaptive EM scheme [9, 11], tamed EM scheme [10, 14], truncated EM scheme [15] and so on.
-They approximated the solutions of McKean-Vlasov SDEs using numerical methods whose coefficients are globally (or locally) Lipschitz continuous or exhibit super-linear growth.
-Recently, there has been a tendency for research to focus on the McKean-Vlasov SDEs driven by fractional Brownian motion. The form is
+Stochastic differential equations (SDEs) whose coefficients depend on the law of the solution (also called McKean-Vlasov SDEs), were first investigated by McKean [1], whose work drew inspiration from Kac’s programme in Kinetic Theory [2].
+Since then, McKean-Vlasov SDEs have evolved into a major area of research in stochastic analysis.
+To approximate the McKean-Vlasov SDE limit, Bossy and Talay [3, 4] proposed a system of $n$ weakly interacting particles—replicating the statistical behavior of the limit system, and systematically studied the numerical approximation of its solutions.
+Building on this foundation, the existence and uniqueness of McKean-Vlasov SDEs were discussed in [5, 6], while the well-posedness and the propagation of chaos for McKean-Vlasov SDEs were investigated by Lacker [7].
+Given the rarity of explicit solutions for McKean-Vlasov SDEs, numerical approximation methods have become indispensable.
+A variety of schemes have been developed to handle different coefficients behaviors, for example the implicit Euler-Maruyama (EM) scheme [8, 9], the adaptive EM scheme [10, 11], the tamed EM scheme [12, 13], and truncated EM scheme [14].
+In recent years, research attention has increasingly shifted to McKean-Vlasov SDEs driven by fractional Brownian motion—a generalization of standard Brownian motion with unique self-similarity and long-memory properties, making it well-suited for modeling phenomena in atmospheric sciences, financial engineering, and other fields.
+These equations typically take the form:
 (1.1) ${\rm d}X_{t}=b(X_{t},\mathcal{L}(X_{t})){\rm d}t+\sigma(\mathcal{L}(X_{t})){\rm d}B_{t}^{H},$
-where the coefficients $b:\mathbb{R}^{d}\times\mathcal{P}_{q}(\mathbb{R}^{d})\to\mathbb{R}^{d},\sigma:\mathcal{P}_{q}(\mathbb{R}^{d})\to\mathbb{R}^{d}\otimes\mathbb{R}^{d}$ are Borel-measurable, here, $\mathcal{P}_{q}(\mathbb{R}^{d})$ is the set of probability measures on $\mathbb{R}^{d}$ with finite $q$-th moment.
+where the coefficients $b:\mathbb{R}^{d}\times\mathcal{P}_{\theta}(\mathbb{R}^{d})\to\mathbb{R}^{d},\sigma:\mathcal{P}_{\theta}(\mathbb{R}^{d})\to\mathbb{R}^{d}\otimes\mathbb{R}^{d}$ are Borel-measurable, here, $\mathcal{P}_{\theta}(\mathbb{R}^{d})$ is the set of probability measures on $\mathbb{R}^{d}$ with finite $\theta$-th moment.
 $\mathcal{L}(X_{t})$ is the law of $X$ at $t$ and $B^{H}=\{B_{t}^{H},t\in[0,T]\}$ is a fractional Brownian motion with Hurst exponent $H\in(0,1)$ defined on the probability space $(\Omega,\mathcal{F},\mathbb{P})$.
-Moreover, $B^{H}$ is a centered Gaussian process with the covariance function
-$R_{H}(t,s)=\mathbb{E}(B_{t}^{H}B_{s}^{H})=\frac{1}{2}(t^{2H}+s^{2H}-|t-s|^{2H}),\ \ \forall t,s\in[0,T].$
-When $H=1/2$, the fractional Brownian motion $\{B_{t}^{H}\}_{t\geq 0}$ corresponds to a standard Bromnian motion.
-The self-similarity and long memory properties of fractional Brownian motion make it applicable in various fields, such as atmospheric sciences, financial engineering and others.
-From then on, scholars conducted studies on (1.1) separately.
-He et al.
-[17] established the theory of propagation of chaos and studied the strong convergence of the EM method when the coefficients of (1.1) satisfy the globally Lipschitz condition.
-However, few studies considered the case where the coefficients of (1.1) grow super-linearly.
-Shen [18] demonstrated that, under certain averaging conditions, the solutions of McKean-Vlasov SDEs driven by fractional Brownian motion can be approximated by the solutions of the associated averaged McKean-Vlasov SDEs driven by fractional Brownian motion, in the sense of the mean square convergence, where the coefficients exhibit super-linear growth.
-Wang et al.
+As a centered Gaussian process, $B_{H}$ has the covariance function
+(1.2) $R_{H}(t,s)=\mathbb{E}(B_{t}^{H}B_{s}^{H})=\frac{1}{2}(t^{2H}+s^{2H}-|t-s|^{2H}),\ \ \forall t,s\in[0,T].$
+Notably, when $H=1/2$, the fractional Brownian motion $\{B_{t}^{H}\}_{t\geq 0}$ reduces to a standard Brownian motion.
+The well-posedness of such equations was first proven by Fan et al.[15].
+Subsequent research have further expanded the theoretical framework of this model.
+For example, under the assumption that the coefficients satisfy globally Lipschitz conditions, He et al.
+[16] established the propagation of chaos for the model described in (1.1) and analyzed the strong convergence of the EM method.
+A critical gap in current literature is the lack of studies on McKean-Vlasov SDEs driven by fractional Brownian motion with super-linearly growing coefficients.
+One notable exception is Shen et al.[17] who showed that under specific averaging conditions, solutions to such equations can be approximated by solutions of an associated averaged equation.
+More recently, Wang et al.
 [18] and Gao et al.
-[20] focus on the EM scheme for delayed McKean-Vlasov SDEs driven by fractional Brownian motion, whose drift coefficient contains a polynomial delay term.
-Actually, in order to describe the impact of the past, stochastic delay differential equations (SDDEs) and neutral stochastic differential delay equations (NSDDEs) play a significant role in stochastic calculus, describing a wide variety of natural and man-made systems (see, e.g., [21, 22, 23]).
-The classical EM method is no longer applicable, as the coefficients of SDEs do not exhibit linear growth.
-It is known that the tamed EM shceme can be used to deal with superlinearly growing coefficients.
-Deng et al.
-[24] proposed two types of explicit tamed EM schemes for NSDDEs with superlinearly growing drift and diffusion coefficients and proved strong convergence for each of them, respectively.
+[19] focused on the EM scheme for delayed McKean-Vlasov SDEs driven by fractional Brownian motion, but their work only considered super-linear growth in the delay term of the drift coefficient, not the coefficient itself.
+To capture the influence of past states on current dynamics, stochastic delay differential equations (SDDEs) and neutral stochastic differential delay equations (NSDDEs) are essential tools in stochastic calculus, with broad applications in modeling natural and engineered systems, see for example [20, 21, 22].
+A key challenge arises when NSDDEs exhibit super-linear growth: the classical EM scheme often fails to converge.
+To address this, a tamed EM scheme was introduced to handle super-linearly growing coefficients.
+For instance, Deng et al.
+[23] proposed two types of explicit tamed EM schemes for NSDDEs with super-linearly growing drift and diffusion coefficients and proved strong convergence for each.
 Cui et al.
-[25] studied the tamed EM methods to approximate the solutions for a type of of McKean-Vlasov NSDDEs in which the drift coefficients exhibit superlinear growth, and obtained the rate of convergence.
-We can refer to [23, 27] for more detail about tamed EM scheme in NSDDEs.
-It is worth noting that Tan and Yuan [28] modified the tamed theta EM scheme and showed the convergence of the numerical solutions, which provides us with research ideas for conducting this paper.
-Drawing inspiration from the tamed approach introduced in [28] and previous models of McKean-Vlasov SDEs driven by fractional Brownian motion, we consider a tamed theta EM scheme for MV-NSDDEs driven by fractional Brownian motion with coefficients of superlinear growth, as an extension of (1.1).
-The form is
+[24] studied the tamed EM methods to approximate the solutions of McKean-Vlasov NSDDEs with super-linear drift, establishing the convergence rate.
+Additional insights into tamed EM schemes for NSDDEs can be found in [22, 25].
+Notably, Tan and Yuan [26] proposed a modified tamed theta EM scheme and demonstrated the convergence of the numerical solutions, providing a key motivation for the present work.
+Drawing inspiration from the existing literatures, we extend the model (1.1) to accommodate coefficients with super-linear growth.
+We propose a tamed theta EM scheme for the McKean-Vlasov NSDDEs, which takes the following form:
 (1.3) ${\rm d}(X_{t}-D(X_{t-\tau}))=b(X_{t},X_{t-\tau},\mathcal{L}(X_{t})){\rm d}t+\sigma(\mathcal{L}(X_{t})){\rm d}B_{t}^{H},$
-where $\tau>0$ is the delay.
-The structure of the rest of the paper is as follows.
-In Section 2, the mathematical preliminaries on the McKean-Vlasov NSDDEs with superlinear growth driven by fractional Brownian motion are presented.
-In Section 3, existence and uniqueness of the solution is given.
-In Section 4, the propagation of chaos for this model is also analyzed.
-In Section 5, the EM scheme is established and the strong convergence rate is given.
+where $\tau>0$ denotes the delay.
+This paper’s main contributions are outlined below:
+  • Unlike the prior work (e.g., [18, 19]) that only considers super-linear growth in the delay term, we analyze McKean-Vlasov equations driven by fractional Brownian motion where the drift coefficient itself exhibits super-linear growth.
+We prove the core properties of such equations, representing a significant advance in this field.
+  • By a delicate analysis involving the fractional Itô formula and intricate inequality techniques, we investigate the properties of the solution, the propagation of chaos in associated particle systems, and the convergence of the proposed numerical scheme.
+The rest of the paper is organized as follows.
+Section 2 presents the mathematical preliminaries, including key definitions and lemmas for McKean-Vlasov NSDDEs with super-linear growth driven by fractional Brownian motion.
+Section 3 establishes the existence and uniqueness of the solution.
+In Section 4, we analyze the propagation of chaos in the particle system associated with our model.
+Finally, Section 5 introduces the tamed theta EM scheme and derives its strong convergence rate.
 2 Preliminaries
-[…]
-$\mathcal{P}_{q}(\mathbb{R}^{d})=\bigg{\{}\mu\in\mathcal{P}(\mathbb{R}^{d}):\int_{\mathbb{R}^{d}}|x|^{q}\mu({\rm d}x)<\infty\bigg{\}},$
-which is a Polish space under the $L^{q}$-Wasserstein distance [30, Definition 6.1]. Then, for $\mu\in\mathcal{P}_{q}(\mathbb{R}^{d})$, define
-$\mathcal{W}_{q}(\mu)=\bigg{(}\int_{\mathbb{R}^{d}}|x|^{q}\mu({\rm d}x)\bigg{)}^{1/q}$
-For $\mu,\nu\in\mathcal{P}_{q}(\mathbb{R}^{d})$, we define the Wasserstein distance on $\mathcal{P}_{q}(\mathbb{R}^{d})$ by
-$\mathbb{W}_{q}(\mu,\nu)=\bigg{(}\inf_{\pi\in\Pi(\mu,\nu)}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}|x-y|^{q}\pi({\rm d}x,{\rm d}y)\bigg{)}^{1/q},$
-where $\Pi(\mu,\nu)$ is the set of probability measures on $\mathbb{R}^{d}\times\mathbb{R}^{d}$ with marginals $\mu$ and $\nu$.
-In other words, $\pi\in\Pi(\mu,\nu)$ is a probability measure on $\mathbb{R}^{d}\times\mathbb{R}^{d}$ such that $\pi(\cdot\times\mathbb{R}^{d})=\mu(\cdot)$ and $\pi(\cdot\times\mathbb{R}^{d})=\nu(\cdot)$.
-For $11/2<H<$ 1, we put
+Throughout this paper, let $(\Omega,\mathcal{F},\mathbb{P})$ be a complete probability space with a filtration $\{\mathcal{F}_{t}\}_{t\geq 0}$ satisfying the usual conditions (i.e., it is increasing and right continuous while $\mathcal{F}_{0}$ contains all $\mathbb{P}$-null sets).
+We use $\lvert\cdot\rvert$ and $\langle\cdot,\cdot\rangle$ for the Euclidean norm and inner product, respectively, and let $a\wedge b:=\min(a,b)$ and $a\vee b:=\max(a,b)$.
+Let $\tau>0$ be a constant and denote by $\mathcal{C}([-\tau,0];\mathbb{R}^{d})$ the space of all continuous functions $\varphi$ from $\tau,0]$ to $\mathbb{R}^{d}$ with the norm $\parallel\varphi\parallel=\sup_{-\tau\leq t\leq 0}|\varphi(t)|$.
+For $p>0$, $L_{\mathcal{F}_{0}}^{p}(\Omega;\mathbb{R}^{d})$ represents the family of $\mathcal{F}_{0}$-measurable random variable with $\mathbb{E}|\cdot|^{p}<\infty$.
+Let $\mathcal{P}_{\theta}(\mathbb{R}^{d}),\theta\geq 1$ be the space of probability measures on $\mathbb{R}^{d}$ with finite $\theta$-th moment.
+Formally, it is defined as
+$\mathcal{P}_{\theta}(\mathbb{R}^{d})=\bigg\{\mu\in\mathcal{P}(\mathbb{R}^{d}):\int_{\mathbb{R}^{d}}|x|^{\theta}\mu({\rm d}x)<\infty\bigg\},$
+where $\mathcal{P}(\mathbb{R}^{d})$ denotes the family of all probability measures on $(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))$, here, $\mathcal{B}(\mathbb{R}^{d})$ is the Borel $\sigma$-field over $\mathbb{R}^{d}$.
+For any $\mu\in\mathcal{P}_{\theta}(\mathbb{R}^{d})$, define the $\theta$-th moment norm of $\mu$ as
+$\mathcal{W}_{\theta}(\mu)=\bigg(\int_{\mathbb{R}^{d}}|x|^{\theta}\mu({\rm d}x)\bigg)^{1/\theta}.$
+Next, for two probability measures $\mu,\nu\in\mathcal{P}_{\theta}(\mathbb{R}^{d})$, the Wasserstein distance between $\mu$ and $\nu$ is defined as
+$\mathbb{W}_{\theta}(\mu,\nu)=\bigg(\inf_{\pi\in\Pi(\mu,\nu)}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}|x-y|^{\theta}\pi({\rm d}x,{\rm d}y)\bigg)^{1/\theta},$
+where $\Pi(\mu,\nu)$—the set of coupling measures between $\mu$ and $\nu$, consists of all probability measures on $\mathbb{R}^{d}\times\mathbb{R}^{d}$ whose marginal measures are $\mu$ and $\nu$, respectively.
+Precisely, a measure $\pi\in\Pi(\mu,\nu)$ satisfies $\pi(A\times\mathbb{R}^{d})=\mu(A)$ and $\pi(\mathbb{R}^{d}\times A)=\nu(A)$ for every Borel set $A\subseteq\mathbb{R}^{d}$.
+The space $\mathcal{P}_{2}(\mathbb{R}^{d})$ is a Polish space when equipped with the $L^{2}$-Wasserstein distance [27, Definition 6.1].
+Let $\{B_{t}^{H}\}_{t\in[0,T]}$ be a $d$-dimensional fractional Brownian motion with Hurst exponent $H\in(0,1)$. For Hurst exponents $1/2<H<1$, we define the kernel function
 (2.1) $\phi(s,t)=\phi_{H}(s,t)=H(2H-1)|s-t|^{2H-2},\ \ s,t\in\mathbb{R},$
-and recall that for $s,t>0$,
-$\int_{0}^{t}\int_{0}^{s}\phi(u,v){\rm d}u{\rm d}v=R_{H}(t,s).$
-Let $\mathcal{S}(\mathbb{R})$ be the Schwartz space of rapidly decreasing smooth functions on $R$, and if $b\in\mathcal{S}(\mathbb{R})$, denote
-$\parallel f\parallel_{H}^{2}:=\int_{\mathbb{R}}\int_{\mathbb{R}}f(s)f(t)\phi(s,t){\rm d}s{\rm d}t<\infty.$
-If we equip $\mathcal{S}(\mathbb{R})$ with the inner product
-$\langle f,g\rangle_{H}:=\int_{\mathbb{R}}\int_{\mathbb{R}}f(s)g(t)\phi(s,t){\rm d}s{\rm d}t,\ \ f,g\in\mathcal{S}(\mathbb{R}),$
-then the completion of $\mathcal{S}(\mathbb{R})$, denoted by $L_{\phi}^{2}(\mathbb{R})$,
+and recall that for $s,t>0$, this kernel satisfies the integral identity
+$\int_{0}^{t}\int_{0}^{s}\phi(u,v){\rm d}u{\rm d}v=R_{H}(t,s),$
+where $R_{H}(t,s)$ is defined as in (1.2).
+Let $\mathcal{S}(\mathbb{R})$ be the Schwartz space (space of rapidly decreasing smooth functions) on $R$.
+For any $f\in\mathcal{S}(\mathbb{R})$, we define its squared norm with respect to the kernel $\phi$ as
+$\|f\|_{H}^{2}:=\int_{\mathbb{R}}\int_{\mathbb{R}}f(s)f(t)\phi(s,t){\rm d}s{\rm d}t<\infty,$
+where $\phi$ is the kernel defined in (2.1). Next, we equip $\mathcal{S}(\mathbb{R})$ with an inner product induced by $\phi$ as
+$\langle f,g\rangle_{H}:=\int_{\mathbb{R}}\int_{\mathbb{R}}f(s)g(t)\phi(s,t){\rm d}s{\rm d}t,\ \ f,g\in\mathcal{S}(\mathbb{R}).$
+The completion of $\mathcal{S}(\mathbb{R})$ with respect to this inner product is denoted by $L_{\phi}^{2}(\mathbb{R})$.
+For the half line $\mathbb{R}_{+}:=[0,+\infty)$, we further define the space $L_{\phi}^{2}(\mathbb{R}_{+})$ as the set of measurable functions $g:\mathbb{R}_{+}\rightarrow\mathbb{R}$ such that their squared $\phi$-norm is finite:
 $L_{\phi}^{2}(\mathbb{R}_{+})=\bigg\{g|g:\mathbb{R}_{+}\to\mathbb{R},\ \ \ |g|_{\phi}^{2}:=\int_{0}^{\infty}\int_{0}^{\infty}g(s)g(t)\phi(s,t){\rm d}s{\rm d}t<\infty\bigg\}.$
+Here, $L_{\phi}^{2}(\mathbb{R}_{+})$ can be understood as the natural restriction of $L_{\phi}^{2}(\mathbb{R})$ to functions supported on $\mathbb{R}_{+}$, and it is also a Hilbert space under the inner product
+$\langle g_{1},g_{2}\rangle_{\phi}:=\int_{0}^{\infty}\int_{0}^{\infty}g_{1}(s)g_{2}(t)\phi(s,t){\rm d}s{\rm d}t.$
 Definition 2.1.
-[32, Definition 3.5.1] Let $g\in L_{\phi}^{2}(\mathbb{R})$. The $\phi$-derivative of a stochastic variable $F\in L^{p}$ in the direction of $\Phi_{g}$ is defined by
-$D_{\Phi_{g}}F(\omega)=\lim_{\delta\to 0}\frac{1}{\delta}\bigg{\{}F\bigg{(}\omega+\delta\int_{0}^{\cdot}(\Phi_{g})(u){\rm d}u\bigg{)}-F(\omega)\bigg{\}},$
-if the limit exists in $L^{p}$. Furthermore, if there is a process $(D_{s}^{\phi}F,s\geq 0)$ such that
-$D_{\Phi_{g}}F=\int_{0}^{\infty}D_{s}^{\phi}Fg_{s}{\rm d}s\ \ \ {\rm almost\ surely},$
-for all $g\in L_{\phi}^{2}(\mathbb{R})$, then $F$ is said to be $\phi$-differentiable.
-For SDEs driven by fractional Brownian motion
-(2.2) ${\rm d}X_{t}=G(t,X_{t}){\rm d}t+F(t,X_{t}){\rm d}B_{t}^{H}.$
-Let $C^{2,1}(\mathbb{R}\times\mathbb{R}_{+};\mathbb{R})$ be the family of all real-valued functions $V(x,t)$ defined on $\mathbb{R}\times\mathbb{R}_{+}$, which is twice continuously differentiable in $x$ and once in $t$.
-Referring to the Itô formula for fractional Brownian motion given by [32, 33],
+[28, Definition 3.5.1] Let $g\in L_{\phi}^{2}(\mathbb{R})$, and $\Phi_{g}$ denote a random variable induced by $g$, i.e., $\Phi_{g}=\int_{\mathbb{R}}g(s){\rm d}B_{s}^{H}$.
+For $F\in L^{p}$, the $\phi$-derivative of $F$ in the direction of $\Phi_{g}$ is defined as
+$D_{\Phi_{g}}F(\omega)=\lim_{\delta\to 0}\frac{1}{\delta}\bigg\{F\bigg(\omega+\delta\int_{0}^{\cdot}g(u){\rm d}u\bigg)-F(\omega)\bigg\},$
+provided the limit exists in $L^{p}$. Furthermore, if there is a process $(D_{s}^{\phi}F,s\geq 0)$ such that for all $g\in L_{\phi}^{2}(\mathbb{R})$
+$D_{\Phi_{g}}F=\int_{0}^{\infty}D_{s}^{\phi}F\cdot g_{s}{\rm d}s\ \ \ {\rm almost\ surely},$
+then $F$ is said to be $\phi$-differentiable.
+We denote $L_{\phi}(0,T)$ as the set of stochastic processes $F$ on $[0,T]$ that satisfy the following properties: $F\in L_{\phi}(0,T)$ if and only if $\mathbb{E}\|F\|_{H}^{2}<\infty$, $F$ is $\phi$-differentiable, the trace $D_{s}^{\phi}F_{t}$ exists for all $0\leq s,t\leq T$, $\mathbb{E}\int_{0}^{T}\int_{0}^{T}|D_{s}^{\phi}F_{t}|^{2}\,ds\,dt<\infty$, and for each sequence of partitions $(\pi_{n},n\in\mathbb{N})$ such that $\|\pi_{n}\|\to 0$ as $n\to\infty$, it holds that
+$\sum_{i,j=0}^{n-1}\mathbb{E}\left[\int_{t_{i}^{(n)}}^{t_{i+1}^{(n)}}\int_{t_{j}^{(n)}}^{t_{j+1}^{(n)}}|D_{s}^{\phi}F_{t_{i}^{(n)}}^{\pi}D_{t}^{\phi}F_{t_{j}^{(n)}}^{\pi}-D_{s}^{\phi}F_{t}D_{t}^{\phi}F_{s}|dsdt\right]\to 0,$
+and
+$\mathbb{E}\|F^{\pi}-F\|_{H}^{2}\to 0,n\to\infty,$
+where $\pi_{n}:0=t_{0}^{(n)}<t_{1}^{(n)}<\cdots<t_{n-1}^{(n)}<t_{n}^{(n)}=T.$
+Consider the $d$-dimensional SDEs driven by a $d$-dimensional fractional Brownian motion $B_{t}^{H}$ with Hurst exponent $H\in(1/2,1)$ as follows
+(2.2) ${\rm d}X_{t}=G(t,X_{t}){\rm d}t+F(t,X_{t}){\rm d}B_{t}^{H},$
+where $G:[0,T]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ and $F:[0,T]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d\times d}$.
+Let $C^{2,1}([0,T]\times\mathbb{R}^{d};\mathbb{R})$ denote the space of real-valued functions $V(t,x)$ that are twice continuously differentiable in $x\in\mathbb{R}^{d}$ and once continuously differentiable in $t\in[0,T]$.
+Based on the fractional Itô formula derived in [28, 29], the differential form of $V(t,X_{t})$ takes the form:
 ${\rm d}V(t,X_{t})=$ $\frac{\partial V(t,X_{t})}{\partial t}{\rm d}t+\frac{\partial V(t,X_{t})}{\partial X}G(t,X_{t}){\rm d}t+\frac{\partial V(t,X_{t})}{\partial X}F(t,X_{t}){\rm d}B_{t}^{H}$
 $+\frac{\partial^{2}V(t,X_{t})}{\partial X^{2}}F(t,X_{t})D_{t}^{\phi}X_{t}{\rm d}t,$
-or formally,
-${\rm d}V(t,X_{t})=$ $\frac{\partial V(t,X_{t})}{\partial t}{\rm d}t+\frac{\partial V(t,X_{t})}{\partial X}G_{t,X_{t}}{\rm d}t+\frac{\partial V(t,X_{t})}{\partial X}F(t,X_{t}){\rm d}B_{t}^{H}$
+where $\frac{\partial V}{\partial X}$ denotes the gradient, $\frac{\partial^{2}V}{\partial X^{2}}$ is the Hessian matrix. Formally, this can also be expressed by the kernel $\phi$ as
+${\rm d}V(t,X_{t})=$ $\frac{\partial V(t,X_{t})}{\partial t}{\rm d}t+\frac{\partial V(t,X_{t})}{\partial X}G({t,X_{t}}){\rm d}t+\frac{\partial V(t,X_{t})}{\partial X}F(t,X_{t}){\rm d}B_{t}^{H}$
 $+\frac{\partial^{2}V(t,X_{t})}{\partial X^{2}}F(t,X_{t})\int_{0}^{t}\phi(t,r)F(r,X_{r}){\rm d}r{\rm d}t.$
-In addition, [34], [35] also discussed the fractional Itô formula. Now, we are going to give the fractional Itô formula as follows.
-Remark 2.1.
-(Fractional Itô formula) Assume that the stochastic process $X_{t}$ follows (2.2) with $11/2<H<$ 1. We have
+Additional discussions on the fractional Itô formula can be found in [30, 31]. We summarize the result below.
+Lemma 2.1.
+(Fractional Itô formula) Let $X_{t}$ be the solution of (2.2) with $1/2<H<1$, and assume that $\mathbb{E}\bigg[\int_{0}^{T}\left\lvert F(s,X_{s})D_{s}^{\phi}\int_{0}^{s}F(u,X_{u}){\rm d}B_{u}^{H}\right\rvert{\rm d}s\bigg]<\infty$, $\frac{\partial V(s,X_{s})}{\partial X}F(s,X_{s})\in L_{\phi}(0,T)$ for $V\in C^{2,1}([0,T]\times\mathbb{R}^{d};\mathbb{R})$.
+Then we have
 $V(t,X_{t})$
-$=$ $V(X_{0})+\int_{0}^{t}\bigg{[}\frac{\partial V(s,X_{s})}{\partial t}+\frac{\partial V(s,X_{s})}{\partial X}G(s,X_{s})$
-$+H(2H-1)\frac{\partial^{2}V(s,X_{s})}{\partial X^{2}}\int_{0}^{s}(s-r)^{2H-2}F^{2}(r,X_{r}){\rm d}r\bigg{]}{\rm d}s+\int_{0}^{t}\frac{\partial V(s,X_{s})}{\partial X}F(s,X_{s}){\rm d}B_{s}^{H}$
-$=$ $V(X_{0})+\int_{0}^{t}\bigg{[}\frac{\partial V(s,X_{s})}{\partial t}+\frac{\partial V(s,X_{s})}{\partial X}G(s,X_{s})+H\frac{\partial^{2}V(s,X_{s})}{\partial X^{2}}s^{2H-1}F^{2}(s,X_{s})\bigg{]}{\rm d}s$
-$+\int_{0}^{t}\frac{\partial V(s,X_{s})}{\partial X}F(s,X_{s}){\rm d}B_{s}^{H}.$
-Since the kernel function $\phi$ is defined only for $11/2<H<$ 1, otherwise, $\phi$ has no meaning.
-We assume that $11/2<H<$ 1 throughout this paper.
-In this paper, we consider the following $d$-dimensional McKean-Vlasov NSDDEs driven by fractional Brownian motion as follows:
-(2.5) $\left\{\begin{array}[]{ll}{\rm d}(X_{t}-D(X_{t-\tau}))=b(X_{t},X_{t-\tau},\mathcal{L}_{X_{t}}){\rm d}t+\sigma(\mathcal{L}_{X_{t}}){\rm d}B_{t}^{H},\ \ t>0,\\ X_{t}=\xi=\{\xi(t):-\tau\leq t\leq 0\}\in L_{\mathcal{F}_{0}}^{p}([-\tau,0];\mathbb{R}^{d}),\ p\geq\theta\geq 1\end{array}\right.$
-where $b:\mathbb{R}^{d}\times\mathbb{R}^{d}\times\mathcal{P}_{q}(\mathbb{R}^{d})\to\mathbb{R}^{d}$, $\sigma:\mathcal{P}_{q}(\mathbb{R}^{d})\to\mathbb{R}^{d}\otimes\mathbb{R}^{d}$, $\mathcal{L}_{X_{t}}$ denotes the law of $X_{t}$.
-Throughout the paper, we shall denote $C$ by a generic positive constant, whose value may change from line to line.
-For any $x,y,\overline{x},\overline{y}$, we impose the following assumptions.
+$=$ $V(0,X_{0})+\int_{0}^{t}\bigg[\frac{\partial V(s,X_{s})}{\partial s}+\frac{\partial V(s,X_{s})}{\partial X}G(s,X_{s})$
+$+H(2H-1)\text{{\rm tr}}\left(\frac{\partial^{2}V(s,X_{s})}{\partial X^{2}}\int_{0}^{s}(s-r)^{2H-2}F(r,X_{r})F^{T}(r,X_{r}){\rm d}r\right)\bigg]{\rm d}s$
+$+\int_{0}^{t}\frac{\partial V(s,X_{s})}{\partial X}F(s,X_{s}){\rm d}B_{s}^{H},$
+where $F^{T}$ is the transpose of $F$, and $\text{{\rm tr}}(\cdot)$ denotes the matrix trace.
+Since the kernel function $\phi$ is only valid for $1/2<H<1$, we restrict our analysis to Hurst exponents satisfying $1/2<H<1$ throughout this paper.
+In this work, we focus on the following $d$-dimensional McKean-Vlasov NSDDEs driven by fractional Brownian motion as follows:
+(2.5) $\left\{\begin{array}[]{ll}{\rm d}(X_{t}-D(X_{t-\tau}))=b(X_{t},X_{t-\tau},\mathcal{L}_{X_{t}}){\rm d}t+\sigma(\mathcal{L}_{X_{t}}){\rm d}B_{t}^{H},\ \ t>0,\\ X_{t}=\xi=\{\xi(t):-\tau\leq t\leq 0\}\in L_{\mathcal{F}_{0}}^{{p}}([-\tau,0];\mathbb{R}^{d}),\ {p}\geq 1\end{array}\right$.
+where $b:\mathbb{R}^{d}\times\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}^{d}$, $\sigma:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}^{d\times d}$, $D:\mathbb{R}^{d}\to\mathbb{R}^{d}$, $\mathcal{L}_{X_{t}}$ denotes the law of $X_{t}$.
+Throughout the paper, we shall denote by $C$ a generic positive constant that may depend on $T,H,\xi,l,\lambda,K_{i}$ etc., but independent of $n,m,s,t,N,\Delta$.
+The value may change from line to line.
 Assumption 2.2.
@@
 Assumption 2.3.
-$D(0)=0$, and there exist positive constants $\lambda<1$ such that
-$|D(x)-D(\overline{x})|\leq\lambda|x-\overline{x}|.$
+$D(0)=0$, and there exists a constant $0<\lambda<1$ such that
+$|D(x)-D(\overline{x})|\leq\lambda|x-\overline{x}|,$
+for $x,\overline{x}\in\mathbb{R}^{d}$.
 Assumption 2.4.
-There exist positive constants $l\geq 1$ and $K_{2},K_{3}$ such that
-$\big{\langle}x-D(y)-\overline{x}+D(\overline{y}),b(x,y,\mu)-b(\overline{x},\overline{y},\nu)\big{\rangle}\leq K_{2}(|x-\overline{x}|^{2}+|y-\overline{y}|^{2}+(\mathbb{W}_{q}(\mu,\nu))^{2}),$
+There exist positive constants $l\geq 1$, $K_{2}$ and $K_{3}$ such that
+$\big\langle x-D(y)-\overline{x}+D(\overline{y}),b(x,y,\mu)-b(\overline{x},\overline{y},\nu)\big\rangle\leq K_{2}(|x-\overline{x}|^{2}+|y-\overline{y}|^{2}+\mathbb{W}_{2}^{2}(\mu,\nu)),$
 and
-$|b(x,y,\mu)-b(\overline{x},\overline{y},\nu)|\leq K_{3}[(1+|x|^{l}+|\overline{x}|^{l}+|y|^{l}+|\overline{y}|^{l})(|x-\overline{x}|+|y-\overline{y}|)+\mathbb{W}_{q}(\mu,\nu)],$
-for $x,y,\overline{x},\overline{y}\in\mathbb{R}^{d}$, $\mu,\nu\in\mathcal{P}_{q}(\mathbb{R}^{d})$.
+$|b(x,y,\mu)-b(\overline{x},\overline{y},\nu)|\leq K_{3}[(1+|x|^{l}+|\overline{x}|^{l}+|y|^{l}+|\overline{y}|^{l})(|x-\overline{x}|+|y-\overline{y}|)+\mathbb{W}_{2}(\mu,\nu)],$
+for $x,y,\overline{x},\overline{y}\in\mathbb{R}^{d}$, $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$.
 Assumption 2.5.
-There exist positive constants $K_{4},K_{5}$ such that
-$\parallel\sigma(\mu)-\sigma(\nu)\parallel\leq K_{4}\mathbb{W}_{q}(\mu,\nu),$
-and,
-$|b(0,0,\mu)|\vee\parallel\sigma(\mu)\parallel\leq K_{5}(1+\mathbb{W}_{q}(\mu,\delta_{0})),$
-for $\mu,\nu\in\mathcal{P}_{q}(\mathbb{R}^{d})$. Then, for initial experience distribution $\delta_{0}$
-$|b(0,0,\delta_{0})|\vee\parallel\sigma(\delta_{0})\parallel\leq K_{6}.$
+There exist a positive constant $K_{4}$ such that
+$\parallel\sigma(\mu)-\sigma(\nu)\parallel\leq K_{4}\mathbb{W}_{2}(\mu,\nu).$
+And for the initial experience distribution $\delta_{0}$, there is a positive constant $K_{5}$ such that
+$|b(0,0,\delta_{0})|\vee\parallel\sigma(\delta_{0})\parallel\leq K_{5}.$
 Remark 2.6.
-Let Assumption 2.4 hold, then, for any $x,y\in\mathbb{R}^{d}$, we have
-$|b(x,y,\mu)|\leq$ $|b(x,y,\mu)-b(0,0,\mu)|+|b(0,0,\mu)|$
-$\leq$ $K_{3}(1+|x|^{l}+|y|^{l})(|x|+|y|)+K_{5}(1+\mathbb{W}_{q}(\mu,\delta_{0}))$
-$\leq$ $C(1+|x|^{l+1}+|y|^{l+1}+\mathbb{W}_{q}(\mu,\delta_{0})),$
-where $C=K_{3}\vee K_{5}$. Moreover, by Young’s inequality, we get
-$(x-D(y))b(x,y,\mu)=$ $(x-D(y)-0+D(0))(b(x,y,\mu)-b(0,0,\delta_{0})+b(0,0,\delta_{0}))$
-$\leq$ $K_{2}(|x|^{2}+|y|^{2}+\mathbb{W}_{q}^{2}(\mu,\delta_{0}))+(|x|+D(y))|b(0,0,\delta_{0})|$
-$\leq$ $C(1+|x|^{2}+|y|^{2}+\mathbb{W}_{q}^{2}(\mu,\delta_{0})),$
-where $C=K_{2}\vee(K_{2}+1)\vee(K_{2}+\lambda)\vee(K_{2}+K_{5}^{2})$.
+Let Assumptions 2.4-2.5 hold, then, for any $x,y\in\mathbb{R}^{d}$, we have
+$|b(x,y,\mu)|\leq$ $|b(x,y,\mu)-b(0,0,\delta_{0})|+|b(0,0,\delta_{0})|$
+$\leq$ $K_{3}[(1+|x|^{l}+|y|^{l})(|x|+|y|)+\mathbb{W}_{2}(\mu,\delta_{0})]+K_{5}$
+$\leq$ $C(1+|x|^{l+1}+|y|^{l+1}+\mathbb{W}_{2}(\mu,\delta_{0})),$
+where $C=2(K_{3}+K_{5})$. Moreover,
+$\|\sigma(\mu)\|\leq$ $\|\sigma(\mu)-\sigma(\delta_{0})\|+\|\sigma(\delta_{0})\|\leq K_{4}\mathbb{W}_{2}(\mu,\delta_{0})+K_{5}\leq C(1+\mathbb{W}_{2}(\mu,\delta_{0})),$
+where $C=K_{4}\vee K_{5}$. Further, let Assumptions 2.3-2.5 hold, then by the Young’s inequality, we get
+$\langle x-D(y),b(x,y,\mu)\rangle=$ $\langle x-D(y)-0+D(0),b(x,y,\mu)-b(0,0,\delta_{0})+b(0,0,\delta_{0})\rangle$
+$\leq$ $K_{2}(|x|^{2}+|y|^{2}+\mathbb{W}_{2}^{2}(\mu,\delta_{0}))+(|x|+|D(y)|)|b(0,0,\delta_{0})|$
+$\leq$ $C(1+|x|^{2}+|y|^{2}+\mathbb{W}_{2}^{2}(\mu,\delta_{0})),$
+where $C=(K_{2}+1/2)\vee K_{5}^{2}$.
+For later use, we state the following lemmas without proof.
 Lemma 2.7.
-[31, Lemma 4.1] Let $p>1$, $\varepsilon>0$ and $x,y\in\mathbb{R}$. Then
+[32, Lemma 4.1] Let $p>1$, $\varepsilon>0$ and $x,y\in\mathbb{R}$. Then
 $|x+y|^{p}\leq\bigg(1+\varepsilon^{\frac{1}{p-1}}\bigg)^{p-1}\bigg(|x|^{p}+\frac{|y|^{p}}{\varepsilon}\bigg).$
-Especially for $\varepsilon=((1-a)/a)^{p-1}$, where $p\geq 2$ and $10<a<$ 1, we get
+Especially for $\varepsilon=[(1-a)/a]^{p-1}$ where $0<a<1$, we get
 $|x+y|^{p}\leq\frac{1}{a^{p-1}}|x|^{p}+\frac{1}{(1-a)^{p-1}}|y|^{p}.$
+Lemma 2.8.
+[33, Lemma 7.1.2] Let $\beta>0$, $\gamma>0$ with $\beta+\gamma>1$. Suppose $a\geq 0$, $b\geq 0$, $u$ is nonnegative and $t^{\gamma-1}u(t)$ is locally integrable on $0\leq t\leq T$. If
+$u(t)\leq a+b\int_{0}^{t}(t-s)^{\beta-1}s^{\gamma-1}u(s){\rm d}s,\ a.e.$
+then
+$u(t)\leq aE_{\beta,\gamma}((b\Gamma(\beta))^{1/v}t),$
+where $v=\beta+\gamma-1>0$, and $E_{\beta,\gamma}(s)$ denotes the generalized Mittag-Leffler function defined by $E_{\beta,\gamma}(s)=\sum_{m=0}^{\infty}c_{m}s^{mv}$ with $c_{0}=1$, $\frac{c_{m+1}}{c_{m}}=\frac{\Gamma(mv+\gamma)}{\Gamma(mv+\gamma+\beta)}$ for all $m\geq 0$, here $\Gamma(\cdot)$ is the Gamma function.
+Remark 2.9.
+Considering the special situation of Lemma 2.8 with $\beta=2H(1/2<H<1)$, $\gamma=1$, $b=1$, it follows immediately that $u(t)$ is bounded, i.e., there exists a positive constant $C$ such that
+$u(t)\leq C.$
+Since for these parameters, we have $\beta+\gamma=2H+1>1$, so $v=2H>0$.
+The Mittag-Leffler function in Lemma 2.8 reduces to $E_{2H,1}(s)=\sum_{m=0}^{\infty}c_{m}s^{2Hm}$ where $s=[\Gamma(2H)]^{1/2H}t$.
+This series converges for all $t\in\mathbb{R}$, as shown by analyzing the asymptotic behavior of its coefficients.
+Using the Stirling’s approximation for the Gamma function, we derive
+$\frac{c_{m+1}}{c_{m}}=\frac{\Gamma(2Hm+1)}{\Gamma(2Hm+1+2H)}\sim\frac{(2Hm)^{2Hm+1/2}e^{-2Hm}}{(2Hm+2H)^{2Hm+2H+1/2}e^{-(2Hm+2H)}}.$
+Simplifying further, this ratio satisfies
+$\frac{c_{m+1}}{c_{m}}\sim(2Hm)^{-2H}e^{2H}.$
+Since for $2H>0$ we have
+$\lim_{m\to\infty}(2Hm)^{-2H}e^{2H}=0,$
+the ratio test guarantees that the series $\sum_{m=0}^{\infty}c_{m}s^{2Hm}$ has an infinite radius of convergence.
+Thus, $E_{2H,1}(\cdot)$ is bounded for all $t\in\mathbb{R}$, implying $u(t)\leq a\cdot E_{2H,1}(\cdot)\leq C.$
 3 Existence and Uniqueness
-Since we are studying equations with super-linearly growing coefficients, it is necessary to consider the Carathéodory approximation.
-Now we define the Carathéodory approximation of (2.5) as follows.
-For every integer $n\geq 2/\tau$, define $X_{t}^{n}$ on $[-\tau,T]$ by
-$X_{t}^{n}=\xi(t),\ \ t\in[-\tau,0],$
+Theorem 3.1.
+Let Assumptions 2.2-2.5 hold. Then there exists a unique solution $\{X_{t}\}_{t\in[-\tau,T]}$ to (2.5). Moreover, for any $\overline{p}\geq 2$, we have
+(3.1) $\mathbb{E}\bigg(\sup_{0\leq t\leq T}|X_{t}|^{\overline{p}}\bigg)\leq C,$
+where $C$ is a positive constant depending on $T,H,\overline{p},\lambda,K_{i}$ and the initial data $\xi$.
+Proof.
+We note that the application of the fractional Itô formula in the following argument is justified by the Lipschitz conditions in Assumptions 2.3-2.5, which ensure the necessary regularity of the Picard iteration sequence $X_{t}^{(k)}$.
+Standard theory of Malliavin calculus for SDEs (e.g.
+[28]) guarantees that these iterates are $\phi$-differentiable and satisfy the integrability conditions required by Lemma 2.1.
+We refer to the proof procedure in [24] and divide the proof into 3 steps.
+Step 1.
+Moment Boundness For $k\geq 0$ and $t\in[-\tau,0]$, define $X^{(k)}_{t}=\xi(t)$.
+For $k=0$ and $t\in[0,T]$, define $X^{(0)}_{t}\equiv\xi(0)$.
+For $k\geq 1$ and $t\in[0,T]$, $X_{t}^{(k)}$ is solved by
+${\rm d}(X_{t}^{(k)}-D(X^{(k)}_{t-\tau}))=b(X_{t}^{(k)},X_{t-\tau}^{(k)},\mathcal{L}_{X_{t}^{(k-1)}}){\rm d}t+\sigma(\mathcal{L}_{X_{t}^{(k-1)}}){\rm d}B_{t}^{H},$
+where $\mathcal{L}_{X_{t}^{(0)}}=\mathcal{L}_{\xi_{0}}$. For $i=1$, by Assumption 2.3 and Lemma 2.7, we get
+$|X_{t}^{(1)}|^{\overline{p}}=$ $|D(X_{t-\tau}^{(1)})+X_{t}^{(1)}-D(X_{t-\tau}^{(1)})|^{\overline{p}}$
+$\leq$ $\frac{1}{\lambda^{{\overline{p}}-1}}|D(X_{t-\tau}^{(1)})|^{\overline{p}}+\frac{1}{(1-\lambda)^{{\overline{p}}-1}}|X_{t}^{(1)}-D(X_{t-\tau}^{(1)})|^{\overline{p}}$
+$\leq$ $\lambda|X_{t-\tau}^{(1)}|^{\overline{p}}+\frac{1}{(1-\lambda)^{{\overline{p}}-1}}|X_{t}^{(1)}-D(X_{t-\tau}^{(1)})|^{\overline{p}}.$
+Then, we conclude that
+(3.2) $\sup_{0\leq s\leq t}|X_{s}^{(1)}|^{\overline{p}}\leq\frac{\lambda}{1-\lambda}\|\xi\|^{\overline{p}}+\frac{1}{(1-\lambda)^{\overline{p}}}\sup_{0\leq s\leq t}|X_{s}^{(1)}-D(X_{s-\tau}^{(1)})|^{\overline{p}}.$
+Using the fractional Itô formula in Lemma 2.1, it follows that
+$|X_{t}^{(1)}-D(X_{t-\tau}^{(1)})|^{\overline{p}}$
+$\leq$ $|\xi(0)-D(\xi(-\tau))|^{\overline{p}}$
+$+{\overline{p}}\int_{0}^{t}|X_{s}^{(1)}-D(X_{s-\tau}^{(1)})|^{{\overline{p}}-2}\langle X_{s}^{(1)}-D(X_{s-\tau}^{(1)}),b(X_{s}^{(1)},X_{s-\tau}^{(1)},\mathcal{L}_{X_{s}^{(0)}})\rangle{\rm d}s$
+$+\int_{0}^{t}{\overline{p}}({\overline{p}}-1)H(2H-1)|X_{s}^{(1)}-D(X_{s-\tau}^{(1)})|^{{\overline{p}}-2}\int_{0}^{s}(s-r)^{2H-2}\|\sigma(\mathcal{L}_{X_{r}^{(0)}})\|^{2}{\rm d}r{\rm d}s$
+$+{\overline{p}}\int_{0}^{t}|X_{s}^{(1)}-D(X_{s-\tau}^{(1)})|^{{\overline{p}}-1}\|\sigma(\mathcal{L}_{X_{s}^{(0)}})\|{\rm d}B_{s}^{H}.$
+$=$ $|\xi(0)-D(\xi(-\tau))|^{\overline{p}}+Q_{1}^{(1)}(t)+Q_{2}^{(1)}(t)+Q_{3}^{(1)}(t).$
+For any positive integer $N$, define the stopping time
+$\rho_{N}^{(1)}=T\wedge\inf\{t\in[0,T]:|X_{t}^{(1)}|\geq N\}.$
+It yields that $\rho_{N}^{(1)}\to T$ a.s. as $N\to\infty$. Applying Assumptions 2.3 and 2.4, the Young’s inequality, we see
+$\mathbb{E}\bigg(\sup_{0\leq s\leq t\wedge\rho_{N}^{(1)}}Q_{1}^{(1)}(s)\bigg)$
+$=$ ${\overline{p}}\mathbb{E}\int_{0}^{t\wedge\rho_{N}^{(1)}}|X_{s}^{(1)}-D(X_{s-\tau}^{(1)})|^{{\overline{p}}-2}\langle X_{s}^{(1)}-D(X_{s-\tau}^{(1)}),b(X_{s}^{(1)},X_{s-\tau}^{(1)},\mathcal{L}_{X_{s}^{(0)}})\rangle{\rm d}s$
+$\leq$ $C\mathbb{E}\int_{0}^{t\wedge\rho_{N}^{(1)}}|X_{s}^{(1)}-D(X_{s-\tau}^{(1)})|^{\overline{p}}{\rm d}s+C\mathbb{E}\int_{0}^{t\wedge\rho_{N}^{(1)}}\bigg[1+|X_{s}^{(1)}|^{2}+|X_{s-\tau}^{(1)}|^{2}+\mathbb{W}_{2}^{2}(\mathcal{L}_{X_{s}^{(0)}},\delta_{0})\bigg]^{{\overline{p}}/2}{\rm d}s$
+$\leq$ $C+C\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r\wedge\rho_{N}^{(1)}}|X_{s}^{(1)}|^{\overline{p}}\bigg){\rm d}r.$
+Since Assumption 2.5 implies $\|\sigma(\mathcal{L}_{X_{r}^{(0)}})\|\leq C\big(1+\mathbb{W}_{2}(\mathcal{L}_{X_{r}^{(0)}},\delta_{0})\big)\leq C$, this together with the Young’s inequality, the Jensen inequality, and the Fubini theorem, it results in
+$\mathbb{E}\bigg(\sup_{0\leq s\leq t\wedge\rho_{N}^{(1)}}Q_{2}^{(1)}(s)\bigg)$
+$=$ $\mathbb{E}\int_{0}^{t\wedge\rho_{N}^{(1)}}{\overline{p}}({\overline{p}}-1)H(2H-1)|X_{s}^{(1)}-D(X_{s-\tau}^{(1)})|^{{\overline{p}}-2}\int_{0}^{s}(s-r)^{2H-2}\|\sigma(\mathcal{L}_{X_{r}^{(0)}})\|^{2}{\rm d}r{\rm d}s$
+$\leq$ $C\mathbb{E}\int_{0}^{t\wedge\rho_{N}^{(1)}}\int_{0}^{s}(s-r)^{2H-2}|X_{s}^{(1)}-D(X_{s-\tau}^{(1)})|^{{\overline{p}}-2}(1+\mathbb{W}_{2}(\mathcal{L}_{X_{r}^{(0)}},\delta_{0}))^{2}{\rm d}r{\rm d}s$
+$\leq$ $C\mathbb{E}\int_{0}^{t\wedge\rho_{N}^{(1)}}|X_{s}^{(1)}-D(X_{s-\tau}^{(1)})|^{\overline{p}}\bigg(\int_{0}^{s}(s-r)^{2H-2}{\rm d}r\bigg){\rm d}s+C\mathbb{E}\int_{0}^{t\wedge\rho_{N}^{(1)}}\int_{0}^{s}(s-r)^{2H-2}{\rm d}r{\rm d}s$
+$=$ $C\mathbb{E}\int_{0}^{t\wedge\rho_{N}^{(1)}}|X_{s}^{(1)}-D(X_{s-\tau}^{(1)})|^{\overline{p}}\bigg(\int_{0}^{s}(s-r)^{2H-2}{\rm d}r\bigg){\rm d}s+C\mathbb{E}\int_{0}^{t\wedge\rho_{N}^{(1)}}\int_{r}^{t}(s-r)^{2H-2}{\rm d}s{\rm d}r$
+$=$ $C\mathbb{E}\int_{0}^{t\wedge\rho_{N}^{(1)}}\frac{s^{2H-1}}{2H-1}|X_{s}^{(1)}-D(X_{s-\tau}^{(1)})|^{\overline{p}}{\rm d}s+C\mathbb{E}\int_{0}^{t\wedge\rho_{N}^{(1)}}\frac{(t-r)^{2H-1}}{2H-1}{\rm d}r.$
+Combining with Assumption 2.3 and the fact that $1/2<H<1$, we obtain
+$\mathbb{E}\bigg(\sup_{0\leq s\leq t\wedge\rho_{N}^{(1)}}Q_{2}^{(1)}(s)\bigg)\leq$ $C\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r\wedge\rho_{N}^{(1)}}|X_{s}^{(1)}|^{\overline{p}}\bigg){\rm d}r+C\mathbb{E}\int_{0}^{t\wedge\rho_{N}^{(1)}}(t-r)^{2H-1}{\rm d}r.$
+For $Q_{3}^{(1)}(s)$, we use the fractional Burkholder-Davis-Gundy (BDG) inequality. For $H\in(1/2,1)$ and $p\geq 1$, there exists a constant $C_{H,p}>0$ such that
+(3.3) $\mathbb{E}\left[\sup_{0\leq s\leq t}\left|\int_{0}^{s}u_{r}{\rm d}B_{r}^{H}\right|^{p}\right]\leq C_{H,p},\mathbb{E}\left[\left(\int_{0}^{t}|u_{r}|^{1/H}{\rm d}r\right)^{pH}\right].$
+Applying this with $p=1$ and $u_{r}=|X_{r}^{(1)}-D(X_{r-\tau}^{(1)})|^{{\overline{p}}-1}\|\sigma(\mathcal{L}_{{X_{r}^{(0)}}})\|$, and noticing the fact that Assumption 2.5 yields $\|\sigma(\mathcal{L}_{X_{r}^{(0)}})\|\leq C\big(1+\mathbb{W}_{2}(\mathcal{L}_{X_{r}^{(0)}},\delta_{0})\big)\leq C$, we get
+$\mathbb{E}\left(\sup_{0\leq s\leq t\wedge\rho_{N}^{(1)}}\left|Q_{3}^{(1)}(s)\right|\right)=\overline{p}\mathbb{E}\bigg(\sup_{0\leq s\leq t\wedge\rho_{N}^{(1)}}\left|\int_{0}^{s}u_{r}{\rm d}B_{r}^{H}\right|\bigg)\leq C\mathbb{E}\left[\left(\int_{0}^{t\wedge\rho_{N}^{(1)}}|u_{r}|^{1/H}{\rm d}r\right)^{H}\right]$
+$\leq$ $C\mathbb{E}\left[\left(\int_{0}^{t\wedge\rho_{N}^{(1)}}|X_{r}^{(1)}-D(X_{r-\tau}^{(1)})|^{({\overline{p}}-1)/H}{\rm d}r\right)^{H}\right].$
+For any $0<\epsilon<1$, the Young’s inequality with exponents $\frac{\overline{p}}{\overline{p}-1}$ and its conjugate gives
+$|X_{r}^{(1)}-D\big(X_{r-\tau}^{(1)}\big)|^{\overline{p}-1}\leq\epsilon|X_{r}^{(1)}-D\big(X_{r-\tau}^{(1)}\big)|^{\overline{p}}+C_{\epsilon},$
+where $C_{\epsilon}$ depends on $\epsilon,\overline{p},H$. Integrating over $0,t\wedge\rho_{N}^{(1)}]$, then
+$\int_{0}^{t\wedge\rho_{N}^{(1)}}|X_{r}^{(1)}-D\big(X_{r-\tau}^{(1)}\big)|^{(\overline{p}-1)/H}\mathrm{d}r\leq 2^{1/H-1}\epsilon^{1/H}\int_{0}^{t\wedge\rho_{N}^{(1)}}|X_{r}^{(1)}-D\big(X_{r-\tau}^{(1)}\big)|^{\overline{p}/H}\mathrm{d}r+2^{1/H-1}C_{\epsilon}^{1/H}t.$
+Since $H\in(1/2,1)$, we use the elementary inequality to get
+$\left(\int_{0}^{t\wedge\rho_{N}^{(1)}}|X_{r}^{(1)}-D\big(X_{r-\tau}^{(1)}\big)|^{(\overline{p}-1)/H}\mathrm{d}r\right)^{H}\leq$ $2^{1-H}\epsilon\left(\int_{0}^{t\wedge\rho_{N}^{(1)}}|X_{r}^{(1)}-D\big(X_{r-\tau}^{(1)}\big)|^{\overline{p}/H}\mathrm{d}r\right)^{H}+2^{1-H}C_{\epsilon}t^{H}$
+$\leq$ $2^{1-H}\epsilon t^{H}\cdot\sup_{0\leq r\leq t\wedge\rho_{N}^{(1)}}|X_{r}^{(1)}-D\big(X_{r-\tau}^{(1)}\big)|^{\overline{p}}+2^{1-H}C_{\epsilon}t^{H}.$
+Taking expectations and using Assumption 2.3
+$\mathbb{E}\left(\sup_{0\leq r\leq t\wedge\rho_{N}^{(1)}}|X_{r}^{(1)}-D\big(X_{r-\tau}^{(1)}\big)|^{\overline{p}}\right)\leq C\left(1+\mathbb{E}\left(\sup_{0\leq r\leq t\wedge\rho_{N}^{(1)}}\left|X_{r}^{(1)}\right|^{\overline{p}}\right)\right).$
+Combining all estimations, we arrive at
+(3.4) $\mathbb{E}\left(\sup_{0\leq s\leq t\wedge\rho_{N}^{(1)}}\left|Q_{3}^{(1)}(s)\right|\right)\leq C\epsilon\mathbb{E}\left(\sup_{0\leq s\leq t\wedge\rho_{N}^{(1)}}\left|X_{s}^{(1)}\right|^{\overline{p}}\right)+C_{\epsilon}.$
+Choosing $\epsilon$ sufficiently small so that $C\epsilon<\frac{1}{2}$, and combining with the estimates for $Q_{1}^{(1)}(t)$ and $Q_{2}^{(1)}(t)$, we get
+$\mathbb{E}\bigg(\sup_{0\leq s\leq t\wedge\rho_{N}^{(1)}}|X_{s}^{(1)}-D(X_{s-\tau}^{(1)})|^{\overline{p}}\bigg)\leq$ $C+C\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r\wedge\rho_{N}^{(1)}}|X_{s}^{(1)}|^{\overline{p}}\bigg){\rm d}r.$
+By the Gronwall inequality together with (3.2), we derive
+$\mathbb{E}\bigg(\sup_{0\leq s\leq t\wedge\rho_{N}^{(1)}}|X_{s}^{(1)}|^{\overline{p}}\bigg)\leq$ $C.$
+The Fatou lemma then leads to
+$\mathbb{E}\bigg(\sup_{0\leq s\leq t}|X_{s}^{(1)}|^{\overline{p}}\bigg)\leq C.$
+For $i=2$, we estimate the corresponding $Q_{1}^{(2)}(s),Q_{2}^{(2)}(s)$ and $Q_{3}^{(2)}(s)$ as follows
+$\mathbb{E}\bigg(\sup_{0\leq s\leq t\wedge\rho_{N}^{(2)}}Q_{1}^{(2)}(s)\bigg)\leq$ $C\bigg[\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r\wedge\rho_{N}^{(2)}}|X_{s}^{(2)}|^{\overline{p}}\bigg){\rm d}r+\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r\wedge\rho_{N}^{(2)}}|X_{s}^{(1)}|^{\overline{p}}\bigg){\rm d}r+1\bigg],$
+$\mathbb{E}\bigg(\sup_{0\leq s\leq t\wedge\rho_{N}^{(2)}}Q_{2}^{(2)}(s)\bigg)\leq$ $C\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r\wedge\rho_{N}^{(2)}}|X_{s}^{(2)}|^{\overline{p}}\bigg){\rm d}r+C\int_{0}^{t}(t-r)^{2H-1}\mathbb{E}\bigg(\sup_{0\leq s\leq r\wedge\rho_{N}^{(2)}}|X_{s}^{(1)}|^{\overline{p}}\bigg){\rm d}r,$
 and
-(3.1) $X_{t}^{n}-D(X_{t-\tau}^{n})=\xi(0)-D(\xi(-\tau))+\int_{0}^{t}b\big{(}X_{s-1/n}^{n},X_{s-\tau}^{n},\mathcal{L}_{X_{s-1/n}^{n}}\big{)}{\rm d}s+\int_{0}^{t}\sigma\big{(}\mathcal{L}_{X_{s-1/n}^{n}}\big{)}{\rm d}B_{s}^{H},$
-for $t\in[0,T]$.
-It is important to note that each $X_{t}^{n}$ can be determined explicitly by the stepwise iterated Itô integrals over the intervals $[0,1/n]$, $(1/n,2/n]$ etc.
-To get the final results, we give the following lemmas.
-Lemma 3.1.
-Let Assumptions 2.2-2.5 hold. Then for $\overline{p}>1/H$ and $n\geq 2/\tau$, we get
-$\mathbb{E}\bigg{(}\sup_{0\leq t\leq T}|X_{t}^{n}|^{\overline{p}}\bigg{)}\leq C.$
-Proof.
-By Assumption 2.4 and Lemma 2.7, we get
-$|X_{t}^{n}|^{\overline{p}}=$ $|D(X_{t-\tau}^{n})+X_{t}^{n}-D(X_{t-\tau}^{n})|^{\overline{p}}$
-$\leq$ $\frac{1}{\lambda^{p-1}}|D(X_{t-\tau}^{n})|^{\overline{p}}+\frac{1}{(1-\lambda)^{{\overline{p}}-1}}|X_{t}^{n}-D(X_{t-\tau}^{n})|^{\overline{p}}$
-$\leq$ $\lambda|X_{t-\tau}^{n}|^{\overline{p}}+\frac{1}{(1-\lambda)^{p-1}}|X_{t}^{n}-D(X_{t-\tau}^{n})|^{\overline{p}},$
-for any $t\in[0,T]$. Then
-$\sup_{0\leq r\leq t}|X_{r}^{n}|^{\overline{p}}\leq$ $\lambda\parallel\xi\parallel^{p}+\lambda\sup_{0\leq r\leq t}|X_{r}^{n}|^{\overline{p}}+\frac{1}{(1-\lambda)^{{\overline{p}}-1}}\sup_{0\leq r\leq t}|X_{r}^{n}-D(X_{r-\tau}^{n})|^{\overline{p}},$
-we may compute that
-(3.2) $\sup_{0\leq r\leq t}|X_{r}^{n}|^{\overline{p}}\leq\frac{\lambda}{1-\lambda}\parallel\xi\parallel^{\overline{p}}+\frac{1}{(1-\lambda)^{\overline{p}}}\sup_{0\leq r\leq t}|X_{r}^{n}-D(X_{r-\tau}^{n})|^{\overline{p}}$
-By Assumption 2.3, we get
-$\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|X_{s}^{n}|^{\overline{p}}\bigg{)}$
-$\leq$ $3^{\overline{p}-1}\mathbb{E}|\xi(0)-D(\xi(-\tau))|^{\overline{p}}+3^{\overline{p}-1}\mathbb{E}\bigg{(}\sup_{s\in[0,t]}\bigg{\lvert}\int_{0}^{s}b\left(X_{r-1/n}^{n},X_{r-\tau}^{n},\mathcal{L}_{X_{r-1/n}^{n}}\right){\rm d}r\bigg{\rvert}^{\overline{p}}\bigg{)}$
-$+3^{\overline{p}-1}\mathbb{E}\bigg{(}\sup_{s\in[0,t]}\bigg{\lvert}\int_{0}^{s}\sigma\left(\mathcal{L}_{X_{r-1/n}^{n}}\right){\rm d}B_{r}^{H}\bigg{\rvert}^{\overline{p}}\bigg{)}$
-$\leq$ $C+C\mathbb{E}\int_{0}^{t}\left[1+|X_{r-1/n}^{n}|^{(l+1)\overline{p}}+|X_{r-\tau}^{n}|^{(l+1){\overline{p}}}+\left(\mathbb{W}_{\theta}(\mathcal{L}_{X_{r-1/n}^{n}},\delta_{0})\right)^{\overline{p}}\right]{\rm d}r$
-$+C\mathbb{E}\bigg{(}\sup_{s\in[0,t]}\bigg{\lvert}\int_{0}^{s}\sigma(\mathcal{L}_{X_{r-1/n}^{n}}){\rm d}B_{r}^{H}\bigg{\rvert}^{\overline{p}}\bigg{)}.$
-Using the techniques of [12, Theorem 3.1].
-Denote $\int_{s}^{t}(t-\beta)^{-\beta}(r-s)^{\beta-1}{\rm d}r:=C(\beta)$.
-Since for $\overline{p}H>1$, we can take some $\beta$ such that $1-H<\beta<1-1/{H}$, then by the stochastic Fubini theorem, the Hölder inequality, and combining with Lemma 2.4, we get
-(3.3) $\mathbb{E}\bigg{(}\sup_{s\in[0,t]}\bigg{\lvert}\int_{0}^{s}\sigma(\mathcal{L}_{X_{r-1/n}^{n}}){\rm d}B_{r}^{H}\bigg{\rvert}^{\overline{p}}\bigg{)}$
-$\leq$ $\frac{C(\beta)^{-{\overline{p}}}}{({\overline{p}}-1-\beta{\overline{p}})^{{\overline{p}}-1}}t^{{\overline{p}}-1-\beta{\overline{p}}}\int_{0}^{t}\mathbb{E}\bigg{|}\int_{0}^{s}(s-r)^{\beta-1}\sigma(\mathcal{L}_{X_{r-1/n}^{n}}){\rm d}B_{r}^{H}\bigg{|}^{{\overline{p}}}{\rm d}s$
-$\leq$ $Ct^{{\overline{p}}-1-\beta{\overline{p}}}\int_{0}^{t}\left(\int_{0}^{s}(s-r)^{\frac{\beta-1}{H}}\parallel\sigma(\mathcal{L}_{X_{r-1/n}^{n}})\parallel^{\frac{1}{H}}{\rm d}r\right)^{\overline{p}H}{\rm d}s$
-$\leq$ $Ct^{\overline{p}H-1}\int_{0}^{t}\parallel\sigma(\mathcal{L}_{X_{r-1/n}^{n}})\parallel^{\overline{p}}{\rm d}r$
-$\leq$ $Ct^{\overline{p}H-1}L^{\overline{p}}\int_{0}^{t}\left[1+\mathbb{W}_{\theta}(\mathcal{L}_{X_{r-1/n}^{n}},\delta_{0})\right]^{\overline{p}}{\rm d}r.$
-Referring to the proof in [6, Proposition 3.4], for the $q$-Wasserstein metric, we get
-$\left(\mathbb{W}_{q}(\mathcal{L}_{X_{r-1/n}^{n}},\delta_{0})\right)^{\overline{p}}=\left(\mathcal{W}_{q}(\mathcal{L}_{X_{r-1/n}^{n}})\right)^{\overline{p}}\leq\mathbb{E}|X_{r-1/n}^{n}|^{\overline{p}}.$
-To sum up, we obtain
-$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}|X_{t}^{n}-D(X_{t-\tau}^{n})|^{\overline{p}}\bigg{)}\leq$ $C+C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{0\leq u\leq s}|X_{u-\tau}^{n}|^{(l+1)\overline{p}}\bigg{)}{\rm d}s,$
-together with (3.2), we get
-$\mathbb{E}\bigg{(}\sup_{0\leq r\leq t}|X_{r}^{n}|^{\overline{p}}\bigg{)}\leq$ $C+C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{0\leq u\leq s}|X_{u-\tau}^{n}|^{(l+1)\overline{p}}\bigg{)}{\rm d}s.$
-We set up a finite sequence $\{u_{1},u_{2},\cdots,u_{[T/\tau]+1}\}$ where $u_{i}=([T/\tau]+2-i)\overline{p}(l+1)^{[T/\tau]+1-i}$, and $[T/\tau]$ denotes the integer part of $T/\tau$.
-It is easy to see that $u_{i}\geq\overline{p}$, $u_{[T/\tau]+1}=\overline{p}$ and $(l+1)u_{i+1}<u_{i}$ for $i=1,2,\cdots,[T/\tau]$.
-For $s\in[0,\tau]$, we immediately get
-$\mathbb{E}\bigg{(}\sup_{0\leq r\leq t}|X_{r}^{n}|^{\overline{p}}\bigg{)}\leq C,$
-since $\xi\in L_{\mathcal{F}_{0}}^{\overline{p}}([-\tau,0];\mathbb{R}^{d})$. For $s\in[0,2\tau]$, by the Hölder inequality, we get
-$\mathbb{E}\bigg{(}\underset{s\in[0,2\tau]}{\sup}|X_{s}^{n}|^{u_{2}}\bigg{)}\leq C+C\int_{0}^{2\tau}\left[\mathbb{E}\bigg{(}\sup_{s\in[0,r]}|X_{s-\tau}^{n}|^{u_{1}}\bigg{)}\right]^{\frac{(l+1)u_{2}}{u_{1}}}{\rm d}r\leq C.$
-By induction, we can derive the assertion. ∎
-Lemma 3.2.
-Let Assumptions 2.2-2.5 hold. Then for $0\leq s\leq t\leq T$ with $t-s<1$, and $p>1/H$, $(l+1)p\leq\overline{p}$, $n\geq 2/\tau$, we have
-$\mathbb{E}|X_{t}^{n}-X_{s}^{n}|^{p}\leq C[(t-s)^{p}+(t-s)^{pH}],$
-where $C$ is a positive constant depending on $p,H,\xi,T,L$ but independent of $s$ and $t$.
-Proof.
-By (3.1), we may compute
-$X_{t}^{n}-X_{s}^{n}=$ $D(X_{t-\tau}^{n})-D(X_{s-\tau}^{n})+\int_{s}^{t}b\left(X^{n}_{r-1/n},X^{n}_{r-\tau},\mathcal{L}_{X_{r-1/n}^{n}}\right){\rm d}r+\int_{s}^{t}\sigma(\mathcal{L}_{X_{r-1/n}^{n}}){\rm d}B_{r}^{H}.$
-By the elementary inequality, the Hölder inequality, Assumptions 2.3, 2.4, 2.5 and using same techniques as (3.3), we get
-$\mathbb{E}|X_{t}^{n}-X_{s}^{n}|^{p}$
-$\leq$ $3^{p-1}\mathbb{E}|D(X_{t-\tau}^{n})-D(X_{s-\tau}^{n})|^{p}+3^{p-1}\mathbb{E}\bigg{\lvert}\int_{s}^{t}b(X_{r-1/n}^{n},X_{r-\tau}^{n},\mathcal{L}_{X_{r-1/n}^{n}}){\rm d}r\bigg{\rvert}^{p}$
-$+3^{p-1}\mathbb{E}\bigg{\lvert}\int_{s}^{t}\sigma(\mathcal{L}_{X_{r-1/n}^{n}}){\rm d}B_{r}^{H}\bigg{\rvert}^{p}$
-$\leq$ $3^{p-1}\bigg{[}\lambda^{p}\mathbb{E}|X_{t-\tau}^{n}-X_{s-\tau}^{n}|^{p}+(t-s)^{p-1}$
-×$\times\mathbb{E}\int_{s}^{t}C^{p}\bigg{(}1+|X_{r-1/n}^{n}|^{l+1}+|X_{r-\tau}^{n}|^{l+1}+\mathbb{W}_{q}(\mathcal{L}_{X_{r-1/n}^{n}},\delta_{0})\bigg{)}^{p}\bigg{]}{\rm d}r$
-$+C(t-s)^{pH-1}\mathbb{E}\int_{s}^{t}\bigg{(}1+\mathbb{W}_{q}(\mathcal{L}_{X_{r-1/n}^{n}},\delta_{0})\bigg{)}^{p}\bigg{]}.$
-For $0\leq s\leq t\leq\tau$, by Assumption 2.2, we get
-$\mathbb{E}|X_{t}^{n}-X_{s}^{n}|^{p}$
-$\leq$ $3^{p-1}\lambda^{p}\mathbb{E}|\xi(t-\tau)-\xi(s-\tau)|^{p}+C(t-s)^{p-1}\int_{s}^{t}\mathbb{E}\bigg{(}\sup_{u\in[-\tau,r]}|X_{u}^{n}|^{(l+1)p}\bigg{)}{\rm d}r$
-$+C(t-s)^{pH-1}\int_{s}^{t}\mathbb{E}\bigg{(}\sup_{u\in[-\tau,r]}|X_{u}^{n}|^{p}\bigg{)}{\rm d}r$
-$\leq$ $3^{p-1}\lambda^{p}K_{0}|t-s|^{p}+C(t-s)^{p-1}\int_{s}^{t}\mathbb{E}\bigg{(}\sup_{u\in[-\tau,r]}|X_{u}^{n}|^{(l+1)p}\bigg{)}{\rm d}r$
-$+C(t-s)^{pH-1}\int_{s}^{t}\mathbb{E}\bigg{(}\sup_{u\in[-\tau,r]}|X_{u}^{n}|^{p}\bigg{)}{\rm d}r.$
-Since $(l+1)p\leq\overline{p}$, we can applying Lemma 3.1 to obtain the assertion.
-For $\tau\leq s\leq t\leq T$, noting that $01-3^{p-1}\lambda^{p}>$ 0, we get the desired assertion by applying Lemma 3.1 again.
-∎
-Theorem 3.3.
-Let Assumption 2.4 hold. Then for $p>1/H$ and $(l+1)p\leq\overline{p}$, (2.5) admits a global unique solution.
-Proof.
-We split the proof into two steps.
-Step 1. Existence. For $m>n\geq 2/\tau$ and $t\in[0,T]$, by (2.5), we may compute
-$\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|X^{m}_{s}-X_{s}^{n}|^{p}\bigg{)}$
-$\leq$ $3^{p-1}\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|D(X_{s-\tau}^{m})-D(X_{s-\tau}^{n})|^{p}\bigg{)}$
-$+3^{p-1}\mathbb{E}\bigg{(}\sup_{s\in[0,t]}\bigg{\lvert}\int_{0}^{s}\bigg{[}b\big{(}X_{r-1/m}^{m},X_{r-\tau}^{m},\mathcal{L}_{X_{r-1/m}^{m}}\big{)}-b\big{(}X_{r-1/n}^{n},X_{r-\tau}^{n},\mathcal{L}_{X_{r-1/n}^{n}}\big{)}\bigg{]}{\rm d}r\bigg{\rvert}^{p}\bigg{)}$
-$+3^{p-1}\mathbb{E}\bigg{(}\sup_{s\in[0,t]}\bigg{\lvert}\int_{0}^{s}\bigg{[}\sigma\big{(}\mathcal{L}_{X_{r-1/m}^{m}}\big{)}-\sigma\big{(}\mathcal{L}_{X_{r-1/n}^{n}}\big{)}\bigg{]}{\rm d}B_{r}^{H}\bigg{\rvert}^{p}\bigg{)}$
-=:absent:=:= : $W_{1}+W_{2}+W_{3},$
-By Assumption 2.3, we get
-$\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|D(X_{s-\tau}^{m})-D(X_{s-\tau}^{n})|^{p}\bigg{)}\leq\lambda^{p}\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|X_{s-\tau}^{m}-X_{s-\tau}^{n}|^{p}\bigg{)}.$
-By the Hölder inequality, Assumption 2.4, we get
-$W_{2}=$ $3^{p-1}\mathbb{E}\bigg{(}\sup_{s\in[0,t]}\bigg{\lvert}\int_{0}^{s}\bigg{[}b\big{(}X_{r-1/m}^{m},X_{r-\tau}^{m},\mathcal{L}_{X_{r-1/m}^{m}}\big{)}-b\big{(}X_{r-1/n}^{n},X_{r-\tau}^{n},\mathcal{L}_{X_{r-1/n}^{n}}\big{)}\bigg{]}{\rm d}r\bigg{\rvert}^{p}\bigg{)}$
-$\leq$ $(3t)^{p-1}\mathbb{E}\int_{0}^{t}\bigg{\lvert}b\big{(}X_{r-1/m}^{m},X_{r-\tau}^{m},\mathcal{L}_{X_{r-1/m}^{m}}\big{)}-b\big{(}X_{r-1/n}^{n},X_{r-\tau}^{n},\mathcal{L}_{X_{r-1/n}^{n}}\big{)}\bigg{\rvert}^{p}{\rm d}r$
-$\leq$ $(3t)^{p-1}K_{3}\mathbb{E}\int_{0}^{t}\bigg{[}\bigg{(}1+|X_{r-1/m}^{m}|^{l}+|X_{r-1/n}^{n}|^{l}+|X_{r-\tau}^{m}|^{l}+|X_{r-\tau}^{n}|^{l}\bigg{)}$
-×$\times\bigg{(}|X_{r-1/m}^{m}-X_{r-1/n}^{n}|+|X_{r-\tau}^{m}-X_{s-\tau}^{n}|\bigg{)}+\mathbb{W}_{q}\big{(}\mathcal{L}_{X_{r-1/m}^{m}},\mathcal{L}_{X_{r-1/n}^{n}}\big{)}\bigg{]}^{p}{\rm d}r$
-$\leq$ $C\mathbb{E}\int_{0}^{t}\bigg{(}1+|X_{r-1/m}^{m}|^{l}+|X_{r-1/n}^{n}|^{l}+|X_{r-\tau}^{m}|^{l}+|X_{r-\tau}^{n}|^{l}\bigg{)}^{p}$
-×$\times\bigg{(}|X_{r-1/m}^{m}-X_{r-1/n}^{n}|+|X_{r-\tau}^{m}-X_{s-\tau}^{n}|\bigg{)}^{p}{\rm d}r$
-$+C\mathbb{E}\int_{0}^{t}|X_{r-1/m}^{m}-X_{r-1/n}^{n}|^{p}{\rm d}r.$
-Since for $l>1$, we have $2pl\leq(l+1)p\leq\overline{p}$, then by the Hölder inequality, Lemmas 3.1, 3.2, we get
-$W_{2}\leq$ $C\int_{0}^{t}\bigg{[}\mathbb{E}\bigg{(}1+|X_{r-1/m}^{m}|^{l}+|X_{r-1/n}^{n}|^{l}+|X_{r-\tau}^{m}|^{l}+|X_{r-\tau}^{n}|^{l}\bigg{)}^{2p}\bigg{]}^{1/2}$
-×$\times\bigg{[}\mathbb{E}\bigg{(}|X_{r-1/m}^{m}-X_{r-1/n}^{n}|+|X_{r-\tau}^{m}-X_{s-\tau}^{n}|\bigg{)}^{2p}\bigg{]}^{1/2}{\rm d}r+C\mathbb{E}\int_{0}^{t}|X_{r-1/m}^{m}-X_{r-1/n}^{n}|^{p}{\rm d}r$
-$\leq$ $C\int_{0}^{t}\bigg{[}\mathbb{E}\bigg{(}|X_{r-1/m}^{m}-X_{r-1/n}^{n}|+|X_{r-\tau}^{m}-X_{r-\tau}^{n}|\bigg{)}^{2p}\bigg{]}^{1/2}{\rm d}r$
-$+C\bigg{[}\bigg{(}\frac{1}{n}-\frac{1}{m}\bigg{)}^{p}+\bigg{(}\frac{1}{n}-\frac{1}{m}\bigg{)}^{pH}\bigg{]}.$
-By Assumption 2.5, Lemmas 3.1, 3.2 and the same techniques as (3.3), we get
-$W_{3}=$ $3^{p-1}\mathbb{E}\bigg{[}\sup_{s\in[0,t]}\bigg{\lvert}\int_{0}^{s}\bigg{(}\sigma(\mathcal{L}_{X_{r-1/m}^{m}})-\sigma(\mathcal{L}_{X_{r-1/n}^{n}})\bigg{)}{\rm d}B_{r}^{H}\bigg{\rvert}^{p}\bigg{]}$
-$\leq$ $C\mathbb{E}\int_{0}^{t}|\sigma(\mathcal{L}_{X_{r-1/m}^{m}})-\sigma(\mathcal{L}_{X_{r-1/n}^{n}})|^{p}{\rm d}r$
-$\leq$ $C\mathbb{E}\int_{0}^{t}K_{4}^{p}(\mathbb{W}_{q}(\mathcal{L}_{X_{r-1/m}^{m}}-\mathcal{L}_{X_{r-1/n}^{n}}))^{p}{\rm d}r$
-$\leq$ $C\int_{0}^{t}\mathbb{E}|X_{r-1/m}^{m}-X_{r-1/n}^{n}|^{p}{\rm d}r$
-$\leq$ $C\int_{0}^{t}\mathbb{E}|X_{r-1/m}^{m}-X_{r-1/n}^{m}|^{p}{\rm d}r+C\int_{0}^{t}\mathbb{E}|X_{r-1/n}^{m}-X_{r-1/n}^{n}|^{p}{\rm d}r$
-$\leq$ $C\bigg{[}\bigg{(}\frac{1}{n}-\frac{1}{m}\bigg{)}^{p}+\bigg{(}\frac{1}{n}-\frac{1}{m}\bigg{)}^{pH}\bigg{]}+C\int_{0}^{t}\mathbb{E}|X_{r-1/n}^{m}-X_{r-1/n}^{n}|^{p}{\rm d}r.$
-Combining $W_{1},W_{2},W_{3}$, we get
-$\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|X^{m}_{s}-X_{s}^{n}|^{p}\bigg{)}$
-$\leq$ $\lambda^{p}\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|X_{s-\tau}^{m}-X_{s-\tau}^{n}|^{p}\bigg{)}+C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{u\in[0,r]}|X_{u}^{m}-X_{u}^{n}|^{p}\bigg{)}{\rm d}r$
-$+C\bigg{[}\bigg{(}\frac{1}{n}-\frac{1}{m}\bigg{)}^{p}+\bigg{(}\frac{1}{n}-\frac{1}{m}\bigg{)}^{pH}\bigg{]}+C\int_{0}^{t}\bigg{[}\mathbb{E}\bigg{(}\sup_{u\in[0,r]}|X_{u-\tau}^{m}-X_{u-\tau}^{n}|\bigg{)}^{2p}\bigg{]}^{1/2}{\rm d}r.$
-Noting that $\lambda<1$, by the Gronwall inequality, we get
-$\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|X^{m}_{s}-X_{s}^{n}|^{p}\bigg{)}$
-$\leq$ $C\bigg{[}\bigg{(}\frac{1}{n}-\frac{1}{m}\bigg{)}^{p}+\bigg{(}\frac{1}{n}-\frac{1}{m}\bigg{)}^{pH}\bigg{]}+C\int_{0}^{t}\bigg{[}\mathbb{E}\bigg{(}\sup_{u\in[0,r]}|X_{u-\tau}^{m}-X_{u-\tau}^{n}|\bigg{)}^{2p}\bigg{]}^{1/2}{\rm d}r.$
-Define a sequence $\{u_{1},u_{2},\cdots,u_{[T/\tau]+1}\}$ where $u_{i}=([T/\tau]+2-i)p2^{[T/\tau]+1-i}$, and $[T/\tau]$ denotes the integer part of $T/\tau$.
-It is easy to see that $u_{i}\geq p$, $u_{[T/\tau]+1}=p$ and $2u_{i+1}<u_{i}$ for $i=1,2,\cdots,[T/\tau]$.
-For $t\in[0,\tau]$, we get
-(3.4) $\mathbb{E}\bigg{(}\sup_{s\in[0,\tau]}|X^{m}_{s}-X_{s}^{n}|^{u_{1}}\bigg{)}\to 0,\ {\rm for}\ m,n\to\infty.$
-For $t\in[0,2\tau]$, by the Hölder inequality, we get
-$\mathbb{E}\bigg{(}\underset{s\in[0,2\tau]}{\sup}|X^{m}_{s}-X_{s}^{n}|^{u_{2}}\bigg{)}$
-[…]
-by (3.4),
-$\mathbb{E}\bigg{(}\underset{s\in[0,2\tau]}{\sup}|X^{m}_{s}-X_{s}^{n}|^{u_{2}}\bigg{)}\to 0,\ {\rm for}\ m,n\to\infty.$
-By induction, we can see that for $t\in[0,T]$,
-(3.5) $\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|X^{m}_{s}-X_{s}^{n}|^{p}\bigg{)}\to 0,\ {\rm for}\ m,n\to\infty.$
-Consequently, $(X^{n})_{n\geq 2/\tau}$ is a Cauchy sequence for $p>1/H$. Denote the limit of $(X^{n})$ by $X$ and letting $m\to\infty$ in (3.5),
-(3.6) $\lim_{n\to\infty}\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|X_{s}-X_{s}^{n}|^{p}\bigg{)}=0.$
-Then, we should show that $X_{t}$ is a solution to (2.5). For $t\in[0,T]$, we get
-$\mathbb{E}|X_{t}-X^{n}_{t-1/n}|^{p}=$ $\mathbb{E}|X_{t}-X^{n}_{t}+X^{n}_{t}-X^{n}_{t-1/n}|^{p}$
-$\leq$ $2^{p-1}\mathbb{E}|X_{t}-X^{n}_{t}|^{p}+2^{p-1}\mathbb{E}|X^{n}_{t}-X^{n}_{t-1/n}|^{p}.$
-By Lemma 3.2 and (3.6), we get
-$\mathbb{E}|X_{t}-X^{n}_{t-1/n}|^{p}\to 0,\ \ {\rm as}\ \ n\to\infty.$
-Taking $n\to\infty$ in (3.1), we get
-$X_{t}-D(X_{t-\tau})=\xi(0)+D(\xi(-\tau))+\int_{0}^{t}b(X_{s},X_{s-\tau},\mathcal{L}_{X_{s}}){\rm d}s+\int_{0}^{t}\sigma(\mathcal{L}_{X_{s}}){\rm d}B_{s}^{H},\ \ t\in[0,T].$
-Step 2. Uniqueness. Let $X_{t}$ and $Y_{t}$ be two solutions of (2.5) with $X_{t}=Y_{t}=\xi,\ \ t\in[-\tau,0]$. By Lemma 2.7 again, we get
-$|X_{t}-Y_{t}|^{p}\leq\lambda|X_{t-\tau}-Y_{t-\tau}|^{p}+\frac{1}{(1-\lambda)^{p-1}}|X_{t}-D(X_{t-\tau})+Y_{t}-D(Y_{t-\tau})|^{p},$
-which implies,
-$\sup_{0\leq r\leq t}|X_{r}-Y_{r}|^{p}$
-$\leq$ $\sup_{-\tau\leq r\leq 0}|X_{r}-Y_{r}|^{p}+\lambda\sup_{0\leq r\leq t}|X_{r}-Y_{r}|^{p}+\frac{1}{(1-\lambda)^{p-1}}\sup_{0\leq r\leq t}|X_{r}-D(X_{r-\tau})+Y_{r}-D(Y_{r-\tau})|^{p}.$
-Since $X_{t}=Y_{t}=\xi,\ \ t\in[-\tau,0]$ and $\lambda\in(0,1)$,
-$\sup_{0\leq r\leq t}|X_{r}-Y_{r}|^{p}\leq\frac{1}{(1-\lambda)^{p}}\sup_{0\leq r\leq t}|X_{r}-D(X_{r-\tau})+Y_{r}-D(Y_{r-\tau})|^{p}=\frac{1}{(1-\lambda)^{p}}\sup_{0\leq r\leq t}|\Pi(r)|^{p},$
-where $\Pi(r)=X_{r}-D(X_{r-\tau})+Y_{r}-D(Y_{r-\tau})$. For $t\in[0,T]$, by the Itô formula and the elementary inequality, we get
-$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}|\Pi(s)|^{p}\bigg{)}\leq$ $p\mathbb{E}\int_{0}^{t}|\Pi(r)|^{p-2}\langle\Pi(r),b(X_{r},X_{r-\tau},\mathcal{L}_{X_{r}})-b(Y_{r},Y_{r-\tau},\mathcal{L}_{Y_{r}})\rangle{\rm d}r$
-$+p(p-1)H\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}\int_{0}^{s}|\Pi(r)|^{p-2}|\sigma(\mathcal{L}_{X_{r}})-\sigma(\mathcal{L}_{Y_{r}})|^{2}{\rm d}r\bigg{)}$
-$+p\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}\int_{0}^{s}|\Pi(r)|^{p-1}|\sigma(\mathcal{L}_{X_{r}})-\sigma(\mathcal{L}_{Y_{r}})|{\rm d}B_{r}^{H}\bigg{)}.$
-$=$ $J_{1}+J_{2}+J_{3}.$
-By Assumption 2.4, the Young’s inequality, we get
-$J_{1}=$ $C\mathbb{E}\int_{0}^{t}|X_{r}-D(X_{r-\tau})+Y_{r}-D(Y_{r-\tau})|^{p}{\rm d}r$
-$+C\mathbb{E}\int_{0}^{t}\bigg{[}|X_{r}-Y_{r}|^{2}+|X_{r-\tau}-Y_{r-\tau}|^{2}+(\mathbb{W}_{q}(\mathcal{L}_{X_{r}},\mathcal{L}_{Y_{r}}))^{2}\bigg{]}^{p/2}{\rm d}r$
-$\leq$ $C\mathbb{E}\int_{0}^{t}\bigg{(}\sup_{0\leq u\leq r}|X_{u}-Y_{u}|^{p}\bigg{)}{\rm d}r+C\mathbb{E}\int_{0}^{t}\bigg{(}\sup_{0\leq u\leq r}|X_{u-\tau}-Y_{u-\tau}|^{p}\bigg{)}{\rm d}r.$
-By Assumption 2.5 and the Young’s inequality, we get
-$J_{2}\leq$ $C\mathbb{E}\int_{0}^{t}|X_{r}-D(X_{r-\tau})+Y_{r}-D(Y_{r-\tau})|^{p}{\rm d}r+C\mathbb{E}\int_{0}^{t}\bigg{[}(\mathbb{W}_{q}(\mathcal{L}_{X_{r}},\mathcal{L}_{Y_{r}}))^{2}\bigg{]}^{p/2}{\rm d}r$
-$\leq$ $C\mathbb{E}\int_{0}^{t}\bigg{(}\sup_{0\leq u\leq r}|X_{u}-Y_{u}|^{p}\bigg{)}{\rm d}r+C\mathbb{E}\int_{0}^{t}\bigg{(}\sup_{0\leq u\leq r}|X_{u-\tau}-Y_{u-\tau}|^{p}\bigg{)}{\rm d}r.$
-In the way in (3.3), by Assumption 2.5, the Young’s inequality, the Hölder inequality, we get
-$J_{3}\leq$ $C\mathbb{E}\int_{0}^{t}\bigg{(}\sup_{0\leq u\leq r}|X_{u}-Y_{u}|^{p}\bigg{)}{\rm d}r+C\mathbb{E}\int_{0}^{t}\bigg{(}\sup_{0\leq u\leq r}|X_{u-\tau}-Y_{u-\tau}|^{p}\bigg{)}{\rm d}r.$
-Combining $J_{1}$-$J_{3}$ and applying the Gronwall inequality, we get
-$\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|X_{s}-Y_{s}|^{p}\bigg{)}\leq C\mathbb{E}\int_{0}^{t}\bigg{(}\sup_{0\leq u\leq r}|X_{u-\tau}-Y_{u-\tau}|^{p}\bigg{)}{\rm d}r.$
-By induction, we derive that $X_{t}=Y_{t},t\in[0,T]$, $\mathbb{P}$-a.s. The proof is complete. ∎
+$\mathbb{E}\bigg(\sup_{0\leq s\leq t\wedge\rho_{N}^{(2)}}Q_{3}^{(2)}(s)\bigg)\leq$ $C\epsilon\mathbb{E}\left(\sup_{0\leq s\leq t\wedge\rho_{N}^{(2)}}\left|X_{s}^{(2)}\right|^{\overline{p}}\right)+C_{\epsilon}\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r\wedge\rho_{N}^{(2)}}|X_{s}^{(2)}|^{\overline{p}}\bigg){\rm d}r$
+$+C\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r\wedge\rho_{N}^{(2)}}|X_{s}^{(1)}|^{\overline{p}}\bigg){\rm d}r.$
+Consequently, choosing $\epsilon$ sufficiently small so that $C\epsilon<\frac{1}{2}$ and using the Fatou lemma, the Gronwall inequality, it holds
+$\mathbb{E}\bigg(\sup_{0\leq s\leq t}|X_{s}^{(2)}|^{\overline{p}}\bigg)\leq C.$
+Finally, (3.1) can be shown by induction.
+Step 2.
+Existence Define the difference processes $\Delta_{s}^{(k+1),(k)}=X^{(k+1)}_{s}-X_{s}^{(k)}$, $\Upsilon_{s}^{(k+1),(k)}=X^{(k+1)}_{s}-D(X_{s-\tau}^{(k+1)})-X_{s}^{(k)}+D(X_{s-\tau}^{(k)})$.
+By Lemma 2.7, we have
+$|\Delta_{s}^{(k+1),(k)}|^{\overline{p}}\leq\lambda|\Delta_{s-\tau}^{(k+1),(k)}|^{\overline{p}}+\frac{1}{(1-\lambda)^{\overline{p}-1}}|\Upsilon_{s}^{(k+1),(k)}|^{\overline{p}},$
+Taking supremum and expectation gives
+(3.5) $\mathbb{E}\left(\sup_{0\leq s\leq t}|\Delta_{s}^{(k+1),(k)}|^{\overline{p}}\right)\leq\frac{1}{(1-\lambda)^{\overline{p}}}\mathbb{E}\left(\sup_{0\leq s\leq t}|\Upsilon_{s}^{(k+1),(k)}|^{\overline{p}}\right).$
+For $t\in[0,T]$, employment of the fractional Itô formula in Lemma 2.1 and the elementary inequality provides
+$\mathbb{E}\bigg(\sup_{0\leq s\leq t}|\Upsilon_{s}^{(k+1),(k)}|^{\overline{p}}\bigg)$
+$\leq$ ${\overline{p}}\mathbb{E}\int_{0}^{t}|\Upsilon_{r}^{(k+1),(k)}|^{{\overline{p}}-2}\langle\Upsilon_{r}^{(k+1),(k)},b(X^{(k+1)}_{r},X^{(k+1)}_{r-\tau},\mathcal{L}_{X^{(k)}_{r}})-b(X^{(k)}_{r},X^{(k)}_{r-\tau},\mathcal{L}_{X^{(k-1)}_{r}})\rangle{\rm d}r$
+$+\mathbb{E}\bigg(\sup_{0\leq s\leq t}\bigg\lvert\int_{0}^{s}{\overline{p}}({\overline{p}}-1)H(2H-1)|\Upsilon_{r}^{(k+1),(k)}|^{{\overline{p}}-2}$
+$\times\bigg(\int_{0}^{r}(r-u)^{2H-2}\|\sigma(\mathcal{L}_{X^{(k)}_{u}})-\sigma(\mathcal{L}_{X^{(k-1)}_{u}})\|^{2}{\rm d}u\bigg){\rm d}r\bigg)$
+$+{\overline{p}}\mathbb{E}\bigg(\sup_{0\leq s\leq t}\int_{0}^{s}|\Upsilon_{r}^{(k+1),(k)}|^{{\overline{p}}-1}\|\sigma(\mathcal{L}_{X^{(k)}_{r}})-\sigma(\mathcal{L}_{X^{(k-1)}_{r}})\|{\rm d}B_{r}^{H}\bigg).$
+$:=$ $J_{1}(t)+J_{2}(t)+J_{3}(t).$
+Applying Assumption 2.4 and the Young’s inequality produces
+$J_{1}(t)=$ ${\overline{p}}\mathbb{E}\int_{0}^{t}|\Upsilon_{r}^{(k+1),(k)}|^{{\overline{p}}-2}\langle\Upsilon_{r}^{(k+1),(k)},b(X^{(k+1)}_{r},X^{(k+1)}_{r-\tau},\mathcal{L}_{X^{(k)}_{r}})-b(X^{(k)}_{r},X^{(k)}_{r-\tau},\mathcal{L}_{X^{(k-1)}_{r}})\rangle{\rm d}r$
+$\leq$ $C\mathbb{E}\int_{0}^{t}|X^{(k+1)}_{r}-D(X^{(k+1)}_{r-\tau})+X^{(k)}_{r}-D(X^{(k)}_{r-\tau})|^{\overline{p}}{\rm d}r$
+$+C\mathbb{E}\int_{0}^{t}\bigg[|X^{(k+1)}_{r}-X^{(k)}_{r}|^{2}+|X^{(k+1)}_{r-\tau}-X^{(k)}_{r-\tau}|^{2}+\mathbb{W}_{2}^{2}(\mathcal{L}_{X^{(k)}_{r}},\mathcal{L}_{X^{(k-1)}_{r}})\bigg]^{{\overline{p}}/2}{\rm d}r$
+$\leq$ $C\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|\Delta_{s}^{(k+1),(k)}|^{\overline{p}}\bigg){\rm d}r+C\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|\Delta_{s}^{(k),(k-1)}|^{\overline{p}}\bigg){\rm d}r,$
+where we used the fact that $\mathbb{W}_{2}^{2}(\mathcal{L}_{X^{(k)}_{r}},\mathcal{L}_{X^{(k-1)}_{r}})\leq\mathbb{E}^{\mathbb{P}}|X^{(k)}_{r}-X^{(k-1)}_{r}|^{2}$.
+Along with Assumption 2.5, the Young’s inequality, the Hölder inequality, and the Fubini theorem, it follows
+$J_{2}(t)=$ $\mathbb{E}\bigg(\sup_{0\leq s\leq t}\bigg\lvert\int_{0}^{s}{\overline{p}}({\overline{p}}-1)H(2H-1)|\Upsilon_{r}^{(k+1),(k)}|^{{\overline{p}}-2}$
+$\times\left(\int_{0}^{r}(r-u)^{2H-2}\|\sigma(\mathcal{L}_{X^{(k)}_{u}})-\sigma(\mathcal{L}_{X^{(k-1)}_{u}})\|^{2}{\rm d}u\bigg){\rm d}r\right)$
+$\leq$ $C\int_{0}^{t}\int_{0}^{r}(r-u)^{2H-2}(\mathbb{E}|\Upsilon_{r}^{(k+1),(k)}|^{\overline{p}}+\mathbb{E}|X^{(k)}_{u}-X^{(k-1)}_{u}|^{\overline{p}}){\rm d}u{\rm d}r$
+$\leq$ $C\int_{0}^{t}\mathbb{E}|\Upsilon_{r}^{(k+1),(k)}|^{\overline{p}}\bigg(\int_{0}^{r}(r-u)^{2H-2}{\rm d}u\bigg){\rm d}r+C\int_{0}^{t}\int_{u}^{t}(r-u)^{2H-2}\mathbb{E}|X^{(k)}_{u}-X^{(k-1)}_{u}|^{\overline{p}}{\rm d}r{\rm d}u$
+$\leq$ $C\int_{0}^{t}\frac{r^{2H-1}}{2H-1}\mathbb{E}|\Upsilon_{r}^{(k+1),(k)}|^{\overline{p}}{\rm d}r+C\int_{0}^{t}\frac{(t-u)^{2H-1}}{2H-1}\mathbb{E}|X^{(k)}_{u}-X^{(k-1)}_{u}|^{\overline{p}}{\rm d}u.$
+Since $1/2<H<1$ and $t\in[0,T]$, we use Assumption 2.3 to get
+$J_{2}(t)\leq$ $C\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|\Delta_{s}^{(k+1),(k)}|^{\overline{p}}\bigg){\rm d}r+C\int_{0}^{t}(t-r)^{2H-1}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|\Delta_{s}^{(k),(k-1)}|^{\overline{p}}\bigg){\rm d}r.$
+In the similar way as dealing with $Q_{3}^{(1)}(t)$ to get (3.4), an application of fractional BDG inequality (3.3) with $p=1$ yields
+$J_{3}(t)=$ ${\overline{p}}\mathbb{E}\left[\sup_{0\leq s\leq t}\left|\int_{0}^{s}|\Upsilon_{r}^{(k+1),(k)}|^{{\overline{p}}-1}\|\sigma(\mathcal{L}_{X^{(k)}_{r}})-\sigma(\mathcal{L}_{X^{(k-1)}_{r}})\|{\rm d}B_{r}^{H}\right|\right]$
+$\leq$ $C\mathbb{E}\left[\left(\int_{0}^{t}\left(|\Upsilon_{r}^{(k+1),(k)}|^{{\overline{p}}-1}\|\sigma(\mathcal{L}_{X^{(k)}_{r}})-\sigma(\mathcal{L}_{X^{(k-1)}_{r}})\|\right)^{1/H}{\rm d}r\right)^{H}\right]$
+For each $r\in[0,t]$, by the Young’s inequality, there exists $0<\epsilon<1$ such that
+$|\Upsilon_{r}^{(k+1),(k)}|^{{\overline{p}}-1}\|\sigma(\mathcal{L}_{X^{(k)}_{r}})-\sigma(\mathcal{L}_{X^{(k-1)}_{r}})\|\leq\epsilon|\Upsilon_{r}^{(k+1),(k)}|^{{\overline{p}}}+C_{\epsilon}\|\sigma(\mathcal{L}_{X^{(k)}_{r}})-\sigma(\mathcal{L}_{X^{(k-1)}_{r}})\|^{{\overline{p}}}.$
+Assumption 2.5 implies that
+$\|\sigma(\mathcal{L}_{X^{(k)}_{r}})-\sigma(\mathcal{L}_{X^{(k-1)}_{r}})\|\leq K_{4}\mathbb{W}_{2}(\mathcal{L}_{X^{(k)}_{r}},\mathcal{L}_{X^{(k-1)}_{r}})\leq K_{4}(\mathbb{E}|\Delta_{r}^{(k),(k-1)}|^{2})^{1/2}.$
+Since ${\overline{p}}\geq 2$, using the fact that $(\mathbb{E}|\Upsilon_{r}^{(k),(k-1)}|^{2})^{{\overline{p}}/2}\leq\mathbb{E}|\Upsilon_{r}^{(k),(k-1)}|^{{\overline{p}}}$, then
+$|\Upsilon_{r}^{(k+1),(k)}|^{{\overline{p}}-1}\|\sigma(\mathcal{L}_{X^{(k)}_{r}})-\sigma(\mathcal{L}_{X^{(k-1)}_{r}})\|\leq\epsilon|\Upsilon_{r}^{(k+1),(k)}|^{{\overline{p}}}+C_{\epsilon}(\mathbb{E}|\Delta_{r}^{(k),(k-1)}|^{2})^{{\overline{p}}/2}$
+$\leq$ $\epsilon|\Upsilon_{r}^{(k+1),(k)}|^{{\overline{p}}}+C_{\epsilon}\mathbb{E}|\Delta_{r}^{(k),(k-1)}|^{{\overline{p}}}.$
+Noting that for $\frac{1}{H}>1$, we can apply the elementary inequality to get
+$\left(\int_{0}^{t}\left(\epsilon|\Upsilon_{r}^{(k+1),(k)}|^{{\overline{p}}}+C_{\epsilon}\mathbb{E}|\Delta_{r}^{(k),(k-1)}|^{{\overline{p}}}\right)^{1/H}{\rm d}r\right)^{H}$
+$\leq$ $2^{1-H}\epsilon\left(\int_{0}^{t}|\Upsilon_{r}^{(k+1),(k)}|^{{\overline{p}}/H}{\rm d}r\right)^{H}+C_{\epsilon}2^{1-H}\left(\int_{0}^{t}\left(\mathbb{E}|\Delta_{r}^{(k),(k-1)}|^{{\overline{p}}}\right)^{1/H}{\rm d}r\right)^{H}.$
+Taking expectation, it holds
+$\mathbb{E}\left[\left(\int_{0}^{t}|\Upsilon_{r}^{(k+1),(k)}|^{{\overline{p}}/H}{\rm d}r\right)^{H}\right]\leq t^{H}\mathbb{E}\left(\sup_{0\leq r\leq t}|\Upsilon_{r}^{(k+1),(k)}|^{{\overline{p}}}\right),$
+and
+$\mathbb{E}\left[\left(\int_{0}^{t}\left(\mathbb{E}|\Delta_{r}^{(k),(k-1)}|^{{\overline{p}}}\right)^{1/H}{\rm d}r\right)^{H}\right]\leq Ct^{H}\int_{0}^{t}\mathbb{E}|\Delta_{r}^{(k),(k-1)}|^{{\overline{p}}}{\rm d}r.$
+Combining the estimates together, we conclude
+(3.6) $J_{3}(t)\leq$ $C\epsilon\mathbb{E}\left(\sup_{0\leq s\leq t}|\Upsilon_{s}^{(k+1),(k)}|^{{\overline{p}}}\right)+C_{\epsilon}\int_{0}^{t}\mathbb{E}\left(\sup_{0\leq s\leq r}|\Delta_{s}^{(k),(k-1)}|^{{\overline{p}}}\right){\rm d}r.$
+Finally, the estimation of $J_{3}(t)$ combined with the estimates for the drift terms $J_{1}(t)$ and $J_{2}(t)$, we obtain
+$\mathbb{E}\left(\sup_{0\leq s\leq t}|\Upsilon_{s}^{(k+1),(k)}|^{\overline{p}}\right)$
+$\leq$ $C\epsilon\mathbb{E}\left(\sup_{0\leq s\leq t}|\Upsilon_{s}^{(k+1),(k)}|^{\overline{p}}\right)+C\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|\Delta_{s}^{(k+1),(k)}|^{\overline{p}}\bigg){\rm d}r$
+$+C\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|\Delta_{s}^{(k),(k-1)}|^{\overline{p}}\bigg){\rm d}r+C\int_{0}^{t}(t-r)^{2H-1}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|\Delta_{s}^{(k),(k-1)}|^{\overline{p}}\bigg){\rm d}r.$
+Choosing $\epsilon$ small enough so that $C\epsilon<\frac{1}{2}$ and absorbing the first term, it implies
+$\mathbb{E}\left(\sup_{0\leq s\leq t}|\Upsilon_{s}^{(k+1),(k)}|^{\overline{p}}\right)\leq$ $C\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|\Delta_{s}^{(k+1),(k)}|^{\overline{p}}\bigg){\rm d}r+C\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|\Delta_{s}^{(k),(k-1)}|^{\overline{p}}\bigg){\rm d}r$
+$+C\int_{0}^{t}(t-r)^{2H-1}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|\Delta_{s}^{(k),(k-1)}|^{\overline{p}}\bigg){\rm d}r.$
+Since singular kernels $\int_{0}^{t}(t-r)^{2H-1}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|\Delta_{s}^{(k),(k-1)}|^{\overline{p}}\bigg){\rm d}r$ are typically “stronger” than regular kernels $\int_{0}^{t}(t-r)^{0}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|\Delta_{s}^{(k),(k-1)}|^{\overline{p}}\bigg){\rm d}r$, we derive from (3.5) that
+$\mathbb{E}\bigg(\sup_{0\leq s\leq t}|\Delta_{s}^{(k+1),(k)}|^{\overline{p}}\bigg)$
+$\leq$ $C\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|\Delta_{s}^{(k+1),(k)}|^{\overline{p}}\bigg){\rm d}r+C\int_{0}^{t}(t-r)^{2H-1}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|\Delta_{s}^{(k),(k-1)}|^{\overline{p}}\bigg){\rm d}r.$
+Using the Gronwall inequality leads to
+$\mathbb{E}\bigg(\sup_{0\leq s\leq t}|\Delta_{s}^{(k+1),(k)}|^{\overline{p}}\bigg)\leq Ce^{Ct}\int_{0}^{t}(t-r)^{2H-1}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|\Delta_{s}^{(k),(k-1)}|^{\overline{p}}\bigg){\rm d}r.$
+By induction, we arrive at
+$\mathbb{E}\bigg(\sup_{0\leq s\leq t}|\Delta_{s}^{(k+1),(k)}|^{\overline{p}}\bigg)\leq\frac{(Ce^{Ct}\Gamma(2H)t^{2H})^{k}}{\Gamma(2kH+1)}\mathbb{E}\bigg(\sup_{0\leq s\leq t}|\Delta_{s}^{(1),(0)}|^{\overline{p}}\bigg).$
+With the properties of the Gamma function and the fact that $\mathbb{E}\bigg(\sup\limits_{0\leq s\leq t}|\Delta_{s}^{(1),(0)}|^{\overline{p}}\bigg)\leq C$ derived by (3.1), we can show that $\{X_{t}^{(k)}\}$ is a Cauchy sequence, then the limit, denoted by $X$, is a solution of (2.5).
+Step 3. Uniqueness Let $X_{t}$ and $Y_{t}$ be two solutions of (2.5) with $X_{t}=Y_{t}=\xi,\ \ t\in[-\tau,0]$. By Lemma 2.7 again, we establish
+$\sup_{0\leq r\leq t}|X_{r}-Y_{r}|^{\overline{p}}\leq\frac{1}{(1-\lambda)^{\overline{p}}}\sup_{0\leq r\leq t}|\Upsilon_{r}|^{\overline{p}},$
+where $\Upsilon_{r}=X_{r}-D(X_{r-\tau})-Y_{r}+D(Y_{r-\tau})$. Similar to the process of Step 2, it holds that
+$\mathbb{E}\bigg(\sup_{0\leq s\leq t}|X_{s}-Y_{s}|^{\overline{p}}\bigg)\leq$ $C\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|X_{s}-Y_{s}|^{\overline{p}}\bigg){\rm d}r+C\int_{0}^{t}(t-r)^{2H-1}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|X_{s}-Y_{s}|^{\overline{p}}\bigg){\rm d}r.$
+Further, we derive
+$\mathbb{E}\bigg(\sup_{0\leq s\leq t}|X_{s}-Y_{s}|^{\overline{p}}\bigg)\leq C\int_{0}^{t}(t-r)^{2H-1}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|X_{s}-Y_{s}|^{\overline{p}}\bigg){\rm d}r.$
+Although the kernel $(t-r)^{2H-1}$ is singular as 2H1>02H-1>0, it is integrable on [0, T].
+Therefore, applying Lemma 2.8 with $a=0$, $\beta=2H$ and $\gamma=1$, we find that the generalized Mittag-Leffler function is bounded, guaranteeing a unique zero solution for the inequality.
+That is,
+$\mathbb{E}\bigg(\sup_{0\leq s\leq t}|X_{s}-Y_{s}|^{\overline{p}}\bigg)=0.$
+Therefore, $X_{t}=Y_{t}$ almost surely for all $t\in[0,T]$. This completes the proof. ∎
+Remark 3.2.
+For the existence and uniqueness of the McKean-Vlasov NSDDEs, we adopt the Picard approximation method to construct a Cauchy sequence and show the existence and uniqueness of the solution.
+Since the super-linearly growing coefficients in our proposed model introduces additional technical complexities that demand refined analytical tools, our analysis differs fundamentally from those of related works.
+For McKean-Vlasov SDEs driven by standard Brownian motions (e.g., [17]) or by fractional Brownian motions without super-linearly growing coefficients (e.g., [18, 19]), the uniqueness is typically proven via direct estimates of the difference between two candidate solutions.
+In contrast, our model involves fractional noise and super-linear coefficients, which necessitates the use of the fractional Itô formula to properly account for the non-Markovian and rough nature of the fractional Brownian motions.
+This application of the fractional Itô formula not only resolves the technical hurdles posed by the fractional noise but also provides a generalizable framework for analyzing McKean-Vlasov SDEs driven by fractional Brownian motions.
 4 Propagation of Chaos
 Since the coefficients of (2.5) depend on the distributions of state variables, we use stochastic interacting particle systems to approximate it.
-Firstly, considering a non-interacting particle system is given by
+Firstly, considering a non-interacting particle system given by
 (4.1) ${\rm d}(X_{t}^{i}-D(X_{t-\tau}^{i}))=b(X_{t}^{i},X_{t-\tau}^{i},\mathcal{L}_{X_{t}^{i}}){\rm d}t+\sigma(\mathcal{L}_{X_{t}^{i}}){\rm d}B_{t}^{H,i},\ \ X_{0}^{i}=Y_{0}^{i}=\xi,\ \ i=1,2,\cdots,N,$
@@
 where the empirical measures is defined by
-$\mu^{X,N}_{t}(\cdot)=\frac{1}{N}\sum_{j=1}^{N}\delta_{X_{t}^{j,N}}(\cdot),$
+$\mu^{X,N}_{t}=\frac{1}{N}\sum_{i=1}^{N}\delta_{X_{t}^{i,N}},$
 here, $\delta_{x}$ denotes the Dirac measure at point $x$. In order to prove the desired result, we give the following Lemma.
 Lemma 4.1.
-[17, Lemma 3.2] For two empirical measures $\mu_{t}^{N}=\frac{1}{N}\sum\limits_{j=1}\limits^{N}\delta_{Z^{1,j,N}_{t}}$ and $\nu_{t}^{N}=\frac{1}{N}\sum\limits_{j=1}\limits^{N}\delta_{Z_{t}^{2,j,N}}$ , then
-$\mathbb{E}\left(\mathbb{W}_{q}(\mu_{t}^{N},\nu_{t}^{N})\right)\leq\mathbb{E}\bigg{(}\frac{1}{N}\sum_{j=1}^{N}\bigg{\lvert}Z_{t}^{1,j,N}-Z_{t}^{2,j,N}\bigg{\rvert}^{q}\bigg{)}^{1/q}.$
+[16, Lemma 3.2] For two empirical measures $\mu_{t}^{N}=\frac{1}{N}\sum\limits_{i=1}\limits^{N}\delta_{Z^{1,i,N}_{t}}$ and $\nu_{t}^{N}=\frac{1}{N}\sum\limits_{i=1}\limits^{N}\delta_{Z_{t}^{2,i,N}}$, then
+$\mathbb{E}\mathbb{W}_{2}(\mu_{t}^{N},\nu_{t}^{N})\leq\mathbb{E}\bigg(\frac{1}{N}\sum_{i=1}^{N}\bigg\lvert Z_{t}^{1,i,N}-Z_{t}^{2,i,N}\bigg\rvert^{2}\bigg)^{1/2}.$
 Next, we consider the theory of the propagation of chaos.
 Lemma 4.2.
-Let Assumptions 2.2-2.5 hold. For $\overline{p}>1/H$, we have
+Let Assumptions 2.2-2.5 hold. For $\overline{p}\geq 2$, we have
 $\mathbb{E}\bigg(\sup_{t\in[-\tau,T]}|X_{t}^{i}|^{\overline{p}}\bigg)+\mathbb{E}\bigg(\sup_{t\in[-\tau,T]}|X_{t}^{i,N}|^{\overline{p}}\bigg)\leq C,\ \ i=1,2,\cdots,N,$
-where $C$ is a positive constant depending on ${\overline{p}},T,L,H,\xi$.
+where $C$ is a positive constant depending on ${\overline{p}},T,L,H,\xi$ but independent of $N$.
 Proof.
-We will proceed in the same way in Lemma 3.1. By Assumption 2.4 and Lemma 2.7, we get
-$|X_{t}^{i,N}|^{\overline{p}}=|D(X_{t-\tau}^{i,N})+X_{t}^{n}-D(X_{t-\tau}^{i,N})|^{\overline{p}}\leq\lambda|X_{t-\tau}^{i,N}|^{\overline{p}}+\frac{1}{(1-\lambda)^{{\overline{p}}-1}}|X_{t}^{i,N}-D(X_{t-\tau}^{i,N})|^{\overline{p}},$
-for any $t\in[0,T]$. We immediately get
-$\sup_{0\leq r\leq t}|X_{r}^{i,N}|^{\overline{p}}\leq\frac{\lambda}{1-\lambda}\parallel\xi\parallel^{\overline{p}}+\frac{1}{(1-\lambda)^{{\overline{p}}}}\sup_{0\leq r\leq t}|X_{r}^{i,N}-D(X_{r-\tau}^{i,N})|^{\overline{p}}$
-By the Itô formula, we get
+We will proceed the proof in the similar way as Theorem 3.1. Firstly, Assumption 2.3 and Lemma 2.7 imply that
+(4.3) $\sup_{0\leq s\leq t}|X_{s}^{i,N}|^{\overline{p}}\leq\frac{\lambda}{1-\lambda}\parallel\xi\parallel^{\overline{p}}+\frac{1}{(1-\lambda)^{{\overline{p}}}}\sup_{0\leq s\leq t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{\overline{p}}.$
+We now concentrate on the estimation of the second term. The application of the fractional Itô formula leads to
 $|X_{t}^{i,N}-D(X_{t-\tau}^{i,N})|^{\overline{p}}$
 $\leq$ $|\xi(0)-D(\xi(-\tau))|^{\overline{p}}$
-$+p\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{{\overline{p}}-2}\big{\langle}X_{s}^{i,N}-D(X_{s-\tau}),b\big{(}X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N}\big{)}\big{\rangle}{\rm d}s$
-$+{\overline{p}}\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{p-1}\sigma\big{(}\mu_{s}^{X,N}\big{)}{\rm d}B_{s}^{H}$
-$+{\overline{p}}({\overline{p}}-1)\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{{\overline{p}}-2}\sigma\big{(}\mu_{s}^{X,N}\big{)}D_{s}^{\phi}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|{\rm d}s$
-=:absent:=:= : $|\xi(0)-D(\xi(-\tau))|^{\overline{p}}+I_{1}(t)+I_{2}(t)+I_{3}(t).$
-Then,
-$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}|X_{t}^{i,N}-D(X_{t-\tau}^{i,N})|^{\overline{p}}\bigg{)}\leq\mathbb{E}|\xi(0)-D(\xi(-\tau))|^{\overline{p}}+\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}I_{1}+I_{2}+I_{3}\bigg{)}$
-By Assumptions 2.3, 2.4, and the Young’s inequality, we get
+$+\overline{p}\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{{\overline{p}}-2}\big\langle X_{s}^{i,N}-D(X_{s-\tau}),b\big(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N}\big)\big\rangle{\rm d}s$
+$+{\overline{p}}\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{\overline{p}-1}\|\sigma\big(\mu_{s}^{X,N}\big)\|{\rm d}B_{s}^{H}$
+$+{\overline{p}}({\overline{p}}-1)H(2H-1)\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{{\overline{p}}-2}\bigg(\int_{0}^{s}(s-r)^{2H-2}\|\sigma(\mu_{r}^{X,N})\|^{2}{\rm d}r\bigg){\rm d}s$
+$=:$ $|\xi(0)-D(\xi(-\tau))|^{\overline{p}}+I_{1}(t)+I_{2}(t)+I_{3}(t).$
+By means of Assumptions 2.3-2.4 and the Young’s inequality, we conclude
 $\mathbb{E}\bigg(\sup_{0\leq s\leq t}I_{1}(s)\bigg)$
 $=$ $\mathbb{E}\int_{0}^{t}{\overline{p}}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{{\overline{p}}-2}\big\langle X_{s}^{i,N}-D(X_{s-\tau}^{i,N}),b\big(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N}\big)\big\rangle{\rm d}s$
-$\leq$ $C{\overline{p}}\mathbb{E}\int_{0}^{t}[|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{p-2}\times(1+|X_{s}^{i,N}|^{2}+|X_{s-\tau}^{i,N}|^{2}+(\mathbb{W}_{q}(\mu_{s}^{X,N},\delta_{0})))^{2}]{\rm d}s$
-$\leq$ $C\mathbb{E}\int_{0}^{t}[|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{\overline{p}}+(1+|X_{s}^{i,N}|^{2}+|X_{s-\tau}^{i,N}|^{2}+(\mathbb{W}_{q}(\mu_{s}^{X,N},\delta_{0}))^{2})^{{\overline{p}}/2}]{\rm d}s$
-$\leq$ $C\mathbb{E}\int_{0}^{t\wedge\nu_{R}}[(1+|X_{s}^{i,N}|^{\overline{p}}+|X_{s-\tau}^{i,N}|^{\overline{p}}+\big{(}\mathbb{W}_{q}(\mu_{s}^{X,N},\delta_{0})\big{)}^{{\overline{p}}}]{\rm d}s.$
-By the Hölder inequality, Young’s inequality and Assumption 2.5 and in the way in (3.3), we get
-$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}I_{2}(s)\bigg{)}\leq$ $C\mathbb{E}\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{{\overline{p}}-1}|\sigma(\mu_{s}^{X,N})|{\rm d}s$
-$\leq$ $C\mathbb{E}\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{\overline{p}}{\rm d}s+C\mathbb{E}\int_{0}^{t}((\mathbb{W}_{q}(\mu_{s}^{X,N},\delta_{0}))^{\overline{q}}+1){\rm d}s.$
-By Assumption 2.5, Remark 2.1, the Young’s inequality, we get
+$\leq$ $C\mathbb{E}\int_{0}^{t}[|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{\overline{p}}+(1+|X_{s}^{i,N}|^{2}+|X_{s-\tau}^{i,N}|^{2}+\mathbb{W}_{2}^{2}(\mu_{s}^{X,N},\delta_{0}))^{{\overline{p}}/2}]{\rm d}s$
+$\leq$ $C\mathbb{E}\int_{0}^{t}[(1+|X_{s}^{i,N}|^{\overline{p}}+|X_{s-\tau}^{i,N}|^{\overline{p}}+\mathbb{W}_{2}^{\overline{p}}(\mu_{s}^{X,N},\delta_{0})]{\rm d}s.$
+Utilizing Assumption 2.5, the Hölder inequality, the Young’s inequality and the similar way as achieving (3.6), it produces
+$\mathbb{E}\bigg(\sup_{0\leq s\leq t}I_{2}(s)\bigg)=$ ${\overline{p}}\mathbb{E}\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{\overline{p}-1}\|\sigma\big(\mu_{s}^{X,N}\big)\|{\rm d}B_{s}^{H}$
+$\leq$ $C\epsilon\mathbb{E}\left(\sup_{0\leq s\leq t}|X_{s}^{i,N}|^{{\overline{p}}}\right)+C_{\epsilon}\mathbb{E}\int_{0}^{t}|X_{s-\tau}^{i,N}|^{\overline{p}}{\rm d}s+C\mathbb{E}\int_{0}^{t}\mathbb{W}_{2}^{\overline{p}}(\mu_{s}^{X,N},\delta_{0}){\rm d}s+C,$
+where $0<\epsilon<1$. Again by Assumption 2.5, the Young’s inequality and the Fubini theorem, it results in
 $\mathbb{E}\bigg(\sup_{0\leq s\leq t}I_{3}(s)\bigg)$
-$\leq$ $C\mathbb{E}\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{\overline{p}}{\rm d}s+C\mathbb{E}\int_{0}^{t}\bigg{[}|\sigma\big{(}\mu_{s}^{X,N}\big{)}|^{2}\int_{0}^{s}H(2H-1)|r-s|^{2H-2}{\rm d}r\bigg{]}^{{\overline{p}}/2}{\rm d}s$
-$\leq$ $C\mathbb{E}\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{\overline{p}}{\rm d}s+2CHT^{2H-1}\mathbb{E}\int_{0}^{t}\lvert\sigma\big{(}\mu_{s}^{X,N}\big{)}\rvert^{\overline{p}}{\rm d}s$
-$\leq$ $C\mathbb{E}\int_{0}^{t}[(|X_{s}^{n}|^{\overline{p}}+|X_{s-\tau}^{n}|^{\overline{p}}+\big{(}\mathbb{W}_{q}(\mu_{s}^{X,N},\delta_{0})\big{)}^{\overline{p}}]{\rm d}s$
-$=$ $C\mathbb{E}\int_{0}^{t}[(|X_{s}^{n}|^{\overline{p}}+|X_{s-\tau}^{n}|^{\overline{p}}+C\mathbb{E}\int_{0}^{t}\bigg{(}\frac{1}{N}\sum_{j=1}^{N}|X_{r}^{j,N}|^{q}\bigg{)}^{{\overline{p}}/q}{\rm d}r,$
-where for the last part, we uses the fact that
-$\mathbb{W}_{\theta}(\mu_{s}^{X,N},\delta_{0})=\mathcal{W}_{q}(\mu_{s}^{X,N})=\bigg{(}\frac{1}{N}\sum_{j=1}^{N}|X_{s}^{j,N}|^{q}\bigg{)}^{1/q}.$
-Since all $j$ are identically distributed, by the Minkowski inequality and Lemma 4.1, we get
-$\mathbb{E}\bigg{(}\frac{1}{N}\sum_{j=1}^{N}|X_{s}^{j,N}|^{q}\bigg{)}^{{\overline{p}}/q}\leq\bigg{(}\frac{1}{N}\sum_{j=1}^{N}\parallel|X_{s}^{j,N}|^{q}\parallel_{L^{1/q}}\bigg{)}^{\overline{p}/q}=\bigg{(}\frac{1}{N}\sum_{j=1}^{N}(\mathbb{E}|X_{s}^{j,N}|)^{q}\bigg{)}^{\overline{p}/q}=\mathbb{E}|X_{s}^{i,N}|^{\overline{p}}.$
+$=$ ${\overline{p}}({\overline{p}}-1)H(2H-1)\mathbb{E}\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{{\overline{p}}-2}\bigg(\int_{0}^{s}(s-r)^{2H-2}\|\sigma(\mu_{r}^{X,N})\|^{2}{\rm d}r\bigg){\rm d}s$
+$\leq$ $C\mathbb{E}\int_{0}^{t}|X_{s}^{i,N}|^{\overline{p}}{\rm d}s+C\mathbb{E}\int_{0}^{t}|X_{s-\tau}^{i,N}|^{\overline{p}}{\rm d}s+C\int_{0}^{t}(t-s)^{2H-1}\mathbb{E}|X_{s}^{i,N}|^{\overline{p}}{\rm d}s,$
+where we have used the fact that
+$\mathbb{W}_{2}(\mu_{s}^{X,N},\delta_{0})=\mathcal{W}_{2}(\mu_{s}^{X,N})=\bigg(\frac{1}{N}\sum_{i=1}^{N}|X_{s}^{i,N}|^{2}\bigg)^{1/2}.$
+Since all $X_{s}^{i,N}$ are identically distributed, the Minkowski inequality and Lemma 4.1 lead to
+$\mathbb{E}\bigg(\frac{1}{N}\sum_{i=1}^{N}|X_{s}^{i,N}|^{2}\bigg)^{{\overline{p}}/2}\leq\bigg(\frac{1}{N}\sum_{i=1}^{N}\parallel|X_{s}^{i,N}|^{2}\parallel_{L^{\overline{p}/2}}\bigg)^{\overline{p}/2}=\bigg(\frac{1}{N}\bigg(\sum_{i=1}^{N}(\mathbb{E}|X_{s}^{i,N}|^{\overline{p}})^{2/\overline{p}}\bigg)^{\overline{p}/2}=\mathbb{E}|X_{s}^{i,N}|^{\overline{p}}.$
 To sum up, we obtain
-$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}|X_{t}^{i,N}-D(X_{t-\tau}^{i,N})|^{\overline{p}}\bigg{)}\leq$ $C+C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{0\leq u\leq s}|X_{u}^{i,N}|^{\overline{p}}\bigg{)}{\rm d}s+C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{0\leq u\leq s}|X_{u-\tau}^{i,N}|^{\overline{p}}\bigg{)}{\rm d}s,$
-Thus, we get
-$\mathbb{E}\bigg{(}\sup_{0\leq r\leq t}|X_{r}^{i,N}|^{\overline{p}}\bigg{)}\leq$ $C+C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{0\leq u\leq s}|X_{u}^{i,N}|^{\overline{p}}\bigg{)}{\rm d}s+C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{0\leq u\leq s}|X_{u-\tau}^{i,N}|^{\overline{p}}\bigg{)}{\rm d}s.$
-For $t\in[0,\tau]$, we immediately get
-$\mathbb{E}\bigg{(}\sup_{0\leq r\leq t}|X_{r}^{i,N}|^{\overline{p}}\bigg{)}\leq C+C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{0\leq u\leq s\wedge\nu_{R}}|X_{u}^{i,N}|^{\overline{p}}\bigg{)}{\rm d}s,$
-since $\xi\in L_{\mathcal{F}_{0}}^{\overline{p}}([-\tau,0];\mathbb{R}^{d})$. We get the desired result by Gronwall inequality. By induction, we can derive the assertion. ∎
+$\mathbb{E}\bigg(\sup_{0\leq s\leq t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{\overline{p}}\bigg)\leq$ $C+C\epsilon\mathbb{E}\left(\sup_{0\leq s\leq t}|X_{s}^{i,N}|^{{\overline{p}}}\right)+C\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|X_{s}^{i,N}|^{\overline{p}}\bigg){\rm d}r$
+$+C\int_{0}^{t}(t-r)^{2H-1}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|X_{s}^{i,N}|^{\overline{p}}\bigg){\rm d}r.$
+Consequently, choosing some $\epsilon$ small enough to ensure $C\epsilon<\frac{1}{2}$ and using (4.3) to get
+(4.4) $\mathbb{E}\bigg(\sup_{0\leq s\leq t}|X_{s}^{i,N}|^{\overline{p}}\bigg)\leq$ $C+C\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|X_{s}^{i,N}|^{\overline{p}}\bigg){\rm d}r+C\int_{0}^{t}(t-r)^{2H-1}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|X_{s}^{i,N}|^{\overline{p}}\bigg){\rm d}r.$
+Since the singular kernel $(t-s)^{2H-1}$ is integrable on the finite interval $[0,T]$, we can bound
+$\int_{0}^{t}(t-s)^{2H-1}{\rm d}s\leq C.$
+Rewriting (4.4) by ignoring the regular kernel as
+$\mathbb{E}\bigg(\sup_{0\leq s\leq t}|X_{s}^{i,N}|^{\overline{p}}\bigg)\leq C+C\int_{0}^{t}(t-r)^{2H-1}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|X_{s}^{i,N}|^{\overline{p}}\bigg){\rm d}r,$
+then, Lemma 2.8 implies
+$\mathbb{E}\bigg(\sup_{0\leq s\leq t}|X_{s}^{i,N}|^{\overline{p}}\bigg)\leq CE_{2H,1}((\Gamma(2H))^{1/2H}t).$
+The Mittag-Leffler function, defined by the power series $E_{2H,1}((\Gamma(2H))^{1/2H}t)$, is an entire function.
+It is a well-established result that this series converges uniformly on any compact subset of the complex plane, and therefore on any finite interval $[0,T]$ (see, e.g.,[34]), which means $\mathbb{E}\bigg(\sup_{t\in[-\tau,T]}|X_{t}^{i,N}|^{\overline{p}}\bigg)$ is bounded.
+Finally, the boundness of $\mathbb{E}\bigg(\sup_{t\in[-\tau,T]}|X_{t}^{i}|^{\overline{p}}\bigg)$ can be shown similarly.
+This completes the proof.
+∎
 Lemma 4.3.
-Let Assumption 2.4 hold. For $\overline{p}>1/H$ and $\overline{p}>2p$, we have
+Let Assumptions 2.2-2.5 hold. Fix any $\overline{p}>4$ in Lemma 4.2. Then, for $p\in[2,\frac{\overline{p}}{2})$, we have
 $\mathbb{E}\bigg(\sup_{t\in[0,T]}|X_{t}^{i}-X_{t}^{i,N}|^{p}\bigg)\leq C\times\left\{\begin{array}[]{ll}N^{-\frac{1}{2}},\ \ p>\frac{d}{2},\\ N^{-\frac{1}{2}}\log(1+N),\ p=\frac{d}{2},\\ N^{-\frac{p}{d}},\ p\in[2,\frac{d}{2}),\end{array}\right$.
-where $C$ is a positive constant depending on $p,T,H,L,q$ but independent of $N$.
+where $C$ is a positive constant depending on $p,T,H,L$ but independent of $N$ and $d$.
 Proof.
 For $i=1,2,\cdots,N$, by (4.1) and (4.2), set
-$\Theta_{t}^{i}=X_{t}^{i}-D(X_{t-\tau}^{i})-X_{t}^{i,N}+D(X_{t-\tau}^{i,N}).$
-By Itô formula, we get
-$|\Theta_{t}^{i}|^{p}-|\Theta_{0}^{i}|^{p}$
-$\leq$ $p\int_{0}^{t}|\Theta_{s}^{i}|^{p-1}\bigg{[}b(X_{s}^{i},X_{s-\tau}^{i},\mathcal{L}_{X_{s}^{i}})-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})\bigg{]}{\rm d}s$
-$+p\int_{0}^{t}|\Theta_{t}^{i}|^{p-1}\bigg{[}\sigma(\mathcal{L}_{X_{s}^{i}})-\sigma(\mu_{s}^{X,N})\bigg{]}{\rm d}B_{s}^{H,i}$
-$+p(p-1)\int_{0}^{t}|\Theta_{s}^{i}|^{p-2}\bigg{(}\sigma(\mathcal{L}_{X_{s}^{i}})-\sigma(\mu_{s}^{X,N})\bigg{)}\bigg{[}\int_{0}^{s}\phi(s,r)(\sigma(\mathcal{L}_{X_{r}^{i}})-\sigma(\mu_{r}^{X,N})){\rm d}r\bigg{]}{\rm d}s$
-=:absent:=:= : $Q_{1}(t)+Q_{2}(t)+Q_{3}(t).$
-For any $t\in[0,T]$, by Assumption 2.4, the Young’s inequality, we get
-$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}Q_{1}(s)\bigg{)}$
-$=$ $p\int_{0}^{t}|\Theta_{s}^{i}|^{p-1}\bigg{[}b(X_{s}^{i},X_{s-\tau}^{i},\mathcal{L}_{X_{s}^{i}})-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})\bigg{]}{\rm d}s$
-$\leq$ $p\mathbb{E}\int_{0}^{t}|\Theta_{s}^{i}|^{p-2}\langle\Theta_{s}^{i},b(X_{s}^{i},X_{s-\tau}^{i},\mathcal{L}_{X_{s}^{i}})-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})\rangle{\rm d}s$
-$\leq$ $C\int_{0}^{t}|\Theta_{s}^{i}|^{p}{\rm d}s+C\int_{0}^{t}\mathbb{E}[|X_{s}^{i}-X_{s}^{i,N}|^{p}+|X_{s-\tau}^{i}-X_{s-\tau}^{i,N}|^{p}+(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N}))^{p}]{\rm d}s.$
-In the same way in (3.3), by Assumption 2.5, the Young’s inequality and the Hölder inequality, we get
-$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}Q_{2}(s)\bigg{)}\leq$ $C\mathbb{E}\int_{0}^{t}|\Theta_{s}^{i}|^{\overline{p}-1}|\sigma(\mathcal{L}_{X_{s}^{i}})-\sigma(\mu_{s}^{X,N})|{\rm d}s$
-$\leq$ $C\int_{0}^{t}|\Theta_{s}^{i}|^{p}{\rm d}s+C\int_{0}^{t}\mathbb{E}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N}))^{p}{\rm d}s.$
-By Assumption 2.5, Remark 2.1, the Young’s inequality, we get
-$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}Q_{3}(s)\bigg{)}\leq$ $C\mathbb{E}\int_{0}^{t}|\Theta_{s}^{i}|^{p}{\rm d}s+2CHT^{2H-1}\mathbb{E}\int_{0}^{t}\lvert\sigma\big{(}\mathcal{L}_{X_{u}^{i}}\big{)}-\sigma\big{(}\mu_{u}^{X,N}\big{)}\rvert^{p}{\rm d}s$
-$\leq$ $C\mathbb{E}\int_{0}^{t}[(|\Theta_{s}^{i}|^{p}+\big{(}\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N})\big{)}^{p}]{\rm d}s.$
+$\Upsilon_{t}^{i}=X_{t}^{i}-D(X_{t-\tau}^{i})-X_{t}^{i,N}+D(X_{t-\tau}^{i,N}).$
+By using the fractional Itô formula to $|\Upsilon_{t}^{i}|^{p}$, we can show
+$|\Upsilon_{t}^{i}|^{p}-|\Upsilon_{0}^{i}|^{p}$
+$\leq$ $p\int_{0}^{t}|\Upsilon_{s}^{i}|^{p-2}\langle\Upsilon_{s}^{i},b(X_{s}^{i},X_{s-\tau}^{i},\mathcal{L}_{X_{s}^{i}})-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})\rangle{\rm d}s$
+$+p\int_{0}^{t}|\Upsilon_{s}^{i}|^{p-2}\langle\Upsilon_{s}^{i},\sigma(\mathcal{L}_{X_{s}^{i}})-\sigma(\mu_{s}^{X,N})\rangle{\rm d}B_{s}^{H,i}$
+$+p(p-1)H(2H-1)\int_{0}^{t}|\Upsilon_{s}^{i}|^{p-2}\bigg(\int_{0}^{s}(s-r)^{2H-2}\|\sigma(\mathcal{L}_{X_{r}^{i}})-\sigma(\mu_{r}^{X,N})\|^{2}{\rm d}r\bigg){\rm d}s$
+$=:$ $H_{1}(t)+H_{2}(t)+H_{3}(t).$
+Assumptions 2.3, 2.4 and the Young’s inequality lead to
+$\mathbb{E}\bigg(\sup_{0\leq s\leq t}H_{1}(s)\bigg)=p\mathbb{E}\int_{0}^{t}|\Upsilon_{s}^{i}|^{p-2}\langle\Upsilon_{s}^{i},b(X_{s}^{i},X_{s-\tau}^{i},\mathcal{L}_{X_{s}^{i}})-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})\rangle{\rm d}s$
+$\leq$ $C\mathbb{E}\int_{0}^{t}|X_{t}^{i}-X_{t}^{i,N}|^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}|X_{s-\tau}^{i}-X_{s-\tau}^{i,N}|^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}\mathbb{W}_{2}^{p}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N}){\rm d}s.$
+As the same way of achieving (3.6), we use Assumption 2.5, the Young’s inequality and the Hölder inequality to get
+$\mathbb{E}\bigg(\sup_{0\leq s\leq t}H_{2}(s)\bigg)=$ $p\mathbb{E}\int_{0}^{t}|\Upsilon_{s}^{i}|^{p-2}\langle\Upsilon_{s}^{i},\sigma(\mathcal{L}_{X_{s}^{i}})-\sigma(\mu_{s}^{X,N})\rangle{\rm d}B_{s}^{H,i}$
+$\leq$ $C\epsilon\mathbb{E}\left(\sup_{0\leq s\leq t}|\Upsilon_{s}^{i}|^{{p}}\right)+C\mathbb{E}\int_{0}^{t}\mathbb{W}_{2}^{p}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N}){\rm d}s.$
+Further, Assumptions 2.3, 2.5, Remark 2.1, the Young’s inequality and Fubini theorem give
+$\mathbb{E}\bigg(\sup_{0\leq s\leq t}H_{3}(s)\bigg)$
+$=$ $p(p-1)H(2H-1)\int_{0}^{t}|\Upsilon_{s}^{i}|^{p-2}\bigg(\int_{0}^{s}(s-r)^{2H-2}\|\sigma(\mathcal{L}_{X_{r}^{i}})-\sigma(\mu_{r}^{X,N})\|^{2}{\rm d}r\bigg){\rm d}s$
+$\leq$ $C\mathbb{E}\int_{0}^{t}|\Upsilon_{s}^{i}|^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}(t-s)^{2H-1}\mathbb{W}_{2}^{p}\big(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N}\big){\rm d}s$
+$\leq$ $C\mathbb{E}\int_{0}^{t}|X_{t}^{i}-X_{t}^{i,N}|^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}|X_{s-\tau}^{i}-X_{s-\tau}^{i,N}|^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}(t-s)^{2H-1}\mathbb{W}_{2}^{p}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N}){\rm d}s.$
 Noting that
-$\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N})\leq\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X})+\mathbb{W}_{q}(\mu_{s}^{X},\mu_{s}^{X,N}),$
-For $q<p$, by Lemma 4.1, we get
-$\mathbb{E}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N}))^{p}=$ $\mathbb{E}[(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N}))^{q}]^{p/{q}}$
-$\leq$ $\mathbb{E}\bigg{(}2^{q-1}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X}))^{q}+2^{q-1}(\mathbb{W}_{q}(\mu_{s}^{X},\mu_{s}^{X,N}))^{q}\bigg{)}^{p/q}$
-$\leq$ $C\mathbb{E}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X}))^{p}+C\mathbb{E}(\mathbb{W}_{q}(\mu_{s}^{X},\mu_{s}^{X,N}))^{p}$
-$\leq$ $C\mathbb{E}\bigg{(}\frac{1}{N}\sum_{j=1}^{N}|X_{s}^{j}-X_{s}^{j,N}|^{q}\bigg{)}^{p/q}+C\mathbb{E}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X}))^{p}$
-$\leq$ $C\mathbb{E}|X_{s}^{i}-X_{s}^{i,N}|^{p}+C\mathbb{E}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X}))^{p}.$
-Applying the Gronwall inequality, we get
-$\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|\Theta_{s}^{i}|^{p}\bigg{)}$
-$\leq$ $C\int_{0}^{t}\mathbb{E}|X_{s}^{i}-X_{s}^{i,N}|^{p}{\rm d}s+C\int_{0}^{t}\mathbb{E}|X_{s-\tau}^{i}-X_{s-\tau}^{i,N}|^{p}{\rm d}s+C\int_{0}^{t}\mathbb{E}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N}))^{p}{\rm d}s$
-$\leq$ $\mathbb{E}\int_{0}^{t}\bigg{(}\sup_{s\in[0,t]}|X_{s}^{i}-X_{s}^{i,N}|^{p}\bigg{)}{\rm d}s+C\int_{0}^{t}\mathbb{E}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N}))^{p}{\rm d}s$
-$\leq$ $\mathbb{E}\int_{0}^{t}\bigg{(}\sup_{s\in[0,t]}|X_{s}^{i}-X_{s}^{i,N}|^{p}\bigg{)}{\rm d}s+C\int_{0}^{t}\mathbb{E}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X}))^{p}{\rm d}s.$
-According to (2.7), we obtain
-$|X_{s}^{i}-X_{s}^{i,N}|^{p}\leq\lambda|X_{t-\tau}^{i}-X_{t-\tau}^{i,N}|^{p}+\frac{1}{(1-\lambda)^{p-1}}|\Theta_{s}^{i}|^{p}.$
-Noting that $10<\lambda<$ 1, by the Gronwall inequality, we get
-$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}|X_{s}^{i}-X_{s}^{i,N}|^{p}\bigg{)}\leq$ $C\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}|\Theta_{s}^{i}|^{p}\bigg{)}\leq C\int_{0}^{t}\mathbb{E}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X}))^{p}{\rm d}s.$
-Since $(\mathbb{W}_{p}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X}))^{p}$ is controlled by the Wasserstein distance estimate([13], Theorem 1), for $\overline{p}>2p$, we get
-$\mathbb{E}(\mathbb{W}_{p}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X}))^{p}\leq$ $C(\mathbb{E}|X_{s}^{i}|^{\overline{p}})^{p/{\overline{p}}}\times\left\{\begin{array}[]{ll}N^{-\frac{1}{2}},\ \ p>\frac{d}{2},\\ N^{-\frac{1}{2}}\log(1+N),\ p=\frac{d}{2},\\ N^{-\frac{p}{d}},\ p\in[2,\frac{d}{2}).\end{array}\right.$
-Then, by Lemma 4.2, we complete the proof. ∎
-5 The tamed theta scheme
-In this section, we consider the tamed theta scheme to approximate solution of (4.2). Since the drift coefficient exhibits super-linear growth, we define
+$\mathbb{W}_{2}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N})\leq\mathbb{W}_{2}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X})+\mathbb{W}_{2}(\mu_{s}^{X},\mu_{s}^{X,N}),$
+then, for $p\geq 2$, Lemma 4.1 leads to
+$\mathbb{E}\mathbb{W}_{2}^{p}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N})=$ $\mathbb{E}[\mathbb{W}_{2}^{2}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N})]^{p/{2}}$
+$\leq$ $\mathbb{E}\bigg(2\mathbb{W}_{2}^{2}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X})+2\mathbb{W}_{2}^{2}(\mu_{s}^{X},\mu_{s}^{X,N})\bigg)^{p/2}$
+$\leq$ $C\mathbb{E}\mathbb{W}_{2}^{p}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X})+C\mathbb{E}\mathbb{W}_{2}^{p}(\mu_{s}^{X},\mu_{s}^{X,N})$
+$\leq$ $C\mathbb{E}\bigg(\frac{1}{N}\sum_{i=1}^{N}|X_{s}^{i}-X_{s}^{i,N}|^{2}\bigg)^{p/2}+C\mathbb{E}\mathbb{W}_{2}^{p}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X})$
+$\leq$ $C\mathbb{E}|X_{s}^{i}-X_{s}^{i,N}|^{p}+C\mathbb{E}\mathbb{W}_{2}^{p}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X}).$
+Putting $H_{1}(t)$-$H_{3}(t)$ together, we conclude that
+$\mathbb{E}\bigg(\sup_{0\leq s\leq t}|\Upsilon_{s}^{i}|^{p}\bigg)\leq$ $C\mathbb{E}\int_{0}^{t}|X_{s}^{i}-X_{s}^{i,N}|^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}|X_{s-\tau}^{i}-X_{s-\tau}^{i,N}|^{p}{\rm d}s$
+$+C\mathbb{E}\int_{0}^{t}\mathbb{W}_{2}^{p}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N}){\rm d}s+C\mathbb{E}\int_{0}^{t}(t-s)^{2H-1}\mathbb{W}_{2}^{p}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N}){\rm d}s$
+$\leq$ $\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|X_{s}^{i}-X_{s}^{i,N}|^{p}\bigg){\rm d}r+C\int_{0}^{t}(t-r)^{2H-1}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|X_{s}^{i}-X_{s}^{i,N}|^{p}\bigg){\rm d}r$
+$+C\int_{0}^{t}(t-r)^{2H-1}\mathbb{E}\mathbb{W}_{2}^{p}(\mathcal{L}_{X_{r}^{i}},\mu_{r}^{X}){\rm d}r.$
+According to Lemma 2.7, it derives
+$|X_{s}^{i}-X_{s}^{i,N}|^{p}\leq\lambda|X_{s-\tau}^{i}-X_{s-\tau}^{i,N}|^{p}+\frac{1}{(1-\lambda)^{p-1}}|\Upsilon_{s}^{i}|^{p}.$
+Noting that $0<\lambda<1$, then
+$\mathbb{E}\bigg(\sup_{0\leq s\leq t}|X_{s}^{i}-X_{s}^{i,N}|^{p}\bigg)\leq$ $C\mathbb{E}\bigg(\sup_{0\leq s\leq t}|\Upsilon_{s}^{i}|^{p}\bigg).$
+That is,
+$\mathbb{E}\bigg(\sup_{0\leq s\leq t}|X_{s}^{i}-X_{s}^{i,N}|^{p}\bigg)$
+$\leq$ $\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|X_{s}^{i}-X_{s}^{i,N}|^{p}\bigg){\rm d}r+C\int_{0}^{t}(t-r)^{2H-1}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|X_{s}^{i}-X_{s}^{i,N}|^{p}\bigg){\rm d}r$
+$+C\int_{0}^{t}(t-r)^{2H-1}\mathbb{E}\mathbb{W}_{2}^{p}(\mathcal{L}_{X_{r}^{i}},\mu_{r}^{X}){\rm d}r.$
+Since $\mathbb{W}_{2}^{p}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X})$ is controlled by the Wasserstein distance estimate([35], Theorem 1), for $2\leq p<\overline{p}/2$, it follows
+$\mathbb{E}\mathbb{W}_{2}^{p}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X})\leq$ $C(\mathbb{E}|X_{s}^{i}|^{\overline{p}})^{p/{\overline{p}}}\times\left\{\begin{array}[]{ll}N^{-\frac{1}{2}},\ \ p>\frac{d}{2},\\ N^{-\frac{1}{2}}\log(1+N),\ p=\frac{d}{2},\\ N^{-\frac{p}{d}},\ p\in[2,\frac{d}{2}).\end{array}\right$.
+Then, similar to (4.4), by Lemmas 2.8 and 4.2, the desired result can be obtained. ∎
+5 The Tamed Theta EM Scheme
+In this section, we consider the tamed theta EM scheme to approximate solution of (4.2). Since the drift coefficient exhibits super-linear growth, we define
 (5.1) $b_{\Delta}(x,y,\mu)=\frac{b(x,y,\mu)}{1+\Delta^{\alpha}|b(x,y,\mu)|},$
-for $x,y\in\mathbb{R}^{d}$, $\mu\in\mathcal{P}_{q}(\mathbb{R}^{d})$ and $\alpha\in(0,1/2]$.
+for $x,y\in\mathbb{R}^{d}$, $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and $\alpha\in(0,1/2]$.
 Assume that for any time $T>0$, there exist positive constants $M,m\in\mathbb{N}$ such that $\Delta=\frac{\tau}{m}=\frac{T}{M}$, where $\Delta\in(0,1)$ is the step-size.
-Consequently, for the time grid $t_{k}=t_{0}+k\Delta$, it holds that $t_{k+1}-\tau=t_{k+1-m}$.
-For $k=-m,\cdots,0,$ we set $Y_{t_{k}}^{i,N}=\xi^{i}(t_{k})$ and $t_{0}=0$.
-For $k=0,1,\cdots,M$, we now define the discrete EM scheme for (4.2) as follows,
+Consequently, for the time grid $t_{k}=k\Delta$, it holds that $t_{k+1}-\tau=t_{k+1-m}$.
+For $k=-m,\cdots,0,$ we set $Y_{t_{k}}^{i,N}=\xi^{i}(t_{k})$.
+For $k=0,1,\cdots,M$, we now define the discrete tamed theta EM scheme for (4.2) as follows,
 (5.2) $Y_{t_{k+1}}^{i,N}-D(Y_{t_{k+1-m}}^{i,N})=$ $Y_{t_{k}}^{i,N}-D(Y_{t_{k-m}}^{i,N})+\theta b_{\Delta}(Y_{t_{k+1}}^{i,N},Y_{t_{k+1-m}}^{i,N},\mu_{t_{k+1}}^{Y,N})\Delta$
 $+(1-\theta)b_{\Delta}(Y_{t_{k}}^{i,N},Y_{t_{k-m}}^{i,N},\mu_{t_{k}}^{Y,N})\Delta+\sigma(\mu_{t_{k}}^{Y,N})\Delta B_{t_{k}}^{H,i},$
-where the empirical measures $\mu_{t_{k}}^{Y,N}(\cdot):=\frac{1}{N}\sum\limits_{j=1}^{N}\delta_{Y_{t_{k}}^{j,N}}(\cdot)$ and $\Delta B_{t_{k}}^{H,i}=B_{t_{k+1}}^{H,i}-B_{t_{k}}^{H,i}$. Let
+where the empirical measures $\mu_{(\cdot)}^{Y,N}:=\frac{1}{N}\sum\limits_{i=1}^{N}\delta_{Y_{(\cdot)}^{i,N}}$ and $\Delta B_{t_{k}}^{H,i}=B_{t_{k+1}}^{H,i}-B_{t_{k}}^{H,i}$. Let
 (5.5) $\overline{Y}_{\Delta}^{i,N}(t)=\left\{\begin{array}[]{ll}\xi^{i}(t),\ \ -\tau\leq t\leq 0,\\ \sum\limits_{k=0}^{[T/\Delta]}Y_{t_{k}}^{i,N}{\rm I}_{[k\Delta,(k+1)\Delta)}(t),\ \ t\geq 0,\end{array}\right$.
@@
 $\overline{Y}_{\Delta+}^{i,N}(t)=\left\{\begin{array}[]{ll}\xi^{i}(t),\ \ -\tau\leq t\leq 0,\\ \sum\limits_{k=0}^{[T/\Delta]}Y_{t_{k+1}}^{i,N}{\rm I}_{[k\Delta,(k+1)\Delta)}(t),\ \ t\geq 0,\end{array}\right$.
-and the continuous EM solution writes as
+and the continuous tamed theta EM solution writes as
 (5.6) $Y_{\Delta}^{i,N}(t)=$ $D(\overline{Y}_{\Delta}^{i,N}(t))+\xi^{i}(0)-D(\xi^{i}(-\tau))+\theta\int_{0}^{t}b_{\Delta}(\overline{Y}_{\Delta+}^{i,N}(s),\overline{Y}_{\Delta+}^{i,N}(s-\tau),\overline{\mu}^{Y_{+},N}(s)){\rm d}s$
 $+(1-\theta)\int_{0}^{t}b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s)){\rm d}s+\int_{0}^{t}\sigma(\overline{\mu}^{Y,N}(s)){\rm d}B_{s}^{H,i},\ \ t>0,$
-where $\overline{\mu}_{s}^{Y,N}:=\frac{1}{N}\sum\limits_{j=1}^{N}\delta_{\overline{Y}_{t}^{j,N}}(\cdot)$, and $Y_{\Delta}^{i,N}(t)=\xi^{i}(t)$ when $t_{k}\in[-\tau,0]$.
+where $\overline{\mu}^{Y,N}(\cdot):=\frac{1}{N}\sum\limits_{i=1}^{N}\delta_{\overline{Y}_{\Delta}^{i,N}}(\cdot)$, $\overline{\mu}^{Y_{+},N}(\cdot):=\frac{1}{N}\sum\limits_{i=1}^{N}\delta_{\overline{Y}_{\Delta+}^{i,N}}(\cdot)$ and $Y_{\Delta}^{i,N}(t)=\xi^{i}(t)$ for $t\in[-\tau,0]$.
 However, this $Y_{\Delta}^{i,N}(t)$ is not $\mathcal{F}_{t}$-adapted, it does not meet the fundamental requirement in the Itô stochastic analysis.
-To avoid calculus, we use the discrete split-step theta scheme (see [28]) described as follows: For $k=-m,\cdots,-1$, set $z_{t_{k}}^{i,N}=\xi^{i}(k\Delta)$.
-For $k=0,1,\cdots,M-1$, we rewrite (5.2) as
-(5.9) $\left\{\begin{array}[]{ll}Y_{t_{k}}^{i,N}=D(Y_{t_{k-m}}^{i,N})+z_{t_{k}}^{i,N}-D(z_{t_{k-m}})+\theta b_{\Delta}(Y_{t_{k}}^{i,N},Y_{t_{k-m}}^{i,N},\mu_{t_{k}}^{Y,N})\Delta,\\ z_{t_{k+1}}^{i,N}=D(z_{t_{k+1-m}}^{i,N})+z_{t_{k}}^{i,N}-D(z_{t_{k-m}}^{i,N})+b_{\Delta}(Y_{t_{k}}^{i,N},Y_{t_{k-m}}^{i,N},\mu_{t_{k}}^{Y,N})\Delta+\sigma(\mu_{t_{k}}^{Y,N})\Delta B_{t_{k}}^{H,i},\end{array}\right.$
+To avoid calculus, we use the discrete split-step theta scheme (see [26]) described as follows: For $k=-m,\cdots,0$, set $Z_{t_{k}}^{i,N}=\xi^{i}(k\Delta)$.
+For $k=0,1,\cdots,M-1$, rewrite (5.2) as
+(5.9) $\left\{\begin{array}[]{ll}Y_{t_{k}}^{i,N}=D(Y_{t_{k-m}}^{i,N})+Z_{t_{k}}^{i,N}-D(Z_{t_{k-m}}^{i,N})+\theta b_{\Delta}(Y_{t_{k}}^{i,N},Y_{t_{k-m}}^{i,N},\mu_{t_{k}}^{Y,N})\Delta,\\ Z_{t_{k+1}}^{i,N}=D(Z_{t_{k+1-m}}^{i,N})+Z_{t_{k}}^{i,N}-D(Z_{t_{k-m}}^{i,N})+b_{\Delta}(Y_{t_{k}}^{i,N},Y_{t_{k-m}}^{i,N},\mu_{t_{k}}^{Y,N})\Delta+\sigma(\mu_{t_{k}}^{Y,N})\Delta B_{t_{k}}^{H,i},\end{array}\right$.
 This scheme can also be rewritten as
-$z_{t_{k+1}}^{i,N}-D(z_{t_{k+1-m}}^{i,N})=z_{t_{0}}^{i,N}-D(z_{t_{-m}}^{i,N})+\sum_{j=0}^{k}b_{\Delta}(Y_{t_{j}}^{i,N},Y_{t_{j-m}}^{i,N},\mu_{t_{j}}^{Y,N})\Delta+\sum_{j=0}^{k}\sigma(\mu_{t_{j}}^{Y,N})\Delta B_{t_{j}}^{H,i}$
-$=$ $\xi^{i}(0)-D(\xi^{i}(-\tau))-\theta b_{\Delta}(\xi^{i}(0),\xi^{i}(-\tau),\mu_{t_{0}}^{i,N})\Delta+\sum_{j=0}^{k}b_{\Delta}(Y_{t_{j}}^{i,N},Y_{t_{j-m}}^{i,N},\mu_{t_{j}}^{Y,N})\Delta$
-$+\sum_{j=0}^{k}\sigma(\mu_{t_{j}}^{Y,N})\Delta B_{t_{j}}^{H,i}.$
-Similarly, we define the corresponding continuous-time split-step tamed theta solution $Z_{\Delta}^{i,N}(t)$ as follows: For any $t\in[-\tau,0]$, $Z_{\Delta}^{i,N}(t)=\xi^{i}(t)$, $Z_{\Delta}^{i,N}(0)=\xi^{i}(0)-\theta b_{\Delta}(\xi^{i}(0)$, $\xi^{i}(-\tau),\mu_{0}^{Y,N})\Delta$.
+$Z_{t_{k+1}}^{i,N}-D(Z_{t_{k+1-m}}^{i,N})$
+$=$ $Z_{t_{0}}^{i,N}-D(Z_{t_{-m}}^{i,N})+\sum_{j=0}^{k}b_{\Delta}(Y_{t_{j}}^{i,N},Y_{t_{j-m}}^{i,N},\mu_{t_{j}}^{Y,N})\Delta+\sum_{j=0}^{k}\sigma(\mu_{t_{j}}^{Y,N})\Delta B_{t_{j}}^{H,i}$
+$=$ $\xi^{i}(0)-D(\xi^{i}(-\tau))-\theta b_{\Delta}(\xi^{i}(0),\xi^{i}(-\tau),\delta_{\xi(0)})\Delta$
+$+\sum_{j=0}^{k}b_{\Delta}(Y_{t_{j}}^{i,N},Y_{t_{j-m}}^{i,N},\mu_{t_{j}}^{Y,N})\Delta+\sum_{j=0}^{k}\sigma(\mu_{t_{j}}^{Y,N})\Delta B_{t_{j}}^{H,i}.$
+Similarly, we define the corresponding continuous-time split-step tamed theta EM solution $Z_{\Delta}^{i,N}(t)$ as follows: For any $t\in[-\tau,0)$, $Z_{\Delta}^{i,N}(t)=\xi^{i}(t)$, $Z_{\Delta}^{i,N}(0)=\xi^{i}(0)-\theta b_{\Delta}(\xi^{i}(0)$, $\xi^{i}(-\tau),\delta_{\xi(0)})\Delta$.
 For any $t\in[0,T]$,
 (5.10) ${\rm d}[Z_{\Delta}^{i,N}(t)-D(Z_{\Delta}^{i,N}(t-\tau))]=b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(t),\overline{Y}_{\Delta}^{i,N}(t-\tau),\overline{\mu}^{Y,N}(t)){\rm d}t+\sigma(\overline{\mu}^{Y,N}(t)){\rm d}B_{t}^{H,i},$
-where $\overline{Y}_{\Delta}^{i,N}(t)$ is defined by (5.5). Combining (5.6), (5.9) and (5.10), we have
+where $\overline{Y}_{\Delta}^{i,N}(t)$ is defined by (5.5). Combining (5.6), (5.9) and (5.10), we see
 $Y_{\Delta}^{i,N}(t)-D(Y_{\Delta}^{i,N}(t-\tau))-\theta b_{\Delta}(Y_{\Delta}^{i,N}(t),Y_{\Delta}^{i,N}(t-\tau),\mu^{Y,N}(t))\Delta=Z_{\Delta}^{i,N}(t)-D(Z_{\Delta}^{i,N}(t-\tau)).$
@@
 (5.11) $\widetilde{Y}_{\Delta}^{i,N}(t)=\widetilde{Y}_{\Delta}^{i,N}(0)+\int_{0}^{t}b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s)){\rm d}s+\int_{0}^{t}\sigma(\overline{\mu}^{Y,N}(s)){\rm d}B_{s}^{H,i}.$
-It is clear that $\widetilde{Y}_{\Delta}^{i,N}$ coincide with $\overline{Y}_{\Delta}^{i,N}(t)-D(\overline{Y}_{\Delta}^{i,N}(t-\tau))-\theta b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(t),\overline{Y}_{\Delta}^{i,N}(t-\tau),\overline{\mu}^{Y,N}(t))$ at grid points $t=k\Delta$, $k=0,1,\cdots,M-1$, this also means that the continuous-time tamed theta solution $Y_{\Delta}^{i,N}(t)$ and the discrete-time tamed theta solution $\overline{Y}_{\Delta}^{i,N}(t)$ take the same values at grid points $t=k\Delta$, $k=0,1,\cdots,M-1$.
+It is clear that $\widetilde{Y}_{\Delta}^{i,N}$ coincide with $\overline{Y}_{\Delta}^{i,N}(t)-D(\overline{Y}_{\Delta}^{i,N}(t-\tau))-\theta b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(t),\overline{Y}_{\Delta}^{i,N}(t-\tau),\overline{\mu}^{Y,N}(t))$ at grid points $t=k\Delta$, $k=0,1,\cdots,M-1$, this also means that the continuous-time tamed theta EM solution $Y_{\Delta}^{i,N}(t)$ and the discrete-time tamed theta EM solution $\overline{Y}_{\Delta}^{i,N}(t)$ take the same values at grid points $t=k\Delta$, $k=0,1,\cdots,M-1$.
 Lemma 5.1.
-There exists a positive constant $K_{7}\geq 1$ such that
-$|b_{\Delta}(x,y,\mu)|\leq\min\{K_{7}\Delta^{-\alpha}(1+|x|+|y|+\mathbb{W}_{q}(\mu,\delta_{0})),|b(x,y,\mu)|\}.$
+There exists a positive constant $K_{6}\geq 1$ such that
+$|b_{\Delta}(x,y,\mu)|\leq\min\{K_{6}\Delta^{-\alpha}(1+|x|+|y|+\mathbb{W}_{2}(\mu,\delta_{0})),|b(x,y,\mu)|\}$
+for $x,y\in\mathbb{R}^{d}$ and $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$.
 Proof.
 For $\forall x,y>0$, by (5.1), it is easy to see that $|b_{\Delta}(x,y,\mu)|\leq|b(x,y,\mu)|$. On the other hand, we immediately get
-$|b_{\Delta}(x,y,\mu)|=\Delta^{-\alpha}\frac{|b(x,y,\mu)|}{\Delta^{-\alpha}+|b(x,y,\mu)|}\leq\Delta^{-\alpha}\leq K_{7}\Delta^{-\alpha}(1+|x|+|y|+\mathbb{W}_{q}(\mu,\delta_{0})).$
+$|b_{\Delta}(x,y,\mu)|=\Delta^{-\alpha}\frac{|b(x,y,\mu)|}{\Delta^{-\alpha}+|b(x,y,\mu)|}\leq\Delta^{-\alpha}\leq K_{6}\Delta^{-\alpha}(1+|x|+|y|+\mathbb{W}_{2}(\mu,\delta_{0})).$
 Lemma 5.2.
 Let Assumption 2.4 hold. One observes
-$\langle x-D(y)-\overline{x}+D(\overline{y}),b_{\Delta}(x,y,\mu)-b_{\Delta}(\overline{x},\overline{y},\nu)\rangle\leq K_{2}(|x-\overline{x}|^{2}+|y-\overline{y}|^{2}+(\mathbb{W}_{q}(\mu,\nu))^{2}),$
+$\langle x-D(y)-\overline{x}+D(\overline{y}),b_{\Delta}(x,y,\mu)-b_{\Delta}(\overline{x},\overline{y},\nu)\rangle\leq K_{7}(|x-\overline{x}|^{2}+|y-\overline{y}|^{2}+\mathbb{W}_{2}^{2}(\mu,\nu)),$
 and
-$|b(x,y,\mu)-b_{\Delta}(x,y,\mu)|^{p}\leq K_{8}\Delta^{\alpha p}(1+|x|^{2(l+1)p}+|y|^{2(l+1)p}+(\mathbb{W}_{q}(\mu,\delta_{0}))^{2p}),$
-where $K_{8}$ is a positive constant.
+$|b(x,y,\mu)-b_{\Delta}(x,y,\mu)|^{p}\leq K_{8}\Delta^{\alpha p}(1+|x|^{2(l+1)p}+|y|^{2(l+1)p}+\mathbb{W}_{2}^{2p}(\mu,\delta_{0})),$
+where $K_{7},K_{8}$ are positive constants.
 Proof.
-We divide it into two case. Case 1, for $b(x,y,\mu)\cdot b(\overline{x},\overline{y},\nu)<0$, we get
-$\langle x-D(y)-\overline{x}+D(\overline{y}),b_{\Delta}(x,y,\mu)-b_{\Delta}(\overline{x},\overline{y},\nu)\rangle$
-$=$ $\bigg{\langle}x-D(y)-\overline{x}+D(\overline{y}),\frac{b(x,y,\mu)}{1+\Delta^{\alpha}|b(x,y,\mu)|}-\frac{b(\overline{x},\overline{y},\nu)}{1+\Delta^{\alpha}|b(\overline{x},\overline{y},\nu)|}\bigg{\rangle}$
-$\leq$ $K_{2}(|x-\overline{x}|^{2}+|y-\overline{y}|^{2}+(\mathbb{W}_{q}(\mu,\nu))^{2}).$
-Case 2, for $b(x,y,\mu)\cdot b(\overline{x},\overline{y},\nu)>0$, we get
-$\langle x-D(y)-\overline{x}+D(\overline{y}),b_{\Delta}(x,y,\mu)-b_{\Delta}(\overline{x},\overline{y},\nu)\rangle$
-$=$ $\bigg{\langle}x-D(y)-\overline{x}+D(\overline{y}),\frac{b(x,y,\mu)-b(\overline{x},\overline{y},\nu)}{(1+\Delta^{\alpha}|b(x,y,\mu)|)(1+\Delta^{\alpha}|b(\overline{x},\overline{y},\nu))}\bigg{\rangle}$
-$+\bigg{\langle}x-D(y)-\overline{x}+D(\overline{y}),\frac{\Delta^{\alpha}[b(x,y,\mu)\cdot b(\overline{x},\overline{y},\nu)-b(x,y,\mu)\cdot b(\overline{x},\overline{y},\nu)]}{(1+\Delta^{\alpha}|b(x,y,\mu)|)(1+\Delta^{\alpha}|b(\overline{x},\overline{y},\nu))}\bigg{\rangle}$
-$\leq$ $K_{2}(|x-\overline{x}|^{2}+|y-\overline{y}|^{2}+(\mathbb{W}_{q}(\mu,\nu))^{2}).$
-Furthermore,
-$|b(x,y,\mu)-b_{\Delta}(x,y,\mu)|^{p}\leq\Delta^{\alpha p}|b(x,y,\mu)|^{2p},$
-the second part is obvious. ∎
+For the first inequality, we write
+$b_{\Delta}(x,y,\mu)-b_{\Delta}(\overline{x},\overline{y},\nu)$
+$=$ $\frac{b(x,y,\mu)-b(\overline{x},\overline{y},\nu)}{(1+\Delta^{\alpha}|b(x,y,\mu)|)(1+\Delta^{\alpha}|b(\overline{x},\overline{y},\nu)|)}$
+$+\frac{\Delta^{\alpha}[|b(\overline{x},\overline{y},\nu)|\cdot b(x,y,\mu)-|b(x,y,\mu)|\cdot b(\overline{x},\overline{y},\nu)]}{(1+\Delta^{\alpha}|b(x,y,\mu)|)(1+\Delta^{\alpha}|b(\overline{x},\overline{y},\nu)|)}.$
+Using Assumption 2.4 for the first term, the Lipschitz continuity of $b$ and its growth condition from Remark 2.6 along with the Young’s inequality for the second term, we obtain the desired result.
+Furthermore, for the second inequality, we observe that
+$|b(x,y,\mu)-b_{\Delta}(x,y,\mu)|^{p}$ $=\left|b(x,y,\mu)-\frac{b(x,y,\mu)}{1+\Delta^{\alpha}|b(x,y,\mu)|}\right|^{p}$
+$=\left|\frac{\Delta^{\alpha}|b(x,y,\mu)|\cdot b(x,y,\mu)}{1+\Delta^{\alpha}|b(x,y,\mu)|}\right|^{p}$
+$\leq\Delta^{\alpha p}|b(x,y,\mu)|^{2p}.$
+Then, with the growth condition of $b$, the second part can be shown. ∎
 Remark 5.3.
-Given Lemmas 5.1, 5.2 hold. We have
-$\langle x-D(y),b_{\Delta}(x,y,\mu)\rangle\leq C(1+|x|^{2}+|y|^{2}+(\mathbb{W}_{q}(\mu,\delta_{0}))^{2}).$
-5.1 Moment bounds
+Let Assumption 2.4 hold. We derive directly from Lemmas 5.1, 5.2 that
+$\langle x-D(y),b_{\Delta}(x,y,\mu)\rangle\leq C(1+|x|^{2}+|y|^{2}+\mathbb{W}_{2}^{2}(\mu,\delta_{0})).$
+5.1 Moment Boundness
 In order to get the desired results, we give the following Lemma.
 Lemma 5.4.
-Let Assumption 2.2-2.5 and Lemmas 5.1, 5.2 hold. For $\overline{p}>1/H$, we have
+Let Assumptions 2.2-2.5 hold. For $\overline{p}\geq 2$, we have
 $\sup_{0\leq t\leq T}\mathbb{E}|Y_{\Delta}^{i,N}(t)|^{\overline{p}}\leq C,$
@@
 Proof.
-For $a>0$, let $\lfloor a\rfloor$ be the integer part of $a$. Applying the Itô formula to $[1+|\widetilde{Y}_{\Delta}^{i,N}(t)|^{2}]^{\overline{p}/2}$, we get
+We begin by considering the case $\overline{p}\geq 4$, using moment estimation techniques for numerical solutions from [36].
+The case $2\leq\overline{p}<4$ then follows directly from the Young’s inequality.
+For $a>0$, let $\lfloor a\rfloor$ be the integer part of $a$.
+Applying the fractional Itô formula to $1+|\widetilde{Y}_{\Delta}^{i,N}(t)|^{2}]^{\overline{p}/2}$ and using the fact that $\mathbb{E}B_{s}^{H,i}=0$, it holds that
 $\mathbb{E}[1+|\widetilde{Y}_{\Delta}^{i,N}(t)|^{2}]^{\overline{p}/2}$
-$\leq$ $\mathbb{E}[1+|\widetilde{Y}_{\Delta}^{i,N}(0)|^{2}]^{\overline{p}/2}+p\mathbb{E}\int_{0}^{t}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\rangle$
-$+\overline{p}(\overline{p}-2)\mathbb{E}\int_{0}^{t}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}\sigma(\overline{\mu}^{Y,N}(s))D_{s}^{\phi}[1+|\widetilde{Y}_{\Delta}(s)|^{2}]{\rm d}s$
-$\leq$ $\mathbb{E}[1+|\widetilde{Y}_{\Delta}^{i,N}(0)|^{2}]^{\overline{p}/2}+\overline{p}(\overline{p}-2)\mathbb{E}\int_{0}^{t}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}|\sigma(\overline{\mu}^{Y,N}(s))|^{2}\int_{0}^{s}\phi(s,r){\rm d}r{\rm d}s$
+$\leq$ $\mathbb{E}[1+|\widetilde{Y}_{\Delta}^{i,N}(0)|^{2}]^{\overline{p}/2}+\overline{p}\mathbb{E}\int_{0}^{t}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\rangle$
+$+\overline{p}(\overline{p}-2)H(2H-1)\mathbb{E}\int_{0}^{t}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}\bigg(\int_{0}^{s}(s-r)^{2H-2}\|\sigma(\overline{\mu}^{Y,N}(r))\|^{2}{\rm d}r\bigg){\rm d}s$
+$\leq$ $\mathbb{E}[1+|\widetilde{Y}_{\Delta}^{i,N}(0)|^{2}]^{\overline{p}/2}+\overline{p}(\overline{p}-2)H(2H-1)\mathbb{E}\int_{0}^{t}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}$
+$\times\bigg(\int_{0}^{s}(s-r)^{2H-2}\|\sigma(\overline{\mu}^{Y,N}(r))\|^{2}{\rm d}r\bigg){\rm d}s$
 $+\overline{p}\mathbb{E}\int_{0}^{t}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}\langle\overline{Y}_{\Delta}^{i,N}(s)-D(\overline{Y}_{\Delta}^{i,N}(s-\tau)),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\rangle$
 $+\overline{p}\mathbb{E}\int_{0}^{t}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(s)-\widehat{Y}_{\Delta}^{i,N}(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\rangle{\rm d}s$
-=:absent:=:= : $\mathbb{E}[1+|\widetilde{Y}_{\Delta}^{i,N}(0)|^{2}]^{\overline{p}/2}+G_{1}(t)+G_{2}(t)+G_{3}(t).$
-where $\widehat{Y}_{\Delta}^{i,N}(s)=\overline{Y}_{\Delta}^{i,N}(s)-D(\overline{Y}_{\Delta}^{i,N}(s-\tau))-\theta b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\Delta$.
-By the properties of $\phi$ and the Young’s inequality, Lemma 5.1, we get
-$G_{1}(t)$
-$\leq$ $CHT^{2H-1}\mathbb{E}\int_{0}^{t}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}|\sigma(\overline{\mu}^{Y,N}(s))|^{2}{\rm d}s$
-$\leq$ $C\mathbb{E}\int_{0}^{t}\big{(}1+[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}}{2}}+|\overline{Y}_{\Delta}^{i,N}(s)|^{\overline{p}}\big{)}{\rm d}s$
-$\leq$ $C\mathbb{E}\int_{0}^{t}\bigg{(}1+|Y_{\Delta}^{i,N}(s)|^{\overline{p}}+|Y_{\Delta}^{i,N}(s-\tau)|^{\overline{p}}+|\theta b_{\Delta}(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))\Delta|^{\overline{p}}+|\overline{Y}_{\Delta}^{i,N}(s)|^{\overline{p}}\bigg{)}{\rm d}s$
-$\leq$ $C\mathbb{E}\int_{0}^{t}\bigg{(}1+|Y_{\Delta}^{i,N}(s)|^{\overline{p}}+|Y_{\Delta}^{i,N}(s-\tau)|^{\overline{p}}+|\overline{Y}_{\Delta}^{i,N}(s)|^{\overline{p}}\bigg{)}{\rm d}s$
+$=:$ $\mathbb{E}[1+|\widetilde{Y}_{\Delta}^{i,N}(0)|^{2}]^{\overline{p}/2}+G_{1}(t)+G_{2}(t)+G_{3}(t).$
+where
+$\widehat{Y}_{\Delta}^{i,N}(s)=\overline{Y}_{\Delta}^{i,N}(s)-D(\overline{Y}_{\Delta}^{i,N}(s-\tau))-\theta b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\Delta.$
+For $t\in[0,T]$, the Young’s inequality, the Fubini theorem and Lemma 5.1 give that
+$G_{1}(t)\leq$ $C\mathbb{E}\int_{0}^{t}\bigg(1+|Y_{\Delta}^{i,N}(s)|^{\overline{p}}+|Y_{\Delta}^{i,N}(s-\tau)|^{\overline{p}}$
+$+|\theta b_{\Delta}(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))\Delta|^{\overline{p}}\bigg){\rm d}s+C\int_{0}^{t}(t-r)^{2H-1}\mathbb{E}|\overline{Y}_{\Delta}^{i,N}(r)|^{\overline{p}}{\rm d}r$
+$\leq$ $C\mathbb{E}\int_{0}^{t}\bigg(1+|Y_{\Delta}^{i,N}(s)|^{\overline{p}}+|Y_{\Delta}^{i,N}(s-\tau)|^{\overline{p}}+|\overline{Y}_{\Delta}^{i,N}(s)|^{\overline{p}}\bigg){\rm d}s+C\int_{0}^{t}(t-s)^{2H-1}\mathbb{E}|\overline{Y}_{\Delta}^{i,N}(s)|^{\overline{p}}{\rm d}s$
 $+C\Delta^{(1-\alpha)\overline{p}}\mathbb{E}\int_{0}^{t}\bigg(1+|Y_{\Delta}^{i,N}(s)|^{\overline{p}}+|Y_{\Delta}^{i,N}(s-\tau)|^{\overline{p}}\bigg)$
-$\leq$ $C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)|^{\overline{p}}{\rm d}s.$
-By Remark 5.3 and the Young’s inequality, we get
-$G_{2}(t)\leq$ $C\mathbb{E}\int_{0}^{t}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}\bigg{(}1+|\overline{Y}_{\Delta}^{i,N}(s)|^{2}+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|^{2}+(\mathbb{W}_{q}(\mu^{\overline{Y},N}(s),\delta_{0}))^{2}\bigg{)}{\rm d}s$
+$\leq$ $C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)|^{\overline{p}}{\rm d}s+C\int_{0}^{t}(t-s)^{2H-1}\bigg(\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)|^{\overline{p}}\bigg){\rm d}s.$
+Remark 5.3 together with the Young’s inequality yield
+$G_{2}(t)\leq$ $C\mathbb{E}\int_{0}^{t}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}\bigg(1+|\overline{Y}_{\Delta}^{i,N}(s)|^{2}+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|^{2}+\mathbb{W}_{2}^{2}(\mu^{\overline{Y},N}(s),\delta_{0})\bigg){\rm d}s$
 $\leq$ $C\mathbb{E}\int_{0}^{t}\bigg([1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}}{2}}+1+|\overline{Y}_{\Delta}^{i,N}(s)|^{\overline{p}}+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|^{\overline{p}}\bigg){\rm d}s$
 $\leq$ $C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)|^{\overline{p}}{\rm d}s.$
-We may compute that
+We may compute
 $G_{3}(t)=$ $\overline{p}\mathbb{E}\int_{0}^{t}[1+|\widehat{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(s)-\widehat{Y}_{\Delta}^{i,N}(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\rangle{\rm d}s$
 $+p\mathbb{E}\int_{0}^{t}\bigg[[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}-[1+|\widehat{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}\bigg]$
-×$\times\langle\widetilde{Y}_{\Delta}^{i,N}(s)-\widehat{Y}_{\Delta}^{i,N}(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\rangle{\rm d}s$
-=:absent:=:= : $\overline{p}G_{31}(t)+\overline{p}G_{32}(t),$
-where $\widetilde{Y}_{\Delta}^{i,N}(t)-\widehat{Y}_{\Delta}^{i,N}(s)=\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u)){\rm d}u+\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}\sigma(\overline{\mu}^{Y,N}(u)){\rm d}B_{u}^{H,i}$.
-By Lemma 5.1 and the Young’s inequality, we get
+$\times\langle\widetilde{Y}_{\Delta}^{i,N}(s)-\widehat{Y}_{\Delta}^{i,N}(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\rangle{\rm d}s$
+$=:$ $\overline{p}G_{31}(t)+\overline{p}G_{32}(t),$
+where
+$\widetilde{Y}_{\Delta}^{i,N}(t)-\widehat{Y}_{\Delta}^{i,N}(s)=\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u)){\rm d}u+\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}\sigma(\overline{\mu}^{Y,N}(u)){\rm d}B_{u}^{H,i}.$
+We use Lemma 5.1 and the Young’s inequality to get
 $G_{31}(t)$
@@
 $\leq$ $\mathbb{E}\int_{0}^{t}[1+|\widehat{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(u))|{\rm d}u$
-×$\times|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))|{\rm d}s$
+$\times|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))|{\rm d}s$
 $\leq$ $\Delta\mathbb{E}\int_{0}^{t}[1+|\widehat{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))|^{2}{\rm d}s$
@@
 $\leq$ $C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)|^{\overline{p}}{\rm d}s.$
-Applying Itô formula to $[1+|\widetilde{Y}_{\Delta}(s)|^{2}]^{\frac{\overline{p}-2}{2}}$, we get
+Again, by the fractional Itô formula
 $[1+|\widetilde{Y}_{\Delta}(s)|^{2}]^{\frac{\overline{p}-2}{2}}$
@@
 $+(p-2)\int_{0}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u))\rangle{\rm d}u$
-$+(\overline{p}-2)(\overline{p}-4)\int_{0}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\sigma(\overline{\mu}^{Y,N}(u))D_{r}^{\phi}[1+|\widetilde{Y}_{\Delta}(u)|^{2}]^{\frac{\overline{p}-2}{2}}{\rm d}u$
+$+\frac{1}{2}H(2H-1)(\overline{p}-2)(\overline{p}-3)\int_{0}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(r)|^{2}]^{\frac{\overline{p}-4}{2}}\int_{0}^{r}(r-u)^{2H-2}\|\sigma(\overline{\mu}^{Y,N}(u))\|^{2}{\rm d}u{\rm d}r$
 $+(\overline{p}-2)\int_{0}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),\sigma(\overline{\mu}^{Y,N}(u))\rangle{\rm d}B_{u}^{H,i}.$
-Then,
-$[1+|\widehat{Y}_{\Delta}(s)|^{2}]^{\frac{\overline{p}-2}{2}}$
+Then, by the Young’s inequality and the Fubini theorem, we have
+$\mathbb{E}[1+|\widehat{Y}_{\Delta}(s)|^{2}]^{\frac{\overline{p}-2}{2}}$
 $\leq$ $[1+|\widehat{Y}_{\Delta}^{i,N}(0)|^{2}]^{\frac{\overline{p}-2}{2}}$
-$+(\overline{p}-2)\int_{0}^{\lfloor\frac{s}{\Delta}\rfloor\Delta}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u))\rangle{\rm d}u$
-$+(\overline{p}-2)(\overline{p}-4)\int_{0}^{\lfloor\frac{s}{\Delta}\rfloor\Delta}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\sigma(\overline{\mu}^{Y,N}(u))D_{u}^{\phi}[1+|\widetilde{Y}_{\Delta}(u)|^{2}]^{\frac{\overline{p}-2}{2}}{\rm d}u$
-$+(\overline{p}-2)\int_{0}^{\lfloor\frac{s}{\Delta}\rfloor\Delta}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),\sigma(\overline{\mu}^{Y,N}(u))\rangle{\rm d}B_{u}^{H,i}.$
+$+(\overline{p}-2)\mathbb{E}\int_{0}^{\lfloor\frac{s}{\Delta}\rfloor\Delta}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u))\rangle{\rm d}u$
+$+\frac{1}{2}H(2H-1)(\overline{p}-2)(\overline{p}-3)\bigg(\int_{0}^{\lfloor\frac{s}{\Delta}\rfloor\Delta}\frac{u^{2H-1}}{2H-1}\mathbb{E}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}}{2}}{\rm d}u$
+$+\int_{0}^{\lfloor\frac{s}{\Delta}\rfloor\Delta}\frac{(\lfloor\frac{s}{\Delta}\rfloor\Delta-u)^{2H-1}}{2H-1}\mathbb{E}|\overline{Y}_{\Delta}^{i,N}(u)|^{\overline{p}}{\rm d}u\bigg)$
+$+(\overline{p}-2)\mathbb{E}\int_{0}^{\lfloor\frac{s}{\Delta}\rfloor\Delta}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),\sigma(\overline{\mu}^{Y,N}(u))\rangle{\rm d}B_{u}^{H,i}.$
 One immediately gets
 $G_{32}(t)\leq$ $(\overline{p}-2)\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u))\rangle{\rm d}u$
-×$\times\langle\widetilde{Y}_{\Delta}^{i,N}(s)-\widehat{Y}_{\Delta}^{i,N}(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\rangle{\rm d}s$
-$+(\overline{p}-2)(\overline{p}-4)\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\sigma(\overline{\mu}^{Y,N}(u))D_{u}^{\phi}[1+|\widetilde{Y}_{\Delta}(u)|^{2}]^{\frac{\overline{p}-2}{2}}{\rm d}u$
-×$\times\langle\widetilde{Y}_{\Delta}^{i,N}(s)-\widehat{Y}_{\Delta}^{i,N}(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\rangle{\rm d}s$
+$\times\langle\widetilde{Y}_{\Delta}^{i,N}(s)-\widehat{Y}_{\Delta}^{i,N}(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))\rangle{\rm d}s$
+$+(\overline{p}-2)(\overline{p}-3)\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}\bigg(\frac{u^{2H-1}}{2H-1}\mathbb{E}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}}{2}}$
+$+\frac{(s-u)^{2H-1}-(\lfloor\frac{s}{\Delta}\rfloor\Delta-u)^{2H-1}}{2H-1}\mathbb{E}|\overline{Y}_{\Delta}^{i,N}(u)|^{\overline{p}}\bigg){\rm d}u$
+$\times\langle\widetilde{Y}_{\Delta}^{i,N}(s)-\widehat{Y}_{\Delta}^{i,N}(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))\rangle{\rm d}s$
 $+(\overline{p}-2)\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),\sigma(\overline{\mu}^{Y,N}(u))\rangle{\rm d}B_{u}^{H,i}$
-×$\times\langle\widetilde{Y}_{\Delta}^{i,N}(s)-\widehat{Y}_{\Delta}^{i,N}(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\rangle{\rm d}s$
-=:absent:=:= : $(\overline{p}-2)G_{321}(t)+(\overline{p}-2)(\overline{p}-4)G_{322}(t)+(\overline{p}-2)G_{323}(t).$
+$\times\langle\widetilde{Y}_{\Delta}^{i,N}(s)-\widehat{Y}_{\Delta}^{i,N}(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))\rangle{\rm d}s$
+$=:$ $(\overline{p}-2)G_{321}(t)+(\overline{p}-2)(\overline{p}-3)G_{322}(t)+(\overline{p}-2)G_{323}(t).$
 By Lemma 5.1, the Young’s inequality, we get
 $G_{321}(t)\leq$ $\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u))\rangle{\rm d}u$
-×$\times\bigg{\langle}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u)){\rm d}u,b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\bigg{\rangle}{\rm d}s$
+$\times\bigg\langle\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u)){\rm d}u,b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))\bigg\rangle{\rm d}s$
 $+\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u))\rangle{\rm d}u$
-×$\times\bigg{\langle}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}\sigma(\overline{\mu}^{Y,N}(u)){\rm d}B_{u}^{H,i},b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\bigg{\rangle}{\rm d}s$
-$\leq$ $\Delta\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-3}{2}}|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))|^{3}{\rm d}u{\rm d}s$
-$+C\mathbb{E}\int_{0}^{t}\bigg{[}\bigg{(}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-3}{2}}|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u),\overline{\mu}^{Y,N}(u))|{\rm d}u$
-×$\times|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))|\bigg{)}^{\frac{\overline{p}}{\overline{p}-1}}+\bigg{\lvert}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}\sigma(\overline{\mu}^{Y,N}(u)){\rm d}B_{u}^{H,i}\bigg{\rvert}^{\overline{p}}\bigg{]}{\rm d}s$
+$\times\bigg\langle\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}\sigma(\overline{\mu}^{Y,N}(u)){\rm d}B_{u}^{H,i},b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))\bigg\rangle{\rm d}s$
+$\leq$ $\Delta\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-3}{2}}|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))|^{3}{\rm d}u{\rm d}s$
+$+C\mathbb{E}\int_{0}^{t}\bigg[\bigg(\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-3}{2}}|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u))|{\rm d}u$
+$\times|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))|\bigg)^{\frac{\overline{p}}{\overline{p}-1}}+\bigg\lvert\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}\sigma(\overline{\mu}^{Y,N}(u)){\rm d}B_{u}^{H,i}\bigg\rvert^{\overline{p}}\bigg]{\rm d}s$
 $\leq$ $C\Delta^{2-3\alpha}\mathbb{E}\int_{0}^{t}(1+|Y_{\Delta}^{i,N}(s)|^{\overline{p}}+|Y_{\Delta}^{i,N}(s-\tau)|^{\overline{p}}+|\overline{Y}_{\Delta}^{i,N}(s)|^{p}+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|^{\overline{p}}){\rm d}s$
@@
 $+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|^{\overline{p}}){\rm d}s$
-$+C\mathbb{E}\int_{0}^{t}\bigg{[}\bigg{(}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-3}{2}}|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))|^{2}{\rm d}u|\bigg{)}^{\frac{\overline{p}}{\overline{p}-1}}$
+$+C\mathbb{E}\int_{0}^{t}\bigg[\bigg(\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-3}{2}}|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))|^{2}{\rm d}u|\bigg)^{\frac{\overline{p}}{\overline{p}-1}}$
 $+\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}\lvert\sigma(\overline{\mu}^{Y,N}(u))\rvert^{\overline{p}}{\rm d}u{\rm d}s$
-$\leq$ $C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(s)|^{\overline{p}}{\rm d}s.$
-Similarly, we get
-$G_{322}(t)\leq C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(s)|^{\overline{p}}{\rm d}s.$
-By Assumption 2.4, Lemma 5.1 and [32, Lemma 3.1.3], noting that $\mathbb{E}\int_{0}^{t}F_{s}{\rm d}B_{s}^{H,i}=0$, we get
-$G_{323}(t)=$ $\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),\sigma(\overline{\mu}^{Y,N}(u))\rangle{\rm d}B_{u}^{H,i}$
-×$\times\bigg{\langle}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u)){\rm d}u,b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\bigg{\rangle}{\rm d}s$
+$\leq$ $C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)|^{\overline{p}}{\rm d}s.$
+Similarly, $G_{322}(t)$ can be estimated by Lemma 2.8 as
+$G_{322}(t)\leq C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)|^{\overline{p}}{\rm d}s+C\int_{0}^{t}(t-s)^{2H-1}\bigg(\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)|^{\overline{p}}\bigg){\rm d}s.$
+Using Assumption 2.4, Lemma 5.1 and [28, Lemma 3.1.3], Hölder inequality, Young’s inequality, we derive
+$G_{323}(t)$
+$=$ $\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),\sigma(\overline{\mu}^{Y,N}(u))\rangle{\rm d}B_{u}^{H,i}$
+$\times\bigg\langle\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u)){\rm d}u,b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))\bigg\rangle{\rm d}s$
 $+\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),\sigma(\overline{\mu}^{Y,N}(u))\rangle{\rm d}B_{u}^{H,i}$
-×$\times\bigg{\langle}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}\sigma(\overline{\mu}^{Y,N}(u)){\rm d}B_{u}^{H,i},b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\bigg{\rangle}{\rm d}s$
-$=$ $\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),\sigma(\overline{\mu}^{Y,N}(u))\rangle{\rm d}B_{u}^{H,i}$
-×$\times\bigg{\langle}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}\sigma(\overline{\mu}^{Y,N}(u)){\rm d}B_{u}^{H,i},b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\bigg{\rangle}{\rm d}s$
-$\leq$ $2HT^{2H-2}\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-3}{2}}|\sigma(\overline{\mu}^{Y,N}(u))|^{2}{\rm d}u$
-×$\times|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))|{\rm d}s$
-$\leq$ $C\Delta^{1-\alpha}\mathbb{E}\int_{0}^{t}(|Y_{\Delta}^{i,N}(s)|^{\overline{p}}+|Y_{\Delta}^{i,N}(s-\tau)|^{\overline{p}}+|\overline{Y}_{\Delta}^{i,N}(s))|^{\overline{p}}){\rm d}s$
-$+C\Delta^{1-\alpha}\Delta^{(1-\alpha)\overline{p}}\mathbb{E}\int_{0}^{t}|\overline{Y}_{\Delta}^{i,N}(s))|^{\overline{p}}{\rm d}s$
-$\leq$ $C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(s)|^{\overline{p}}{\rm d}s.$
-To sum up, we derive
-$G_{3}(t)\leq C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(s)|^{\overline{p}}{\rm d}s.$
-Combining $G_{1}-G_{3}$, we get
-(5.12) $\sup_{0\leq u\leq t}\mathbb{E}|\widetilde{Y}_{\Delta}^{i,N}(u)|^{\overline{p}}\leq\sup_{0\leq u\leq t}\mathbb{E}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\overline{p}/2}\leq C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)|^{\overline{p}}{\rm d}s.$
-Noting that $|x-y|^{\overline{p}}\geq 2^{1-\overline{p}}|x|^{\overline{p}}-|y|^{\overline{p}}$, by Assumption 2.3, Lemma 5.1, we get
+$\times\bigg\langle\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}\sigma(\overline{\mu}^{Y,N}(u)){\rm d}B_{u}^{H,i},b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))\bigg\rangle{\rm d}s$
+$\leq$ $\int_{0}^{t}\bigg\{\bigg[\mathbb{E}\bigg(\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),\sigma(\overline{\mu}^{Y,N}(u))\rangle{\rm d}B_{u}^{H,i}\bigg)^{2}\bigg]^{1/2}$
+$\times\bigg[\mathbb{E}\bigg(\bigg\langle\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u)){\rm d}u,b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\bigg\rangle\bigg)^{2}\bigg]^{1/2}\bigg\}{\rm d}s$
+$+$ $\int_{0}^{t}\bigg\{\bigg[\mathbb{E}\bigg(\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),\sigma(\overline{\mu}^{Y,N}(u))\rangle{\rm d}B_{u}^{H,i}\bigg)^{2}\bigg]^{1/2}$
+$\times\bigg[\mathbb{E}\bigg(\bigg\langle\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}\sigma(\overline{\mu}^{Y,N}(u)){\rm d}B_{u}^{H,i},b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))\bigg\rangle\bigg)^{2}\bigg]^{1/2}\bigg\}{\rm d}s$
+$\leq$ $2H{\Delta}^{2H-2}\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-3}{2}}|\sigma(\overline{\mu}^{Y,N}(u))|^{2}{\rm d}u$
+$\times\bigg\{\bigg(\mathbb{E}\Delta^{4-4\alpha}(1+|\overline{Y}_{\Delta}^{i,N}(s)|+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|)^{4}\bigg)^{1/4}\bigg(\mathbb{E}\Delta^{-4\alpha}(1+|\overline{Y}_{\Delta}^{i,N}(s)|+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|)^{4}\bigg)^{1/4}$
+$+\bigg(\mathbb{E}\Delta^{4}(1+|\overline{Y}_{\Delta}^{i,N}(s)|)^{4}\bigg)^{1/4}\bigg(\mathbb{E}\Delta^{-4\alpha}(1+|\overline{Y}_{\Delta}^{i,N}(s)|+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|)^{4}\bigg)^{1/4}{\rm d}s$
+$\leq$ $2H{\Delta}^{2H-2}\mathbb{E}\int_{0}^{t}\bigg(\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-3}{2}}|\sigma(\overline{\mu}^{Y,N}(u))|^{2}{\rm d}u\times C\Delta^{1-\alpha}\mathbb{E}\int_{0}^{t}(1+|Y_{\Delta}^{i,N}(s)|^{\overline{p}})\bigg){\rm d}s$
+$\leq$ $C\Delta^{2H-1-\alpha}\int_{0}^{t}[C\Delta(1+\sup_{0\leq v\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(v)|^{\overline{p}})]\times[C(1+\sup_{0\leq v\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(v)|^{\overline{p}})]\mathrm{d}s$
+$=$ $C\Delta^{2H-\alpha}\int_{0}^{t}(1+\sup_{0\leq v\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(v)|^{\overline{p}})^{2}\mathrm{d}s$
+$\leq$ $C\Delta^{2H-1-\alpha}(1+\sup_{0\leq v\leq t}\mathbb{E}|Y_{\Delta}^{i,N}(v)|^{\overline{p}})\int_{0}^{t}(1+\sup_{0\leq v\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(v)|^{\overline{p}})\mathrm{d}s$
+$\leq$ $C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)|^{\overline{p}}{\rm d}s.$
+To sum up, we conclude
+$G_{3}(t)\leq C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)|^{\overline{p}}{\rm d}s+C\int_{0}^{t}(t-s)^{2H-1}\bigg(\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)|^{\overline{p}}\bigg){\rm d}s.$
+Combining $G_{1}(t)-G_{3}(t)$ together, it shows
+(5.12) $\sup_{0\leq u\leq t}\mathbb{E}|\widetilde{Y}_{\Delta}^{i,N}(u)|^{\overline{p}}\leq$ $\sup_{0\leq u\leq t}\mathbb{E}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\overline{p}/2}\leq C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)|^{\overline{p}}{\rm d}s$
+$+C\int_{0}^{t}(t-s)^{2H-1}\bigg(\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)|^{\overline{p}}\bigg){\rm d}s.$
+Noting that $|x-y|^{\overline{p}}\geq 2^{1-\overline{p}}|x|^{\overline{p}}-|y|^{\overline{p}}$, Assumption 2.3 and Lemma 5.1 give
 $|\widetilde{Y}_{\Delta}^{i,N}(u)|^{\overline{p}}\geq$ $2^{1-\overline{p}}|Y_{\Delta}^{i,N}(t)-D(Y_{\Delta}^{i,N}(t-\tau))|^{\overline{p}}-|\theta b_{\Delta}(Y_{\Delta}^{i,N}(t),Y_{\Delta}^{i,N}(t-\tau),\mu^{Y,N}(t))\Delta|^{\overline{p}}$
 $\geq$ $2^{1-\overline{p}}[2^{1-\overline{p}}|Y_{\Delta}^{i,N}(t)|^{\overline{p}}-|D(Y_{\Delta}^{i,N}(t-\tau))|^{\overline{p}}]-|\theta b_{\Delta}(Y_{\Delta}^{i,N}(t),Y_{\Delta}^{i,N}(t-\tau),\mu^{Y,N}(t))\Delta|^{\overline{p}}$
-$\geq$ $(2^{2-2\overline{p}}-4^{\overline{p}-1}\theta^{\overline{p}}K_{7}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}})|Y_{\Delta}^{i,N}(t)|^{\overline{p}}-(2^{1-\overline{p}}\lambda^{\overline{p}}+4^{\overline{p}-1}\theta^{p}K_{7}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}})|Y_{\Delta}^{i,N}(t-\tau)|^{\overline{p}}$
-$-4^{\overline{p}-1}\theta^{\overline{p}}K_{7}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}}(\mathbb{W}_{q}(\mu^{Y,N}(t)),\delta_{0})^{\overline{p}}-4^{\overline{p}-1}\theta^{\overline{p}}K_{7}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}}.$
+$\geq$ $(2^{2-2\overline{p}}-4^{\overline{p}-1}\theta^{\overline{p}}K_{6}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}})|Y_{\Delta}^{i,N}(t)|^{\overline{p}}-(2^{1-\overline{p}}\lambda^{\overline{p}}+4^{\overline{p}-1}\theta^{p}K_{6}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}})|Y_{\Delta}^{i,N}(t-\tau)|^{\overline{p}}$
+$-4^{\overline{p}-1}\theta^{\overline{p}}K_{6}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}}\mathbb{W}_{2}^{\overline{p}}(\mu^{Y,N}(t),\delta_{0})-4^{\overline{p}-1}\theta^{\overline{p}}K_{6}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}}.$
 Thus, we obtain
-$\sup_{0\leq u\leq t}\mathbb{E}|Y_{\Delta}^{i,N}(t)|^{\overline{p}}\leq(2^{2-2\overline{p}}-3\times 4^{\overline{p}-1}\theta^{\overline{p}}K_{7}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}}-2^{1-\overline{p}}\lambda^{\overline{p}})^{-1}$
-×$\times\bigg{[}\sup_{0\leq u\leq t}\mathbb{E}|\widetilde{Y}_{\Delta}^{i,N}(u)|^{\overline{p}}+4^{\overline{p}-1}\theta^{\overline{p}}K_{7}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}}+(2^{1-\overline{p}}\lambda^{\overline{p}}+4^{\overline{p}-1}\theta^{\overline{p}}K_{7}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}})\mathbb{E}\parallel\xi\parallel^{\overline{p}}\bigg{]}.$
-Then, we derive from (5.12),
-$\sup_{0\leq u\leq t}\mathbb{E}|Y_{\Delta}^{i,N}(t)|^{\overline{p}}\leq C+C\int_{0}^{t}\sup_{0\leq u\leq t}\mathbb{E}|Y_{\Delta}^{i,N}(t)|^{\overline{p}}.$
-Applying the Gronwall inequality, we complete the proof. ∎
+$\sup_{0\leq u\leq t}\mathbb{E}|Y_{\Delta}^{i,N}(t)|^{\overline{p}}\leq(2^{2-2\overline{p}}-3\times 4^{\overline{p}-1}\theta^{\overline{p}}K_{6}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}}-2^{1-\overline{p}}\lambda^{\overline{p}})^{-1}$
+$\times\bigg[\sup_{0\leq u\leq t}\mathbb{E}|\widetilde{Y}_{\Delta}^{i,N}(u)|^{\overline{p}}+4^{\overline{p}-1}\theta^{\overline{p}}K_{6}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}}+(2^{1-\overline{p}}\lambda^{\overline{p}}+4^{\overline{p}-1}\theta^{\overline{p}}K_{6}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}})\mathbb{E}\|\xi\|^{\overline{p}}\bigg].$
+Further, it follows from (5.12) that
+$\sup_{0\leq u\leq t}\mathbb{E}|Y_{\Delta}^{i,N}(u)|^{\overline{p}}\leq C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)|^{\overline{p}}{\rm d}s+C\int_{0}^{t}(t-s)^{2H-1}\bigg(\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)|^{\overline{p}}\bigg){\rm d}s.$
+Consequently, the desired result can be shown by applying Lemma 2.8. This completes the proof. ∎
 5.2 Strong Convergence Rate
-Lemma 5.5.
-Let Assumptions 2.2-2.5 hold. For $p>1/H$, then
+Lemma $5.$
+Let Assumptions 2.2-2.5 hold. For $p\geq 2$, then
 $\mathbb{E}\bigg(\sup_{0\leq k\leq M-1}\sup_{t_{k}\leq t\leq t_{k+1}}|Y_{\Delta}^{i,N}(t)-Y_{\Delta}^{i,N}(t_{k})|^{p}\bigg)\leq C\Delta^{[(1-\alpha)\wedge H]p},$
@@
 $+2^{p-1}\mathbb{E}\bigg(\sup_{t_{k}\leq t\leq t_{k+1}}\bigg\lvert\int_{t_{k}}^{t}\sigma(\overline{\mu}^{Y,N}(s)){\rm d}B_{s}^{H,i}\bigg\rvert^{p}\bigg)$
-=:absent:=:= : $J_{1}+J_{2}.$
-By the Hölder inequality and Lemmas 5.1, 5.4, we get
-$J_{1}\leq$ $2^{p-1}\Delta^{p-1}\mathbb{E}\int_{t_{k}}^{t_{k+1}}|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))|^{p}{\rm d}s$
-$\leq$ $2^{p-1}\Delta^{p-1}\mathbb{E}\int_{t_{k}}^{t_{k+1}}K_{6}\Delta^{-\alpha p}(1+|\overline{Y}_{\Delta}^{i,N}(s)|+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|+\mathbb{W}_{q}(\overline{\mu}^{Y,N}(s),\delta_{0}))^{p}{\rm d}s$
+$:=$ $W_{1}(t)+W_{2}(t).$
+With the Hölder inequality and Lemmas 5.1, 5.4, it follows
+$W_{1}(t)=$ $2^{p-1}\mathbb{E}\bigg(\sup_{t_{k}\leq t\leq t_{k+1}}\bigg\lvert\int_{t_{k}}^{t}b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s)){\rm d}s\bigg\rvert^{p}\bigg)$
+$\leq$ $2^{p-1}\Delta^{p-1}\mathbb{E}\int_{t_{k}}^{t_{k+1}}|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))|^{p}{\rm d}s$
+$\leq$ $2^{p-1}\Delta^{p-1}\mathbb{E}\int_{t_{k}}^{t_{k+1}}K_{6}\Delta^{-\alpha p}(1+|\overline{Y}_{\Delta}^{i,N}(s)|+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|+\mathbb{W}_{2}^{p}(\overline{\mu}^{Y,N}(s),\delta_{0})){\rm d}s$
 $\leq$ $C\Delta^{(1-\alpha)p}+C\Delta^{(1-\alpha)p}\mathbb{E}\int_{t_{k}}^{t_{k+1}}(|\overline{Y}_{\Delta}^{i,N}(s)|^{p}+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|^{p}){\rm d}s$
 $\leq$ $C\Delta^{(1-\alpha)p}.$
-Using the same technique as in (3.3), along with the Hölder inequality, the Fubini theorem and Lemmas 2.5, 5.4, we get
-$J_{2}\leq C\Delta^{pH-1}\int_{t_{k}}^{t_{k+1}}\mathbb{E}(1+\mathbb{W}_{\theta}(\overline{\mu}_{s}^{Y,N},\delta_{0}))^{p}{\rm d}s\leq C\Delta^{pH}.$
-Recalling that $\alpha\in(0,1/2]$, we obtian
+Using the same techniques as the proof process of [15, Theorem 3.1], along with the Hölder inequality, the Fubini theorem and Lemmas 2.5, 5.4, we derive
+$W_{2}(t)=$ $2^{p-1}\mathbb{E}\bigg(\sup_{t_{k}\leq t\leq t_{k+1}}\bigg\lvert\int_{t_{k}}^{t}\sigma(\overline{\mu}^{Y,N}(s)){\rm d}B_{s}^{H,i}\bigg\rvert^{p}\bigg)$
+$\leq$ $C\Delta^{pH-1}\mathbb{E}\int_{t_{k}}^{t_{k+1}}(1+\mathbb{W}_{2}(\overline{\mu}_{s}^{Y,N},\delta_{0}))^{p}{\rm d}s\leq C\Delta^{pH}.$
+Recalling that $\alpha\in(0,1/2]$, then we conclude
 $\mathbb{E}\bigg(\sup_{t_{k}\leq t\leq t_{k+1}}|\widetilde{Y}_{\Delta}^{i,N}(t)-\widetilde{Y}_{\Delta}^{i,N}(t_{k})|^{p}\bigg)\leq C\Delta^{[(1-\alpha)\wedge H]p}.$
-Denoting by $\widetilde{D}(t,t_{k}):=D(Y_{\Delta}^{i,N}(t-\tau))-D(Y_{\Delta}^{i,N}(t_{k}-\tau))$, and $\widetilde{b}(t,t_{k}):=b_{\Delta}(Y_{\Delta}^{i,N}(t),Y_{\Delta}^{i,N}(t-\tau),\mu^{Y,N}(t))-b_{\Delta}(Y_{\Delta}^{i,N}(t_{k}),Y_{\Delta}^{i,N}(t_{k}-\tau),\mu^{Y,N}(t_{k}))$, together with Assumption 2.3, we get
+Denote $\widetilde{D}(t,t_{k}):=D(Y_{\Delta}^{i,N}(t-\tau))-D(Y_{\Delta}^{i,N}(t_{k}-\tau))$, and $\widetilde{b}(t,t_{k}):=b_{\Delta}(Y_{\Delta}^{i,N}(t),Y_{\Delta}^{i,N}(t-\tau),\mu^{Y,N}(t))-b_{\Delta}(Y_{\Delta}^{i,N}(t_{k}),Y_{\Delta}^{i,N}(t_{k}-\tau),\mu^{Y,N}(t_{k}))$, together with Assumption 2.3
 (5.13) $|\widetilde{Y}_{\Delta}^{i,N}(t)-\widetilde{Y}_{\Delta}^{i,N}(t_{k})|^{p}$
@@
 $\geq$ $2^{2-2p}|Y_{\Delta}^{i,N}(t)-Y_{\Delta}^{i,N}(t_{k})|^{p}-2^{1-p}\lambda^{p}|Y_{\Delta}^{i,N}(t-\tau)-Y_{\Delta}^{i,N}(t_{k}-\tau)|^{p}-\theta^{p}\Delta^{p}|\widetilde{b}(t,t_{k})|^{p}.$
-For $0\leq t\leq t_{1}=\Delta$, then $t-\tau<t_{1}-\tau<0$, we derive from (5.13) and Lemma 5.4,
+For $0\leq t<t_{1}=\Delta$, then $t-\tau<t_{1}-\tau<0$, we derive from (5.13) and Lemma 5.4,
 $\mathbb{E}\bigg(\sup_{0\leq t\leq t_{1}}|Y_{\Delta}^{i,N}(t)-Y_{\Delta}^{i,N}(t_{0})|^{p}\bigg)\leq C\mathbb{E}\bigg(\sup_{0\leq t\leq t_{1}}|\widetilde{Y}_{\Delta}^{i,N}(t)-\widetilde{Y}_{\Delta}^{i,N}(t_{0})|^{p}\bigg)+C\Delta^{(1-\alpha)p}\leq C\Delta^{[(1-\alpha)\wedge H]p}.$
@@
 $\leq$ $C\Delta^{[(1-\alpha)\wedge H]p}.$
-By induction,
-$\mathbb{E}\bigg{(}\sup_{t_{k}\leq t\leq t_{k+1}}|Y_{\Delta}^{i,N}(t)-Y_{\Delta}^{i,N}(t_{1})|^{p}\bigg{)}\leq$ $C\mathbb{E}\bigg{(}\sup_{t_{k}\leq t\leq t_{k+1}}|\widetilde{Y}_{\Delta}^{i,N}(t)-\widetilde{Y}_{\Delta}^{i,N}(t_{k})|^{p}\bigg{)}+C\Delta^{(1-\alpha)p}$
-$\leq$ $C\Delta^{[(1-\alpha)\wedge H]p}.$
-we get the desired assertion. ∎
+By induction, the desired assertion can be shown. ∎
 Lemma 5.6.
-Let Assumptions 2.2-2.5 and Lemmas 5.1, 5.2 hold, then it holds that for $p>1/H$ and $\overline{p}>2(l+1)p$,
+Let Assumptions 2.2-2.5 hold. Fix any $\overline{p}>4(l+1)$ in Lemma 5.4. Then it holds that for $2\leq p<\overline{p}/2(l+1)$,
 $\mathbb{E}\bigg(\sup_{0\leq t\leq T}|X_{t}^{i,N}-Y_{\Delta}^{i,N}(t)|^{p}\bigg)\leq C\Delta^{\alpha p},$
@@
 Proof.
-For $t\in[0,T]$, denote $I(t)=\widetilde{Y}_{\Delta}^{i,N}(t)-X_{t}^{i,N}+D(X_{t-\tau}^{i,N})$, then
-$I(t)=$ $I(0)+\int_{0}^{t}(b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})){\rm d}s$
+For $t\in[0,T]$, denote $\Lambda(t)=\widetilde{Y}_{\Delta}^{i,N}(t)-X_{t}^{i,N}+D(X_{t-\tau}^{i,N})$, then
+$\Lambda(t)=$ $\Lambda(0)+\int_{0}^{t}(b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})){\rm d}s$
 $+\int_{0}^{t}(\sigma(\overline{\mu}^{Y,N}(s))-\sigma(\mu_{s}^{X,N})){\rm d}B_{s}^{H,i},$
-where $I_{0}=-\theta b_{\Delta}(\xi(0),\xi(-\tau),\delta_{0})\Delta$. The Itô formula yields,
-$|I(t)|^{p}\leq$ $|I(0)|^{p}$
-$+p\int_{0}^{t}|I(s)|^{p-2}\langle I(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})\rangle{\rm d}s$
-$+HT^{2H-1}p(p-1)\int_{0}^{t}|I(s)|^{p-2}|\sigma(\overline{\mu}^{Y,N}(s))-\sigma(\mu_{s}^{X,N})|^{2}{\rm d}s$
-$+p\int_{0}^{t}|I(s)|^{p-2}\langle I(s),\sigma(\overline{\mu}^{Y,N}(s))-\sigma(\mu_{s}^{X,N})\rangle{\rm d}B_{s}^{H,i}$
-$\leq$ $|I(0)|^{p}+\sum_{i=1}^{6}H_{i}(t),$
+where $\Lambda({0})=-\theta b_{\Delta}(\xi(0),\xi(-\tau),\delta_{0})\Delta$. The fractional Itô formula yields,
+$|\Lambda(t)|^{p}\leq$ $|\Lambda(0)|^{p}$
+$+p\int_{0}^{t}|\Lambda(s)|^{p-2}\langle\Lambda(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})\rangle{\rm d}s$
+$+H(2H-1)p(p-1)\int_{0}^{t}|\Lambda(s)|^{p-2}\int_{0}^{s}(s-r)^{2H-2}\|\sigma(\overline{\mu}^{Y,N}(r))-\sigma(\mu_{r}^{X,N})\|^{2}{\rm d}r{\rm d}s$
+$+p\int_{0}^{t}|\Lambda(s)|^{p-2}\langle\Lambda(s),\sigma(\overline{\mu}^{Y,N}(s))-\sigma(\mu_{s}^{X,N})\rangle{\rm d}B_{s}^{H,i}$
+$\leq$ $|\Lambda(0)|^{p}+\sum_{i=1}^{6}Z_{i}(t),$
 where
-$H_{1}(t)=$ $p\int_{0}^{t}|I(s)|^{p-2}\langle I(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))$
+$Z_{1}(t)=$ $p\int_{0}^{t}|\Lambda(s)|^{p-2}\langle\Lambda(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))$
 $-b(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))\rangle{\rm d}s$
-$H_{2}(t)=$ $p\int_{0}^{t}|I(s)|^{p-2}\langle I(s),b(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))$
+$Z_{2}(t)=$ $p\int_{0}^{t}|\Lambda(s)|^{p-2}\langle\Lambda(s),b(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))$
 $-b(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))\rangle{\rm d}s$
-$H_{3}(t)=$ $p\int_{0}^{t}|I(s)|^{p-2}\langle I(s),b(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))$
-$-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})\rangle{\rm d}s,$
-$H_{4}(t)=$ $HT^{2H-1}p(p-1)\int_{0}^{t}|I(s)|^{p-2}|\sigma(\overline{\mu}^{Y,N}(s))-\sigma(\mu^{Y,N}(s))|^{2}{\rm d}s$
-$H_{5}(t)=$ $HT^{2H-1}p(p-1)\int_{0}^{t}|I(s)|^{p-2}|\sigma(\mu^{Y,N}(s))-\sigma(\mu_{s}^{X,N})|^{2}{\rm d}s$
-$H_{6}(t)=$ $p\int_{0}^{t}|I(s)|^{p-2}\langle I(s),\sigma(\overline{\mu}^{Y,N}(s))-\sigma(\mu_{s}^{X,N})\rangle{\rm d}B_{s}^{H,i}.$
+$Z_{3}(t)=$ $p\int_{0}^{t}|\Lambda(s)|^{p-2}\langle\Lambda(s),b(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})\rangle{\rm d}s,$
+$Z_{4}(t)=$ $H(2H-1)p(p-1)\int_{0}^{t}|\Lambda(s)|^{p-2}\int_{0}^{s}(s-r)^{2H-2}\|\sigma(\overline{\mu}^{Y,N}(r))-\sigma(\mu^{Y,N}(r))\|^{2}{\rm d}r{\rm d}s$
+$Z_{5}(t)=$ $H(2H-1)p(p-1)\int_{0}^{t}|\Lambda(s)|^{p-2}\int_{0}^{s}(s-r)^{2H-2}\|\sigma(\mu^{Y,N}(r))-\sigma(\mu_{r}^{X,N})\|^{2}{\rm d}r{\rm d}s$
+$Z_{6}(t)=$ $p\int_{0}^{t}|\Lambda(s)|^{p-2}\langle\Lambda(s),\sigma(\overline{\mu}^{Y,N}(s))-\sigma(\mu_{s}^{X,N})\rangle{\rm d}B_{s}^{H,i}.$
 By Assumption 2.3, Lemmas 5.1, 5.2, 5.4, and the Young’s inequality, the Hölder inequality, we get
-$\mathbb{E}\bigg{(}\sup_{0\leq u\leq t}H_{1}(u)\bigg{)}$
-$\leq$ $C\mathbb{E}\int_{0}^{t}|I(s)|^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))$
+$\mathbb{E}\bigg(\sup_{0\leq u\leq t}Z_{1}(u)\bigg)$
+$\leq$ $C\mathbb{E}\int_{0}^{t}|\Lambda(s)|^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))$
 $-b(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))|^{p}{\rm d}s$
@@
 $+\theta^{p}\Delta^{p}|b_{\Delta}(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))|^{p}\bigg){\rm d}s$
-$+C\Delta^{\alpha p}\mathbb{E}\int_{0}^{t}(1+|\overline{Y}_{\Delta}^{i,N}(s)|^{2(l+1)p}+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|^{2(l+1)p}+(\mathbb{W}_{q}(\overline{\mu}^{Y,N},\delta_{0}))^{2p}){\rm d}s$
-$\leq$ $C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(s)-X_{s}^{i,N}|^{p}+C\Delta^{(1-\alpha)p}+C\Delta^{\alpha p},$
-where for the last term we used the fact that $\overline{p}>2(l+1)p$. By the elementary inequality, the Young’s inequality, the Hölder inequality, Assumptions 2.3, 2.4 and Lemmas 5.1, 5.4, 5.5, we get
-$\mathbb{E}\bigg{(}\sup_{0\leq u\leq t}H_{2}(u)\bigg{)}$
-$\leq$ $C\mathbb{E}\int_{0}^{t}|I(s)|^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}|b(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))-b(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))|^{p}{\rm d}s$
+$+C\Delta^{\alpha p}\mathbb{E}\int_{0}^{t}(1+|\overline{Y}_{\Delta}^{i,N}(s)|^{2(l+1)p}+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|^{2(l+1)p}+\mathbb{W}_{2}^{2p}(\overline{\mu}^{Y,N},\delta_{0})){\rm d}s$
+$\leq$ $C\int_{0}^{t}\mathbb{E}\left(\sup_{0\leq u\leq s}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}\right){\rm d}s+C\Delta^{(1-\alpha)p}+C\Delta^{\alpha p},$
+where for the last term we used the fact that $2(l+1)p<\overline{p}$.
+Combining the elementary inequality, the Young’s inequality, the Hölder inequality, Assumptions 2.3, 2.4 and Lemmas 5.1, 5.4, 5.5, it follows
+$\mathbb{E}\bigg(\sup_{0\leq u\leq t}Z_{2}(u)\bigg)$
+$\leq$ $C\mathbb{E}\int_{0}^{t}|\Lambda(s)|^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}|b(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))-b(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))|^{p}{\rm d}s$
 $\leq$ $C\mathbb{E}\int_{0}^{t}\bigg(|Y_{\Delta}^{i,N}(s)-X_{s}^{i,N}|^{p}+|Y_{\Delta}^{i,N}(s-\tau)-X_{s-\tau}^{i,N}|^{p}+\theta^{p}\Delta^{p}|b_{\Delta}(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))|\bigg){\rm d}s$
 $+C\mathbb{E}\int_{0}^{t}\bigg[(1+|\overline{Y}_{\Delta}^{i,N}(s)|^{l}+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|^{l}+|Y_{\Delta}^{i,N}(s)|^{l}+|Y_{\Delta}^{i,N}(s-\tau)|^{l})^{p}$
-×$\times(|\overline{Y}_{\Delta}^{i,N}(s)-Y_{\Delta}^{i,N}(s)|+|\overline{Y}_{\Delta}^{i,N}(s-\tau)-Y_{\Delta}^{i,N}(s-\tau)|+(\mathbb{W}_{q}(\overline{\mu}^{Y,N}(s),\mu^{Y,N})(s)))^{p}\bigg{]}{\rm d}s$
+$\times(|\overline{Y}_{\Delta}^{i,N}(s)-Y_{\Delta}^{i,N}(s)|+|\overline{Y}_{\Delta}^{i,N}(s-\tau)-Y_{\Delta}^{i,N}(s-\tau)|+\mathbb{W}_{2}^{p}(\overline{\mu}^{Y,N}(s),\mu^{Y,N})(s))\bigg]{\rm d}s$
 $\leq$ $C\mathbb{E}\int_{0}^{t}\bigg(|Y_{\Delta}^{i,N}(s)-X_{s}^{i,N}|^{p}+|Y_{\Delta}^{i,N}(s-\tau)-X_{s-\tau}^{i,N}|^{p}\bigg){\rm d}s+C\Delta^{(1-\alpha)p}$
 $+C\int_{0}^{t}[\mathbb{E}(1+|\overline{Y}_{\Delta}^{i,N}(s)|^{l}+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|^{l}+|Y_{\Delta}^{i,N}(s)|^{l}+|Y_{\Delta}^{i,N}(s-\tau)|^{l})^{2p}]^{1/2}$
-×$\times[\mathbb{E}(|\overline{Y}_{\Delta}^{i,N}(s)-Y_{\Delta}^{i,N}(s)|+|\overline{Y}_{\Delta}^{i,N}(s-\tau)-Y_{\Delta}^{i,N}(s-\tau)|)^{2p}]^{1/2}{\rm d}s$
+$\times[\mathbb{E}(|\overline{Y}_{\Delta}^{i,N}(s)-Y_{\Delta}^{i,N}(s)|+|\overline{Y}_{\Delta}^{i,N}(s-\tau)-Y_{\Delta}^{i,N}(s-\tau)|)^{2p}]^{1/2}{\rm d}s$
 $+C\int_{0}^{t}\mathbb{E}(|\overline{Y}_{\Delta}^{i,N}(s)-Y_{\Delta}^{i,N}(s)|)^{p}{\rm d}s$
-$\leq$ $C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}{\rm d}s+C\Delta^{(1-\alpha)p}+C\Delta^{[(1-\alpha)\wedge H]p}.$
-By Assumption 2.4 and Lemmas 5.1, 5.4, we get
-$\mathbb{E}\bigg{(}\sup_{0\leq u\leq t}H_{3}(u)\bigg{)}$
-$\leq$ $C\mathbb{E}\int_{0}^{t}|I(s)|^{p-2}[|Y_{\Delta}^{i,N}(s)-X_{s}^{i,N}|^{2}+|Y_{\Delta}^{i,N}(s-\tau)-X_{s-\tau}^{i,N}|^{2}+(\mathbb{W}_{q}(\mu^{Y,N}(s),\mu^{X,N}(s)))^{2}]{\rm d}s$
-$+C\mathbb{E}\int_{0}^{t}|I(s)|^{p-2}|\theta b_{\Delta}(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))\Delta||b(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))$
-$-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})|{\rm d}s$
+$\leq$ $C\int_{0}^{t}\mathbb{E}\left(\sup_{0\leq u\leq s}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}\right){\rm d}s+C\Delta^{(1-\alpha)p}+C\Delta^{[(1-\alpha)\wedge H]p}.$
+Again, by Assumption 2.4 and Lemmas 5.1, 5.4
+$\mathbb{E}\bigg(\sup_{0\leq u\leq t}Z_{3}(u)\bigg)$
+$\leq$ $C\mathbb{E}\int_{0}^{t}|\Lambda(s)|^{p-2}[|Y_{\Delta}^{i,N}(s)-X_{s}^{i,N}|^{2}+|Y_{\Delta}^{i,N}(s-\tau)-X_{s-\tau}^{i,N}|^{2}+\mathbb{W}_{2}^{2}(\mu^{Y,N}(s),\mu^{X,N}(s))]{\rm d}s$
+$+C\mathbb{E}\int_{0}^{t}|\Lambda(s)|^{p-2}|\theta b_{\Delta}(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))\Delta|$
+$\cdot|b(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})|{\rm d}s$
 $\leq$ $C\mathbb{E}\int_{0}^{t}[|Y_{\Delta}^{i,N}(s)-X_{s}^{i,N}|^{p}+\theta^{p}\Delta^{p}|b_{\Delta}(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))|^{p}]{\rm d}s$
-$+C\Delta^{1-\alpha}\mathbb{E}\int_{0}^{t}|I(s)|^{p-2}[(1+|Y_{\Delta}^{i,N}(s)|^{l}+|Y_{\Delta}^{i,N}(s-\tau)|^{l}+|X_{s}^{i,N}|^{l}+|X_{s-\tau}^{i,N}|^{l})$
-×$\times(|Y_{\Delta}^{i,N}(s)-X_{s}^{i,N}|+|Y_{\Delta}^{i,N}(s-\tau)-X_{s-\tau}^{i,N}|)+\mathbb{W}_{q}(\mu^{Y,N}(s),\mu_{s}^{X,N})]{\rm d}s$
-$\leq$ $C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}{\rm d}s+C\Delta^{(1-\alpha)(1+p)}.$
-By the Young’s inequality, Assumption 2.5, we get
-$\mathbb{E}\bigg{(}\sup_{0\leq u\leq t}H_{4}(u)\bigg{)}+\mathbb{E}\bigg{(}\sup_{0\leq u\leq t}H_{5}(u)\bigg{)}$
-$\leq$ $C\mathbb{E}\int_{0}^{t}|I(s)|^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}(\mathbb{W}_{q}(\overline{\mu}^{Y,N}(s)),\mu^{Y,N}(s))^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}(\mathbb{W}_{q}(\mu^{Y,N}(s)),\mu_{s}^{X,N})^{p}{\rm d}s$
-$\leq$ $C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}{\rm d}s+C\Delta^{(1-\alpha)p}+C\Delta^{[(1-\alpha)\wedge H]p}.$
-In the way in (3.3), by the Young’s inequality, we get
-$\mathbb{E}\bigg{(}\sup_{0\leq u\leq t}|H_{6}(u)|^{p}\bigg{)}$
-$\leq$ $C\mathbb{E}\int_{0}^{t}|I(s)|^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}(\mathbb{W}_{q}(\overline{\mu}^{Y,N}(s)),\mu^{Y,N}(s))^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}(\mathbb{W}_{q}(\mu^{Y,N}(s)),\mu_{s}^{X,N})^{p}{\rm d}s$
-$\leq$ $C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}{\rm d}s+C\Delta^{(1-\alpha)p}+C\Delta^{[(1-\alpha)\wedge H]p}.$
-Combining $H_{1}-H_{6}$, we get
-$\mathbb{E}\bigg{(}\sup_{0\leq u\leq t}|I(u)|^{p}\bigg{)}\leq C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{0\leq u\leq s}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}\bigg{)}{\rm d}s+C\Delta^{\alpha p}.$
-Recalling the definition of $I(t)$ and by Assumption 2.3, we may compute
-$|I(t)|^{p}\geq$ $2^{1-p}|Y_{\Delta}^{i,N}(t)-X_{t}^{i,N}-D(Y_{\Delta}^{i,N}(t-\tau))+D(X_{t-\tau}^{i,N})|^{p}$
+$+C\Delta^{1-\alpha}\mathbb{E}\int_{0}^{t}|\Lambda(s)|^{p-2}[(1+|Y_{\Delta}^{i,N}(s)|^{l}+|Y_{\Delta}^{i,N}(s-\tau)|^{l}+|X_{s}^{i,N}|^{l}+|X_{s-\tau}^{i,N}|^{l})$
+$\times(|Y_{\Delta}^{i,N}(s)-X_{s}^{i,N}|+|Y_{\Delta}^{i,N}(s-\tau)-X_{s-\tau}^{i,N}|)+\mathbb{W}_{2}(\mu^{Y,N}(s),\mu_{s}^{X,N})]{\rm d}s$
+$\leq$ $C\int_{0}^{t}\mathbb{E}\left(\sup_{0\leq u\leq s}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}\right){\rm d}s+C\Delta^{(1-\alpha)(1+p)}.$
+Further, the Young’s inequality, Assumption 2.5, and Lemmas 5.1, 5.5 lead to
+$\mathbb{E}\bigg(\sup_{0\leq u\leq t}Z_{4}(u)\bigg)+\mathbb{E}\bigg(\sup_{0\leq u\leq t}Z_{5}(u)\bigg)$
+$\leq$ $C\mathbb{E}\int_{0}^{t}|\Lambda(s)|^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}(t-s)^{2H-1}\mathbb{W}_{2}^{p}(\overline{\mu}^{Y,N}(s),\mu^{Y,N}(s)){\rm d}s$
+$+C\mathbb{E}\int_{0}^{t}(t-s)^{2H-1}\mathbb{W}_{2}^{p}(\mu^{Y,N}(s),\mu_{s}^{X,N}){\rm d}s$
+$\leq$ $C\int_{0}^{t}\mathbb{E}\left(\sup_{0\leq u\leq s}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}\right){\rm d}s+C\int_{0}^{t}(t-s)^{2H-1}\mathbb{E}\bigg(\sup_{0\leq u\leq s}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}\bigg){\rm d}s$
+$+C\Delta^{(1-\alpha)p}+C\Delta^{[(1-\alpha)\wedge H]p}.$
+In the similar way as achieving (3.6), the Young’s inequality results in
+$\mathbb{E}\bigg(\sup_{0\leq u\leq t}|Z_{6}(u)|^{p}\bigg)\leq C\Delta^{(1-\alpha)p}+C\Delta^{[(1-\alpha)\wedge H]p}$
+$+C\epsilon\mathbb{E}\bigg(\sup_{0\leq u\leq t}|\Lambda(u)|^{p}\bigg)+C\int_{0}^{t}\mathbb{E}\left(\sup_{0\leq u\leq s}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}\right){\rm d}s.$
+Collecting the above estimates $Z_{1}-Z_{6}$ together
+$\mathbb{E}\bigg(\sup_{0\leq u\leq t}|\Lambda(u)|^{p}\bigg)\leq$ $C\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq u\leq s}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}\bigg){\rm d}s$
+$+C\int_{0}^{t}(t-s)^{2H-1}\mathbb{E}\left(\sup_{0\leq u\leq s}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}\right){\rm d}s+C\Delta^{\alpha p}.$
+Recalling the definition of $\Lambda(t)$ and using Assumption 2.3, we may compute
+$|\Lambda(t)|^{p}\geq$ $2^{1-p}|Y_{\Delta}^{i,N}(t)-X_{t}^{i,N}-D(Y_{\Delta}^{i,N}(t-\tau))+D(X_{t-\tau}^{i,N})|^{p}$
 $-|\theta b_{\Delta}(Y_{\Delta}^{i,N}(t),Y_{\Delta}^{i,N}(t-\tau),\mu^{Y,N}(t))\Delta|^{p}$
@@
 $-|\theta b_{\Delta}(Y_{\Delta}^{i,N}(t),Y_{\Delta}^{i,N}(t-\tau),\mu^{Y,N}(t))\Delta|^{p}.$
-By Lemma 5.1, 5.4, we derive
-$\mathbb{E}\bigg{(}\sup_{0\leq u\leq t}|X_{u}^{i,N}-Y_{\Delta}^{i,N}(u)|^{p}\bigg{)}\leq$ $\mathbb{E}\bigg{(}\sup_{0\leq u\leq t}|I(u)|^{p}\bigg{)}+C\Delta^{(1-\alpha)p}$
-$\leq$ $C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{0\leq u\leq s}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}\bigg{)}{\rm d}s+C\Delta^{\alpha p}.$
-Applying the Gronwall inequality, we can derive the assertion. ∎
+This gives from Lemmas 5.1, 5.4 that
+$\mathbb{E}\bigg(\sup_{0\leq u\leq t}|X_{u}^{i,N}-Y_{\Delta}^{i,N}(u)|^{p}\bigg)\leq\mathbb{E}\bigg(\sup_{0\leq u\leq t}|\Lambda(u)|^{p}\bigg)+C\Delta^{(1-\alpha)p}$
+$\leq$ $C\Delta^{\alpha p}+C\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq u\leq s}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}\bigg){\rm d}s$
+$+C\int_{0}^{t}(t-s)^{2H-1}\mathbb{E}\left(\sup_{0\leq u\leq s}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}\right){\rm d}s.$
+Since the function $E_{2H,1}((\Gamma(2H))^{1/2H}t)$ converges, we can utilize Lemma 2.8 to get the assertion. ∎
 Theorem 5.7.
-Let Assumptions 2.2-2.5 hold. For $p>1/H$ and $\overline{p}\geq 2(l+1)p$, it holds that
-$\mathbb{E}\bigg{(}\sup_{t\in[0,T]}|X_{t}^{i}-Y_{t}^{i.N}|^{p}\bigg{)}\leq C\times\left\{\begin{array}[]{ll}N^{-\frac{1}{2}}+\Delta^{\alpha p},\ \ p>\frac{d}{2},\\ N^{-\frac{1}{2}}\log(1+N)+\Delta^{\alpha p},\ p=\frac{d}{2},\\ N^{-\frac{p}{d}}+\Delta^{\alpha p},\ p\in[2,\frac{d}{2}),\end{array}\right.$
+Let Assumptions 2.2-2.5 hold. Fix any $\overline{p}>4(l+1)$. Then, for $p\in[2,\frac{\overline{p}}{2(l+1)})$, it holds that
+$\mathbb{E}\bigg(\sup_{t\in[0,T]}|X_{t}^{i}-Y_{\Delta}^{i,N}(t)|^{p}\bigg)\leq C\times\left\{\begin{array}[]{ll}N^{-\frac{1}{2}}+\Delta^{\alpha p},\ \ p>\frac{d}{2},\\ N^{-\frac{1}{2}}\log(1+N)+\Delta^{\alpha p},\ p=\frac{d}{2},\\ N^{-\frac{p}{d}}+\Delta^{\alpha p},\ p\in[2,\frac{d}{2}),\end{array}\right$.
 where $C$ is a constant independent of $N$ and $\Delta$.
@@
 By Lemmas 4.3 and 5.6, the result is obvious. ∎
+Remark 5.8.
+This paper introduces a tamed theta EM scheme for a class of McKean-Vlasov SDEs driven by fractional Brownian motions.
+In contrast to the existing literature, the proposed framework generalizes several tamed EM schemes studied in works such as [13] and [37].
+The convergence rate is rigorously derived by employing fractional Itô calculus, the fractional BDG inequality, and other pertinent inequalities.
+This methodology provides a new perspective for handling this category of fractional McKean-Vlasov equations.